### v5

# Preprint submitted to Journal

Abstract
This paper proposes an algebraic framework for analyzing event execution intervals and sequences, introducing the concept of Optional Intervals Event (OIE) as a core abstraction. The $\mathrm{OIE}$ 4-tuple realizes the mapping of real-world events, storing all feasible execution intervals and sub-event relationships. Based on this, two sequential operations are defined: Complete Sequential Addition ($\oplus$) models concurrent events with a certain degree of equal opportunity within a shared time domain, Complete Sequential Multiplication ($\otimes$) models sequential events with a strict order. And the algebraic properties and orbit spaces of these operations are analyzed.
Addressing the traditional challenge of “simultaneous start”, this paper argues that pointwise temporal equality of timestamps is an unattainable ideal due to physical limits of observational precision. Instead, using sequential operation, “simultaneous start” is reformulated as a certain degree of equal opportunity within an equipotential time window, rather than observed instantaneous equality. This paradigm shift emphasizes integrating subjective algebraic structures to move the perspective from passive observation to active constraint-aware planning, providing a more rigorous formal foundation for event scheduling, physical simulation, and computational modeling that can handle inherent observational uncertainties.
keywords:
Optional intervals event; Complete sequential addition; Complete sequential multiplication; Orbit space
1 Introduction
Let’s first discuss two questions.
1. How can we accurately describe the so-called “simultaneous start of two events”?
The following expressions are often used “Event X and event Y started at the same time”. But to what precision are they simultaneous? Are the values of these two starting timestamps exactly the same for any number of decimal places?
As of the end of 2024, the error of the most advanced atomic clock in the world is $8\times 10^{-19}$ seconds(see Aeppli et al. (2024)), that is to say, it cannot be guaranteed that the so-called simultaneously starting moments of the two observed events can be accurate to $10^{-1000}$ seconds. So, strictly speaking, this scenario is not truly “at the same time”, we should pursue a more rigorous expression. Can we use an operation *, represented by the following n-ary finitary operation
$(event_{1},\ event_{2},\ event_{3})$
, to indicate that $event_{1}$, $event_{2}$, and $event_{3}$ “start simultaneously” with an error undetectable by current technological means?
2. Why both head-to-head races and one-by-one races are capable of determining the first-place finisher.
In a 100 metres race, there are usually 8 athletes competing, it is a head-to-head type of race. In this paper, we use an n-ary finitary operation $\oplus$” to represent a race in which all athletes participate together. Suppose the 8 athletes are arranged according to the lane order, with the $i$-th athlete $athlete_{i}$ in the $i$-th lane. Then a 100 metres race can be expressed as
$\oplus(athlete_{1},\ athlete_{2},\ athlete_{3},\ \cdots,\ athlete_{8})$
If we swap the lanes of the athletes in lane 1 and lane 2, the race can be expressed as
$\oplus(athlete_{2},\ athlete_{1},\ athlete_{3},\ \cdots,\ athlete_{8})$
Swapping the lanes does not affect the starting method of each athlete, they all start running after hearing the starting gun. So intuitively, this operation exhibits a certain kind of transformation invariance.
Downhill skiing is a winter sport. Each skier competes on a track in a certain order one by one, and there can only be one person on the track at a time, so this is a sequential type of race. We use an n-ary finitary operator $\otimes$” to represent a race in which all skiers participate. Then a race with 6 skiers can be expressed as
$\otimes(skier_{1},\ skier_{2},\ skier_{3},\ \cdots,\ skier_{6})$
The skier represented on the left side of parameters starts the race earlier than the skier represented on the right side. Intuitively, this operation does not appear to exhibit transformation invariance.
Although 100 metres race and downhill skiing have the above differences, they are both races of speed and can determine the ranking of each person and ultimately decide the ownership of the medals. So why is this?
The essence of a speed race is “sorting the time costs from smallest to largest”. Both the 100 metres and downhill skiing can achieve speed race rankings. Is it because of some of their own properties that the comparison of time costs can be realized through them? What are the physical/mathematical properties among them that we haven’t summarized yet?
In the fields of physics and computer science, a large amount of research focuses on events, time and sequences(see Vaandrager (1991)). Most studies first observe crucial indicator moments, such as the starting timestamp and ending timestamp, and then introduce mathematical methods and models for analysis. People’s traditional understanding of “observation” depends on the result of event execution, but this can easily trap observers in a situation where “small-scale structures are indistinguishable, and large-scale structures cannot be fully grasped”.
This paper introduces a paradigm shift from passive observation to constraint-aware subjective planning of events. By leveraging mathematical frameworks of set theory, group theory, and algebraic structures (e.g., Cartesian product, orbit space, permutational equivalence), it establishes a novel formal algebraic methodology for the analysis of event execution, sequence relationships, and object motion.
The subsequent sections are structured as follows:
  • Section 2: Analysis of Event Properties – Formalizes the concept of an Event as an entity with three execution states (unstarted, running, finished), defines its starting and ending timestamps, and introduces the notions of atomic and composite events. The fundamental property of persistence of execution status is established.
  • Section 3: Optional Intervals Event (OIE) – Proposes the core abstraction of Optional Intervals Event, 4-tuples $(\mathcal{C},\mathcal{F},\mathcal{I},\mathcal{A})$ that map to real-world events. These 4-tuples store all feasible execution intervals and the relationships among sub-events, serving as formal tools for pre-execution planning. The concepts of atomic ($\textsf{Atom}\mathrm{OIE}$) and composite ($\textsf{Comp}\mathrm{OIE}$) $\mathrm{OIE}$ are defined.
  • Section 4: Two Sequential Operations – Introduces two n-ary finitary operations on $\mathrm{OIE}$ instances. Complete Sequential Addition $(\oplus)$ models concurrent events where all participants have a certain degree of equal opportunity to start and end within a specified domain. Complete Sequential Multiplication $(\otimes)$ models strictly ordered sequential events. Detailed operational rules, helper functions, and the handling of infeasible combinations (resulting in the void $\mathrm{OIE}$ instance $\mathrm{oie}_{\mathrm{void}}$) are provided.
  • Section 5: Algebraic Properties of Sequential Operations – Analyzes the algebraic structure of the two operations, proving properties such as closure, general non-commutativity, and permutation equivalence. The orbit space for each operation is defined, for a finite set of $\mathrm{OIE}$ instances complete sequential addition($\oplus$) yielding a single-orbit space due to permutational equivalence while complete sequential multiplication($\otimes$) - constrained by strict sequential order - potentially yielding multiple orbits.
  • Section 6: Implementation of OIE – Describes how $\mathrm{OIE}$ instances guide the execution of their mapped events. Two types of implementation are defined: one using only the overall interval set ($\mathcal{I}$) and another using the detailed schedule ($\mathcal{F}$) from a sequential operation’s result. A projection mechanism is established to link sequential operations with general $n$-ary finitary operations, explaining how competitions (like 100m dash and downhill skiing) deterministically produce a winner via the permutation-invariant $\min$ operation. In a computer science application, we map the stage structure of Merge Sort to a combination of algebraic operations with distinct commutativity properties for the first time, and analyze the feasibility of parallel sorting algebraically. In an example of probability theory, we have analyzed the use of $\mathrm{OIE}$ in classical probability, providing a more rigorous description for drawing lots.
  • Section 7: Full CSA Diagram – Presents the Full Complete Sequential Addition Diagram, a Cayley-table-like representation of the $\oplus$ operation within a finite commutative semigroup featuring an absorbing element. This diagram visually encodes the algebraic properties of complete sequential addition and is shown to have correspondences with natural language conjunction (“and”) and programmatic composition.
  • Section 8: In-depth Insights – Reflects on the epistemological implications of the framework. It reinterprets “simultaneous starting” not as observable pointwise equality but as planned equipotentiality within a domain, demoting the observer’s role from arbiter to consumer of pre-validated plans. It also discusses the generalization of the $\mathrm{OIE}$ model to multi-dimensional physical quantities beyond time.
  • Section 9: Comparison – In the field of computer science, we have conducted comparisons between the content of this paper and classical algebraic and formal models for event processing. In the fields of physics and philosophy, we have further drawn comparisons between our interpretation and various physical and philosophical perspectives with respect to the understanding of simultaneous initiation.
  • Section 10: Derived Implications – Outlines promising directions for future work, including introducing the category theory, extensions of the sequential operations (e.g., “incomplete” variants), application to discrete time models, variable generalization beyond time, integration of probability, exploration of other operation types (subtraction, division), further study of the Full CSA diagram’s structure, applicability to non-classical mechanics.
  • Appendix: Simulation Code and Diagram Generation Code We provide the urls of simulation code and operation table code.
The appendices provide references to simulation code for the proposed operations and diagram generation.
2 Analysis of the properties of event
Event is a noun with complex meanings in both practical and philosophical aspects. It is a fundamental concept frequently used in the fields of physics, computer science, engineering and social sciences (see Casati and Varzi (2023)). Generally, we refer to the event’s class as “Event”. An instance of Event, as a specific and concrete entity, is represented by the term “event”. In this Section, we discuss some basic properties of Event.
Subsection 2.1 clarifies the three execution states of an event, the definitions and core properties of starting and ending timestamps, and standardizes the representation form of time intervals.
Subsection 2.2 defines the concepts of atomic events and composite events, and provides the calculation method for the start and end timestamps of composite events.
Subsection 2.3 proposes the abstract concept of Event with Undetermined Interval ($\mathrm{Event}^{*}$), realizes flexible mapping of real-world events by replacing fixed time intervals with undetermined variables, and distinguishes its atomic and composite types.
2.1 Execution status, starting timestamp and ending timestamp of event
According to the execution situation, an Event instance event has three execution status: unstarted, running and finished. These 3 status can be transformed based on the occurrence, progress and conclusion of the event. For example:
Before event is executed, it remains in the unstarted status all the time.
After event is executed and completed, it stays in the finished status permanently.
The moment event starts to be executed, its execution status changes from unstarted to running.
The moment event ends, its execution status changes from running to finished.
Since the Event symbolizes motion and is dynamic rather than static, a certain status of an Event instance must persist within a time interval. It cannot be in a particular status in an isolated way at a discrete moment that is different from the status at all neighboring moments. We call this the “persistence of the event execution status”.
Property 1 (Persistence of event execution status).
Suppose there exists an Event instance event. If event is in one of the three execution status at timestamp $t_{0}$, then there must be a real number $\epsilon>0$, such that, among the left punctured neighborhood set
$\mathcal{S}_{left}=\{t|t\in(t_{0}-\epsilon,t_{0})\}$
and the right punctured neighborhood set
$\mathcal{S}_{right}=\{t|t\in(t_{0},t_{0}+\epsilon)\}$
, at least one of the two sets contains all real numbers corresponding to timestamp instants at which the event is in the same status as at timestamp $t_{0}$.
According to Property 1, the following situation does not exist:
“The event is in one of the 3 execution status at timestamp $t_{0}$, and there exists a real number $\epsilon>0$ such that the event is in other status at all time instants corresponding to the real numbers in the set $\{t|t\in(t_{0}-\epsilon,t_{0})\cup(t_{0},t_{0}+\epsilon)\}$.”
To indicate that an Event instance event is in a certain status during a certain time interval, for bounded real numbers, we use the form of “left-closed and right-open intervals”. For the special cases of -$\infty$ and $\infty$, we use open intervals. For example:
An Event instance $\textsf{event}_{i}$ is in the unstarted status in the interval $(-\infty,t_{1})$, in the running status in the interval $t_{1},t_{2})$, and in the finished status in the interval $t_{2},+\infty)$. This indicates that $\textsf{event}_{i}$ undergoes status transitions at timestamps $t_{1}$ and $t_{2}$.
This paper only focuses on Event instances with bounded starting timestamp and ending timestamp. That is, for the events we consider, it is impossible for them to be in the running status within the following 3 types of time intervals:
$(-\infty,t_{\alpha}),\ [t_{\beta},+\infty),\ (-\infty,+\infty),$
$t_{\alpha}\neq+\infty\wedge t_{\beta}\neq$-$\infty$
Definition 1 (Starting and ending timestamp of an event).
Given an Event instance named event.
If the status of event at the timestamp $t_{0}$ is running, and for all real numbers $t$ in the set $\{t|t<t_{0}\}$, the status of the event at the corresponding moment is unstarted, then $t_{0}$ is defined as the “starting timestamp” of event. We denote it as $\mathrm{TS}$. For an instance of $\mathrm{TS}$, we use $\mathrm{ts}$” to represent it.
If for all real numbers $t$ in the set $\{t|t\geq t_{0}^{{}^{\prime}}\}$, the status of the event event at the corresponding moment is finished, and there exists a real number $\epsilon>0$ such that for all real numbers $t$ in the set $\{t|t\in(t_{0}^{{}^{\prime}}-\epsilon,t_{0}^{{}^{\prime}})\}$, the status of event at the corresponding moment is running, then $t_{0}^{{}^{\prime}}$ is defined as the “ending timestamp” of event. We denote it as $\mathrm{TE}$. For an instance of $\mathrm{TE}$, we use $\mathrm{te}$” to represent it.
Definition 2 (The interval 2-tuple of an event).
Given an Event instance named event, $\mathrm{ts}$ is the starting timestamp of event, $\mathrm{te}$ is the ending timestamp of event. Then, $\mathrm{ts},\mathrm{te})$ is called “The interval of event”, the 2-tuple $(\mathrm{ts},\mathrm{te})$ is called “the interval 2-tuple of event”.
The naming methods for the noun definitions and instances in Definition 1 will be used throughout this paper. A noun definition represents what a class of instances is. “What it is” is an abstract level question, while an instance is a specific object created using the noun definition. Therefore, instances are usually used in specific practical actions. For example, consider the following expression:
There is an Event instance named event, which $\mathrm{TS}$ instance is $\mathrm{ts}$.
In this sentence, Event and $\mathrm{TS}$ are both nouns defined earlier, they are abstract. In the field of computer object-oriented programming, they are called classes. While event and $\mathrm{ts}$ are instances, which are concrete. In the field of computer object-oriented programming, they are called objects.
The following are several properties of Event that we consider.
Property 2 (Uniqueness of starting timestamp and ending timestamp).
An Event instance event has and must have exactly one starting timestamp and one ending timestamp. There cannot be multiple starting timestamps or ending timestamps for event.
Property 3 (Ordering of starting timestamp and ending timestamp).
The starting timestamp of an Event instance must be less than its ending timestamp.
Property 4 (Stability of event execution status).
An Event instance is in the running status at every moment within the interval $\mathrm{ts},\mathrm{te})$; It is in the unstarted status at every moment within the interval $(-\infty,\mathrm{ts})$; It is in the finished status at every moment within the interval $\mathrm{te},+\infty)$.
2.2 Atomic events and composite events
Multiple Event instances may be combined into one Event instance. However, in many cases, we analyze Event instances under the assumption that they cannot be split.
Atomic events and composite events are fundamental concepts in the fields of probability theory, mathematical statistics, and computer science. Due to the extensive application of probability theory and statistics in physics and many engineering disciplines, they also appear in many other areas.
The atomicity and compositionality of an event are defined based on the performance of its quantifiable indicators during execution. If the selected quantifiable indicator remains unchanged during the execution of an event, this event is defined as an atomic event based on that indicator. Conversely, if the indicator changes, this event is a composite event. For a composite event, it can be decomposed according to each stable value of a certain indicator during its change process, with each value corresponding to an atomic event.
In this paper, the type for atomic Event instances is denoted as AtomEvent, and an instance of it is denoted as atomEvent. The type for composite Event instances is denoted as CompEvent, and an instance of it is denoted as compEvent.
And an CompEvent instance compEvent can be expressed as following:
$\textsf{comp}\textsf{Event}=\{\textsf{event}_{1},\textsf{event}_{2},\textsf{event}_{3},\ldots,\textsf{event}_{n}\}$
Definition 3 (Composite event’s starting timestamp and ending timestamp).
Given an CompEvent instance compEvent, with form
$\textsf{comp}\textsf{Event}=\{\textsf{event}_{1},\textsf{event}_{2},\textsf{event}_{3},\ldots,\textsf{event}_{n}\}$
we define the minimum starting timestamp of all its sub-events as the $\mathrm{TS}$ instance of compEvent
$\mathrm{ts}_{\textsf{comp}\textsf{Event}}=\min\{\mathrm{ts}_{\textsf{event}_{i}}|i\in\{1,2,\cdots,n\}\}$
and the maximum ending timestamp of all its sub-events as the $\mathrm{TE}$ instance of compEvent
$\mathrm{te}_{\textsf{comp}\textsf{Event}}=\max\{\mathrm{te}_{\textsf{event}_{i}}|i\in\{1,2,\cdots,n\}\}$
2.3 Event with undetermined interval
In this subsection, we introduce the first abstraction of real-world events in this paper, the concept of “Event with Undetermined Interval” (denoted as $\mathrm{Event}^{*}$). This abstraction is achieved by replacing the fixed time interval of an event with undetermined variables, and thereby creating a bijective mapping between the abstract entity back and the original event. While preserving other attributes of the event (such as subject and space) this approach variabilizes the temporal property, enabling flexible assignment or combination of different time intervals in subsequent analysis and operations. Using an example of “A doctor submitting a paper”, we illustrate how a real-world event can be abstracted into an $\mathrm{Event}^{*}$ instance and expressed mathematically, laying the groundwork for the concepts in later chapters.
Definition 4 (Event with undetermined interval).
An abstract entity derived from an Event instance event by specifying the event’s interval in the form of an undetermined variable is defined as an “Event with Undetermined Interval”, denoted as $\mathrm{Event}^{*}$”. And there exists a bijective mapping between $\mathrm{Event}^{*}$ and Event, denoted as
$\mathrm{Event}^{*}\leftrightarrow\textsf{Event}$
Event is typically abstracted to include properties such as subject, space, etc. In particular, it should be emphasized that an event has a clear and fixed interval Definition 4 precisely focuses on variabilizing the property of “interval”. In many cases, we can provide multiple variables for it to choose from.
For example, suppose there is an event
$\qquad A\ doctor\ submitted\ a\ paper\ to\ a\ journal\ at\ a\ university\ in\ Beijing$
[…]
which denoted as $\textsf{event}_{Dr}$, the interval of this event is “December 1, 2025, between 10:00 a.m. and 12:00 a.m.”, donated as
[ 1764583200, 1764590400)[\ 1764583200,\ 1764590400\ )
If we set this interval as “to be determined” while keeping all other properties unchanged, the expression thus becomes as following
$\qquad A\ doctor\ submitted\ a\ paper\ to\ a\ journal\ at\ a\ university\ in\ Beijing$
$on\ December1,2025,\ between\ [\ ts,\ te\ ).$
Using definition 4, we can abstract this expression as an $\mathrm{Event}^{*}$ instance $\mathrm{event}^{*}_{Dr}$, and
$\mathrm{event}^{*}_{Dr}\leftrightarrow\textsf{event}_{Dr}$
Definition 5 (Atomic and composite $\mathrm{Event}^{*}$).
An $\mathrm{Event}^{*}$ instance is an atomic $\mathrm{Event}^{*}$ instance if it bijectively maps to an AtomEvent instance, denoted as $\textsf{A}\mathrm{Event}^{*}$”; while an $\mathrm{Event}^{*}$ instance is a composite $\mathrm{Event}^{*}$ instance if it bijectively maps to a CompEvent instance, denoted as $\textsf{C}\mathrm{Event}^{*}$”.
3 Optional intervals event and analysis
Before an event occurs, is it prearranged? Does the planner fully grasp every moment of its execution? What pre-execution planning is required for such arranged events, and can we extract and mathematically formalize key elements from these plans? This section explores these questions and introduces novel event-related concepts.
Section 3.1 formally defines Optional Intervals Event ($\mathrm{OIE}$) as a 4-tuple ($\mathcal{C},\mathcal{F},\mathcal{I},\mathcal{A}$) that maps to real-world events and stores feasible execution intervals and sub-event relationships, along with its properties and classifications ($\textsf{Atom}\mathrm{OIE}$$\textsf{Comp}\mathrm{OIE}$).
Section 3.2 introduces $\mathrm{OIE}\mathcal{S}$(a set of OIE instances) and focuses on the set of feasible interval 2-tuple sets derived from its members.
Section 3.3 defines the void OIE instance ($\mathrm{OIE}_{\mathrm{void}}$) which cannot map to any real-world event and serves as the result of infeasible sequential operations.
To clarify the description of complex mathematical types involving 2-tuples, the following conventions are adopted (“Type Name”: abstract category with capitalized first letter; “Type Instance”: specific object with lowercase first letter):
Type Name Type Instance Core Meaning
$\mathrm{2Tuple}$ $\mathrm{2tuple}$ 2-tuple (basic core concept, a single 2-tuple)
$\mathrm{2Tuple}\mathcal{S}$ $\mathrm{2tuple}\mathcal{S}$ Set of 2-tuples
$\mathrm{2Tuple}\mathcal{SS}$ $\mathrm{2tuple}\mathcal{SS}$ Set of sets of 2-tuples
$\mathrm{2Tuple}\mathcal{T}$ $\mathrm{2tuple}\mathcal{T}$ Tuple of 2-tuples
$\mathrm{2Tuple}\mathcal{TS}$ $\mathrm{2tuple}\mathcal{TS}$ Set of tuples of 2-tuples
3.1 Optional intervals event
Now, building on the concept of Event, we propose an important new abstract concept, a 4-tuple. Each of its proper instances can be mapped to an Event instance, while storing the constituent events and all optional intervals of the Event instance it maps to, and can guiding the start and end timestamps of the corresponding Event instance.
3.1.1 Definition
Definition 6 (Optional intervals event).
A class of 4-tuples that map to Event instances and store their available interval information, termed “Optional Intervals Event” (abbreviated as $\mathrm{OIE}$), with the corresponding instance written as $\mathrm{oie}$”.
For an $\mathrm{OIE}$ instance $\mathrm{oie}$, if there exists an Event instance event which forms a mapping from $\mathrm{oie}$ to event, the mapping is express as:
$\mathrm{oie}\rightarrow\textsf{event}$
Its structure is defined as:
$\quad{{(\mathcal{C},\mathcal{F},\mathcal{I},\mathcal{A})}}$
$\mathcal{C}$(1st element): A tuple of other $\mathrm{OIE}$ instances as components, which constitute the current instance.
$\mathcal{F}$(2nd element): A set of tuples of 2-tuples, which stores the details of the feasible interval combinations of the Event instances mapped by the elements of $\mathcal{C}$
$\mathcal{I}$(3rd element): A set of 2-tuples, which stores the overall intervals which can be derived via $\mathcal{F}$
$\mathcal{A}$(4th element): A set of $\textsf{A}\mathrm{Event}^{*}$ instances, which map all atomic events of the real-world event forming the mapping corresponding to the current $\mathrm{OIE}$ instance.
Since the “Optional Intervals Event($\mathrm{OIE}$)” is an entirely novel concept proposed in this paper, we briefly elaborate on it here.
$\mathrm{OIE}$ serves as an abstract planning tool whose core function is to pre-plan all potential execution intervals for a real-world Event instance. Simply put, it can be analogized to a schedule planner that stores all feasible working time windows in advance before the event’s actual execution. Specifically, this planner is structured as a 4-tuple, consisting of four core parts:
1. A bill of immediate components.
2. A detailed schedule.
3. An overall schedule.
4. All practical tasks guided by the atomic components.
The 1st element ($\mathcal{C}$) of an $\mathrm{OIE}$ instance is analogous to “a bill of components”, specifying the subitems that compose the current instance. Its type is a tuple of $\mathrm{OIE}$ instances, which can be an empty tuple “()()”. Some formal examples of $\mathcal{C}$ are provided as follows:
$\mathcal{C}_{\mathrm{oie}_{A}}=(\mathrm{oie}_{1},\mathrm{oie}_{2})$
$\mathcal{C}_{\mathrm{oie}_{B}}=(\mathrm{oie}_{2},\mathrm{oie}_{3},\mathrm{oie}_{1})$
$\mathcal{C}_{\mathrm{oie}_{1}}=\mathcal{C}_{\mathrm{oie}_{2}}=\mathcal{C}_{\mathrm{oie}_{3}}=()$
The $\mathcal{C}$ of $\mathrm{oie}_{1}$, $\mathrm{oie}_{2}$ and $\mathrm{oie}_{3}$ is “()”, because $\mathcal{C}$ is merely a component, while these 3 $\mathrm{OIE}$ instances map atomic Event instances, and atomic events cannot be split into subcomponents. In other words, there is no more small components for these $\mathrm{OIE}$ instances, so their 1st elements are “()”.
The 2nd element ($\mathcal{F}$) of an $\mathrm{OIE}$ instance is analogous to “a detailed schedule set” whose cardinality corresponds to the number of planned schemes. It records the feasible execution interval combinations of each component for every scheme. It is a $\mathrm{2Tuple}\mathcal{TS}$ instance (a set of tuples of 2-tuples), which can also be $\emptyset$. Its form is as follows:
$\mathcal{F}_{\mathrm{oie}}=\{$
$\qquad\left(\ (\mathrm{ts}_{11},\ \mathrm{te}_{11}),\ (\mathrm{ts}_{12},\ \mathrm{te}_{12}),\ \cdots,\ (\mathrm{ts}_{1n},\ \mathrm{te}_{1n})\ \right),$
$\qquad\left(\ (\mathrm{ts}_{21},\ \mathrm{te}_{21}),\ (\mathrm{ts}_{22},\ \mathrm{te}_{22}),\ \cdots,\ (\mathrm{ts}_{2n},\ \mathrm{te}_{2n})\ \right),$
$\qquad\cdots,$
$\qquad\left(\ (\mathrm{ts}_{\lambda{1}},\ \mathrm{te}_{\lambda{1}}),\ (\mathrm{ts}_{\lambda{2}},\ \mathrm{te}_{\lambda{2}}),\ \cdots,\ (\mathrm{ts}_{\lambda{n}},\ \mathrm{te}_{\lambda{n}})\ \right)$
$(\mathrm{ts}_{11},\mathrm{te}_{11})$ is an available interval of the 1st component $\mathrm{OIE}$ instance in $\mathcal{C}_{\mathrm{oie}}$ under schedule plan 1, and $(\mathrm{ts}_{21},\mathrm{te}_{21})$ is an available interval of the 1st component $\mathrm{OIE}$ instance in $\mathcal{C}_{\mathrm{oie}}$ under schedule plan 2. Continuing this traversal to the end, $(\mathrm{ts}_{\lambda{n}},\mathrm{te}_{\lambda{n}})$ is an available interval of the $n$th component $\mathrm{OIE}$ instance in $\mathcal{C}_{\mathrm{oie}}$ under schedule plan $\lambda$, and it satisfies $\mathrm{ts}_{\lambda{n}}<\mathrm{te}_{\lambda{n}}$. The number of such plans equals the cardinality of $\mathcal{F}_{\mathrm{oie}}$, i.e., $\left|{\mathcal{F}_{\mathrm{oie}}}\right|$.
The 3rd item ($\mathcal{I}$) of an $\mathrm{OIE}$ instance is analogous to “an overall schedule”, it is formed by extracting the final start and end timestamps of each scheme to summarize its execution interval, and then integrating the execution intervals of all schemes into one set. It is a $\mathrm{2Tuple}\mathcal{S}$ instance(a set of 2-tuples), which can also be $\emptyset$. When $\mathcal{I}_{\mathrm{oie}}\neq\emptyset$, its form is as follows:
$\mathcal{I}_{\mathrm{oie}}=\{\ (\mathrm{ts}_{{1}},\ \mathrm{te}_{{1}}),\ (\mathrm{ts}_{{2}},\ \mathrm{te}_{{2}}),\ \cdots,\ (\mathrm{ts}_{\Omega},\ \mathrm{te}_{\Omega})\ \},$
$\land$ $\forall(\mathrm{ts}_{i},\mathrm{te}_{i})\in\mathcal{I}_{\mathrm{oie}}:\ \mathrm{ts}_{i}<\mathrm{te}_{i}$
For each element $(\mathrm{ts}_{i},\mathrm{te}_{i})$ in this set, the interval $\mathrm{ts}_{i},\mathrm{te}_{i})$ constructed using the first and second items can serve as the available interval of the Event instance mapped by the $\mathrm{OIE}$ instance. Some formal examples of the 3rd item are as follows:
[…]
$\mathcal{I}_{\mathrm{oie}_{\Omega}}=\emptyset$
1759740600, 1759741200 and 1759741800 are all timestamps, they are respectively 08:50, 09:00 and 09:10 on November 6, 2025 (Beijing Time). For $\mathrm{oie}_{1}$, suppose there is an Event instance $\textsf{event}_{1}$ in the real world that satisfies
$\mathrm{oie}_{1}\rightarrow\textsf{event}_{1}$
Then, $\textsf{event}_{1}$ has two feasible intervals, which are
[1759740600, 1759741200),[1759741200, 1759741800)[1759740600,\ 1759741200),\ [1759741200,\ 1759741800)
Similarly, for $\mathrm{oie}_{2}$, suppose there is an Event instance $\textsf{event}_{2}$ in real world that satisfies
$\mathrm{oie}_{2}\rightarrow\textsf{event}_{2}$
and $\textsf{event}_{2}$ has only one feasible interval, which is
$[1759741200,1759741800)$
$\mathrm{oie}_{\Omega}$ is a special case, whose 3rd element $\mathcal{I}_{\mathrm{oie}_{\Omega}}$ is $\emptyset$. However, if there were an Event instance $\textsf{event}_{\Omega}$ mapped to by $\mathrm{oie}_{\Omega}$, we would have no available interval for it to execute.
The 4th element ($\mathcal{A}$) of an $\mathrm{OIE}$ instance is analogous to “atomic practical tasks”, which are mapped by the atomic components of elements in $\mathcal{C}_{\mathrm{oie}}$. When analyzing the feasibility of a complex event, we typically decompose it—based on real world conditions—into indivisible atomic events. If even one of the atomic events cannot be executed, the entire complex event becomes infeasible. This element serves two main purposes: first, it reinforces the connection between the abstract $\mathrm{OIE}$ instance and the real world Event instance; second, it plays a critical role in the feasibility evaluations involved in the operations discussed in later sections. Some formal examples of the 4th item are as follows:
$\mathcal{A}_{\mathrm{oie}_{1}}=\{\mathrm{event}^{*}_{1}\},$
$\mathcal{A}_{\mathrm{oie}_{2}}=\{\mathrm{event}^{*}_{2}\},$
$\mathcal{A}_{\mathrm{oie}_{\Omega}}=\emptyset$
To more clearly illustrate $\mathrm{OIE}$ and subsequent content, we will present a detailed example that runs through every key concept in this article.
”An academic conference where N doctors submit their papers”
This conference has a submission window, which includes a start timestamp and a deadline(ending timestamp). Suppose it starts at 00:00 on August 1, 2025, and ends at 00:00 on August 2, 2025(This is just an assumption). For simplicity, we use the hour numbers of August 1 to represent the start and end timestamp of each hour. Thus, this window is denoted as
[ 0, 24)[\ 0,\ 24\ )
where “0” refers to 00:00 on August 1, “24” corresponds to 00:00 on August 2, and the right parenthesis after 24 indicates that 00:00 on August 2 is not included.
Suppose there are 3 doctors: $Dr_{A}$, $Dr_{B}$ and $Dr_{C}$. $Dr_{A}$ can submit his paper during either 0:00-1:00 or 21:00-22:00, and choosing either interval requires using the entire duration of that time slot. $Dr_{B}$’s available submission intervals are 0:00-1:00, 13:00-14:00, and 20:00-22:00. $Dr_{C}$’s available intervals, meanwhile, are 0:00-1:00 and 19:00-22:00.
  • $Dr_{A}$: 0:00–1:00, 21:00–22:00
  • $Dr_{B}$: 0:00–1:00, 13:00–14:00, 20:00–22:00
  • $Dr_{C}$: 0:00–1:00, 19:00–22:00
Furthermore, the scenario where $Dr_{A}$ submits a manuscript between 21:00-22:00, $Dr_{B}$ between 20:00-22:00, and $Dr_{C}$ between 19:00-22:00 is impossible, but that involving any two of them is feasible. Similarly, it is impossible for all 3 doctors to submit their manuscripts during the 00:00–01:00 interval, while the submission of any two within this window is feasible. In reality, they may only have one shared computer available around 22:00 or 00:00, making simultaneous operation for 3 doctors unfeasible but feasible for two.
We use 3 Event instances to represent the submission of each doctor
$\textsf{event}_{Dr_{A}},\ \textsf{event}_{Dr_{B}},\ \textsf{event}_{Dr_{C}}$
And we use 4 $\mathrm{OIE}$ instances to map the corresponding events
$\mathrm{oie}_{A}\rightarrow\textsf{event}_{Dr_{A}}$
$\mathrm{oie}_{B}\rightarrow\textsf{event}_{Dr_{B}}$
$\mathrm{oie}_{C}\rightarrow\textsf{event}_{Dr_{C}}$
$\mathrm{oie}_{A_{\theta}}\rightarrow\textsf{event}_{Dr_{A}}$
$\mathrm{oie}_{A_{\theta}}$ is a special $\mathrm{OIE}$ instance that requires $Dr_{A}$ to submit their papers exclusively within the time slot 21:00–22:00.
In this scenario, paper submission constitutes an indivisible atomic event, so the first item of theirs can be expressed as
$\mathcal{C}_{\mathrm{oie}_{A}}=()$
$\mathcal{C}_{\mathrm{oie}_{B}}=()$
$\mathcal{C}_{\mathrm{oie}_{C}}=()$
$\mathcal{C}_{\mathrm{oie}_{A_{\theta}}}=()$
In our analogy, the third items of theirs reflect the final summary of the intervals corresponding to each sub-event. They can be expressed as follows:
[…]
In this example, each doctor’s paper submission does not need to be divided into several smaller events. Therefore, each tuple in the second item only requires one element. They can be expressed as follows:
[…]
The 4th item of each $\mathrm{OIE}$ instance reflects the atomic practical Event instances with undetermined execution time, mapped by each atomic $\mathrm{OIE}$ components. They can be expressed as follows:
$\mathcal{A}_{\mathrm{oie}_{A}}=\{\mathrm{event}^{*}_{Dr_{A}}\}$
$\mathcal{A}_{\mathrm{oie}_{B}}=\{\mathrm{event}^{*}_{Dr_{B}}\}$
$\mathcal{A}_{\mathrm{oie}_{C}}=\{\mathrm{event}^{*}_{Dr_{C}}\}$
$\mathcal{A}_{\mathrm{oie}_{A_{\theta}}}=\{\mathrm{event}^{*}_{Dr_{A}}\}$
Finally, the forms of these 4 $\mathrm{OIE}$ instances are as follows:
$\mathrm{oie}_{A}={{(}}$
$\quad(),$
[…]
$\quad\{\ \mathrm{event}^{*}_{Dr_{A}}\ \}$
$\mathrm{oie}_{B}={{(}}$
$\quad(),$
[…]
$\quad\{\ \mathrm{event}^{*}_{Dr_{B}}\ \}$
$\mathrm{oie}_{C}={{(}}$
$\quad(),$
[…]
$\quad\{\ \mathrm{event}^{*}_{Dr_{C}}\ \}$
$\mathrm{oie}_{A_{\theta}}={{(}}$
$\quad(),$
[…]
$\quad\{\ \mathrm{event}^{*}_{Dr_{A}}\ \}$
3.1.2 Properties
In this subsubsection, we will discuss the properties among the 4 elements of the $\mathrm{OIE}$, as well as the different types of $\mathrm{OIE}$.
As defined in our definition of the $\mathrm{OIE}$(Definition 6), the 3rd element is a relatively intuitive element to understand, it is a $\mathrm{2Tuple}\mathcal{S}$ instance(a set of 2-tuples). And from an application perspective, it represents the collection of all feasible intervals corresponding to the mapped Event instance. In contrast, the 2nd element is a somewhat a bit more complex set, so we first discuss the relationship between the 2nd and 3rd elements of $\mathrm{OIE}$.
We need a helper function to calculate a $\mathrm{2Tuple}$ instance(a 2-tuple) from a $\mathrm{2Tuple}\mathcal{T}$ instance as a feature value.
Definition 7 (The bound 2-tuple of a finite non-empty $\mathrm{2Tuple}\mathcal{T}$ instance & helper function & permutation invariance).
1) definition:
Let $\mathrm{2tuple}\mathcal{T}_{\theta}$ be a finite-length non-empty $\mathrm{2Tuple}\mathcal{T}$ instance in the ordered enumeration form
$\mathrm{2tuple}\mathcal{T}_{\theta}=(\ (x_{1},\ y_{1}),\ (x_{2},\ y_{2}),\ \dots,\ (x_{n},\ y_{n})\ )$
Let
$L=min\{x_{i}|i\in\{1,2,\cdots,n\}\}$
$R=max\{y_{i}|i\in\{1,2,\cdots,n\}\}$
then, the 2-tuple $(L,R)$ is called “The bound 2-tuple of a non-empty $\mathrm{2Tuple}\mathcal{T}$ instance with limited length”, denoted as $\mathrm{Bound2Tuple}$”.
2) helper function:
$\mathfrak{f}\mathrm{Bound2tuple}(\mathfrak{p}\mathrm{2tuple}\mathcal{T})\rightarrow$
$\qquad\text{The }\ \mathrm{Bound2Tuple}\text{ instance of }\mathfrak{p}\mathrm{2tuple}\mathcal{T}$
$\mathfrak{p}\mathrm{2tuple}\mathcal{T}$ A finite-length non-empty $\mathrm{2Tuple}\mathcal{T}$ instance
Return The bound 2-tuple of $\mathfrak{p}\mathrm{2tuple}\mathcal{T}$
3) permutation invariance:
Suppose there are two finite-length non-empty $\mathrm{2Tuple}\mathcal{T}$ instances $\mathrm{2tuple}\mathcal{T}$ and $\mathrm{2tuple}\mathcal{T}^{{}^{\prime}}$ with the same length $n>0$, and they satisfy permutation equivalence via a permutation matrix $\mathcal{M}$
$\mathrm{2tuple}\mathcal{T}\stackrel{{\scriptstyle\mathcal{M}}}{{\sim}}\mathrm{2tuple}\mathcal{T}^{{}^{\prime}}$
Then the following holds
$\mathfrak{f}\mathrm{Bound2tuple}(\mathrm{2tuple}\mathcal{T})=\mathfrak{f}\mathrm{Bound2tuple}(\mathrm{2tuple}\mathcal{T}^{{}^{\prime}})$
Proof of the permutation invariance in Definition 7.
Let
$\mathrm{2tuple}\mathcal{T}=(\ (x_{1},\ y_{1}),\ (x_{2},\ y_{2}),\ \cdots,\ (x_{n},\ y_{n})\ ),$
$\mathrm{2tuple}\mathcal{T}^{{}^{\prime}}=(\ (x_{1}^{{}^{\prime}},\ y_{1}^{{}^{\prime}}),\ (x_{2}^{{}^{\prime}},\ y_{2}^{{}^{\prime}}),\cdots,\ (x_{n}^{{}^{\prime}},\ y_{n}^{{}^{\prime}})\ ))$
be two non-empty $\mathrm{2Tuple}\mathcal{T}$ instances with finite length $n>0$ that satisfy permutation equivalence via a permutation matrix $\mathcal{M}$
$\mathrm{2tuple}\mathcal{T}\stackrel{{\scriptstyle\mathcal{M}}}{{\sim}}\mathrm{2tuple}\mathcal{T}^{{}^{\prime}}$
This means $\mathrm{2tuple}\mathcal{T}$ and $\mathrm{2tuple}\mathcal{T}^{{}^{\prime}}$ have same elements, i.e.:
$\{(x_{1},y_{1}),\ (x_{2},y_{2}),\ \cdots,\ (x_{n},y_{n})\}=\{(x_{1}^{{}^{\prime}},y_{1}^{{}^{\prime}}),(x_{2}^{{}^{\prime}},y_{2}^{{}^{\prime}}),\cdots,(x_{n}^{{}^{\prime}},y_{n}^{{}^{\prime}})\}$
Consequently, the set of all first items is preserved:
$\{x_{1},x_{2},\cdots,x_{n}\}=\{x_{1}^{{}^{\prime}},x_{2}^{{}^{\prime}},\cdots,x_{n}^{{}^{\prime}}\}$
And the set of all second items is preserved:
$\{y_{1},y_{2},\cdots,y_{n}\}=\{y_{1}^{{}^{\prime}},y_{2}^{{}^{\prime}},\cdots,y_{n}^{{}^{\prime}}\}$
Since the bound 2-tuple (L,R)(L,\ R) is defined as:
$L=min\{x_{i}|i\in\{1,2,\cdots,n\}\}$
$R=max\{y_{i}|i\in\{1,2,\cdots,n\}\}$
then
$min\{x_{i}^{{}^{\prime}}|i\in\{1,2,\cdots,n\}\}=min\{x_{i}|i\in\{1,2,\cdots,n\}\}=L$
$max\{y_{i}^{{}^{\prime}}|i\in\{1,2,\cdots,n\}\}=max\{y_{i}|i\in\{1,2,\cdots,n\}\}=R$
i.e.:
$\mathfrak{f}\mathrm{Bound2tuple}(\mathrm{2tuple}\mathcal{T})=\mathfrak{f}\mathrm{Bound2tuple}(\mathrm{2tuple}\mathcal{T}^{{}^{\prime}})$
proof completed. ∎
Using the helper function of Definition 7, We can present the relationship between the $\mathcal{F}$(2nd element) and $\mathcal{I}$(3rd element) of $\mathrm{OIE}$ instance, and provide the property that derives the $\mathcal{I}$(3rd element) using the $\mathcal{F}$(2nd element).
Property 5 (The relationship between the $\mathcal{F}$(2nd element) and $\mathcal{I}$(3rd element) of $\mathrm{OIE}$).
Suppose there is an $\mathrm{OIE}$ instance $\mathrm{oie}$, with form
$\mathrm{oie}={{(}}\mathcal{C}_{\mathrm{oie}},\ \mathcal{F}_{\mathrm{oie}},\ \mathcal{I}_{\mathrm{oie}},\ \mathcal{A}_{\mathrm{oie}}{{)}}$
If $\mathcal{F}_{\mathrm{oie}}\neq\emptyset$:
$\mathcal{I}_{\mathrm{oie}}=\bigg\{\ \mathfrak{f}\mathrm{Bound2tuple}(\mathrm{2tuple}\mathcal{T}_{i})\ \bigg|\mathrm{2tuple}\mathcal{T}_{i}\in\mathcal{F}_{\mathrm{oie}}\ \bigg\}$
Else $\mathcal{F}_{\mathrm{oie}}=\emptyset$:
$\mathcal{I}_{\mathrm{oie}}=\emptyset$
We use the scenario of “Three doctors’ submitting papers” to give an example that illustrates Property 5. We construct the following $\mathrm{OIE}$ instance.
$\mathrm{oie}_{{Dr_{B}}\_and\_{Dr_{A}}}={{(}}$
$\quad(\ \mathrm{oie}_{B},\ \mathrm{oie}_{A}\ ),$
$\quad\{$
$\qquad(\ (0,1),\ (0,1)\ ),$
[…]
$\quad\},$
[…]
$\quad\{\ \mathrm{event}^{*}_{Dr_{B}},\ \mathrm{event}^{*}_{Dr_{A}}\ \}$
Analyze this example from the literal and practical perspectives.
The practical meaning of $\mathrm{oie}_{{Dr_{B}}\_and\_{Dr_{A}}}$ is $Dr_{B}$ and $Dr_{A}$ submit papers”. This sentence is a textual description of an event, which involves two participants. Each participant, as a subject, leads an event—specifically, the respective paper submissions of the two doctors.
$\mathcal{F}_{\mathrm{oie}_{{Dr_{B}}\_and\_{Dr_{A}}}}$(the 2nd element of $\mathrm{oie}_{{Dr_{B}}\_and\_{Dr_{A}}}$):
$\quad(\ (0,1),\ (0,1)\ ),$
[…]
it means
[…]
$01:00.\displaystyle\textsf{event}_{Dr_{B}}\ runs13:0-4:0,\ while\ \textsf{event}_{Dr_{A}}\ runs00:0$-1:0.
[…]
$2:00.\displaystyle\textsf{event}_{Dr_{B}}\ runs20:0-2:0,\ while\ \textsf{event}_{Dr_{A}}\ runs21:0-2:0$.
The above are feasible details of plans. In the real world, one of these plans will be selected to occur.
$\mathcal{I}_{\mathrm{oie}_{{Dr_{B}}\_and\_{Dr_{A}}}}$(the 3rd element of $\mathrm{oie}_{{Dr_{B}}\_and\_{Dr_{A}}}$):
$\{\ (0,\ 1),\ (0,\ 22),\ (0,\ 14),\ (13,\ 22),\ (20,\ 22)\ \}$
it means
$00.\displaystyle The\ entire\ event,\ which\ consists\ of\ \textsf{event}_{Dr_{B}}\ and\ \textsf{event}_{Dr_{A}},\ runs00:0-1:$0.
$2:00.\displaystyle The\ entire\ event,\ which\ consists\ of\ \textsf{event}_{Dr_{B}}\ and\ \textsf{event}_{Dr_{A}},\ runs00$:0-2:0.
[…]
$2:00.\displaystyle The\ entire\ event,\ which\ consists\ of\ \textsf{event}_{Dr_{B}}\ and\ \textsf{event}_{Dr_{A}},\ runs13$:0-2:0.
$2:00.\displaystyle The\ entire\ event,\ which\ consists\ of\ \textsf{event}_{Dr_{B}}\ and\ \textsf{event}_{Dr_{A}},\ runs20$:0-2:0.
Further analyze this example from the dimensions of formalization and internal logic. $\mathrm{oie}_{{Dr_{B}}\_and\_{Dr_{A}}}$ maps this real-world event
$\mathrm{oie}_{{Dr_{B}}\_and\_{Dr_{A}}}\rightarrow$
${}^{\prime\prime}The\ real-world\ event\ that\ consists\ of\ {Dr_{B}}^{\prime}s\ and\ {Dr_{A}}^{\prime}s\ submitting\ papers^{\prime\prime}$
The 1st element of this $\mathrm{OIE}$ instance $(\mathrm{oie}_{B},\mathrm{oie}_{A})$ expresses $Dr_{B}$ and $Dr_{A}$” in the declarative sentence.
Using Definition 7, we process each item of the 2nd element of $\mathrm{oie}_{{Dr_{B}}\_and\_{Dr_{A}}}$ to implement the logic process of Property 5.
$\mathfrak{f}\mathrm{Bound2tuple}(\ ((0,1),\ (0,1))\ )\rightarrow(0,1)$
[…]
Put all the result 2-tuples on the right into a single set, the result is
$\{\ (0,\ 1),\ (0,\ 22),\ (0,\ 14),\ (13,\ 22),\ (20,\ 22)\ \}$
This is precisely the $\mathcal{I}$(the 3rd item) of $\mathrm{oie}_{{Dr_{B}}\_and\_{Dr_{A}}}$.
In fact, a careful observation reveals that $\mathrm{oie}_{{Dr_{B}}\_and\_{Dr_{A}}}$ is constructed from $\mathrm{oie}_{Dr_{B}}$ and $\mathrm{oie}_{Dr_{A}}$ through certain methods, analogous to how the real-world event “Dr. A and Dr. B submit papers” is composed of the two individual events “Dr. A submits a paper” and “Dr. B submits a paper”. Both our mathematical description and people’s verbal expression elaborate on this from a formalization perspective. How $\mathrm{oie}_{{Dr_{B}}\_and\_{Dr_{A}}}$ is constructed from $\mathrm{oie}_{Dr_{B}}$ and $\mathrm{oie}_{Dr_{A}}$ will be elaborated in detail in the next chapter, which focuses on operations. In this chapter, we concentrate on the $\mathrm{OIE}$ itself.
Here we present other two $\mathrm{OIE}$ instances as examples
$\mathrm{oie}_{{Dr_{A}}\_and\_{Dr_{B}}}={{(}}$
$\quad(\ \mathrm{oie}_{A},\ \mathrm{oie}_{B}\ ),$
$\quad\{$
$\qquad(\ (0,1),\ (0,1)\ ),$
[…]
$\quad\},$
[…]
$\quad\{\ \mathrm{event}^{*}_{Dr_{A}},\ \mathrm{event}^{*}_{Dr_{B}}\ \}$
and
$\mathrm{oie}_{void}={{(}}\ (),\ \emptyset,\ \emptyset,\ \emptyset\ {{)}}$
They both satisfy Property 5 when subjected to the same verification method.
Next, we specify the equality condition for two $\mathrm{OIE}$ instances.
Definition 8 (Equality of $\mathrm{OIE}$ instances).
Suppose $\mathrm{oie}_{1}$ and $\mathrm{oie}_{2}$ are $\mathrm{OIE}$ instances, with forms
$\mathrm{oie}_{1}={{(\mathcal{C}_{1},\mathcal{F}_{1},\mathcal{I}_{1},\mathcal{A}_{1})}},$
$\mathrm{oie}_{2}={{(\mathcal{C}_{2},\mathcal{F}_{2},\mathcal{I}_{2},\mathcal{A}_{2})}}$
If
$\mathcal{C}_{1}=\mathcal{C}_{2}\ \land\ \mathcal{F}_{1}=\mathcal{F}_{2}\ \land\ \mathcal{I}_{1}=\mathcal{I}_{2}\ \land\ \mathcal{A}_{1}=\mathcal{A}_{2}$
Then $\mathrm{oie}_{1}$ and $\mathrm{oie}_{2}$ are equal, denoted as
$\mathrm{oie}_{1}=\mathrm{oie}_{2}$
Else if
$\mathcal{C}_{1}\neq\mathcal{C}_{2}\ \lor\ \mathcal{F}_{1}\neq\mathcal{F}_{2}\ \lor\ \mathcal{I}_{1}\neq\mathcal{I}_{2}\ \lor\ \mathcal{A}_{1}\neq\mathcal{A}_{2}$
they are not equal, denoted as
$\mathrm{oie}_{1}\neq\mathrm{oie}_{2}$
3.1.3 $\textsf{Atom}\mathrm{OIE}$ and $\textsf{Comp}\mathrm{OIE}$
$\mathrm{OIE}$ can also have the type of being either atomic or composite. In subsection 2.2, we have discussed atomic events($\textsf{Atom}\mathrm{OIE}$) and composite events($\textsf{Comp}\mathrm{OIE}$). In this subsubsection, we discuss the atomic and composite classification of optional intervals events.
Definition 9 (Atomic $\mathrm{OIE}$).
$\mathrm{OIE}$ instances that map to AtomEvent instances are called “atomic $\mathrm{OIE}$” and are denoted as $\textsf{Atom}\mathrm{OIE}$”. An instance is written as $\textsf{atom}\mathrm{Oie}$”.
Definition 10 (Composite $\mathrm{OIE}$).
$\mathrm{OIE}$ instances that map to CompEvent instances are called “composite $\mathrm{OIE}$” and are denoted as $\textsf{Comp}\mathrm{OIE}$”. An instance is written as $\textsf{comp}\mathrm{Oie}$”.
$\textsf{Atom}\mathrm{OIE}$ is also a type of $\mathrm{OIE}$, their instances cannot be further split due to its responding real world event. Similarly, $\textsf{Comp}\mathrm{OIE}$ is also a type of $\mathrm{OIE}$, and their instances can be further split.
In the example “Submitting papers of 3 doctors”
$\mathrm{oie}_{A}={{(}}$
$\quad(),$
[…]
$\quad\{\ \mathrm{event}^{*}_{Dr_{A}}\ \}$
$\mathrm{oie}_{B}={{(}}$
$\quad(\mathrm{oie}_{B}),$
[…]
$\quad\{\ \mathrm{event}^{*}_{Dr_{B}}\ \}$
$\mathrm{oie}_{C}={{(}}$
$\quad(\mathrm{oie}_{C}),$
[…]
$\quad\{\ \mathrm{event}^{*}_{Dr_{C}}\ \}$
all are $\textsf{Atom}\mathrm{OIE}$ instance.
And the $\mathrm{OIE}$ instance
$\mathrm{oie}_{{Dr_{B}}\_and\_{Dr_{A}}}={{(}}$
$\quad(\ \mathrm{oie}_{B},\ \mathrm{oie}_{A}\ ),$
$\quad\{$
$\qquad(\ (0,1),\ (0,1)\ ),$
[…]
$\quad\},$
[…]
$\quad\{\ \mathrm{event}^{*}_{Dr_{B}},\ \mathrm{event}^{*}_{Dr_{A}}\ \}$
is an $\textsf{Comp}\mathrm{OIE}$ instance.
$\textsf{Atom}\mathrm{OIE}$ and $\textsf{Comp}\mathrm{OIE}$ have the following properties.
Property 6 (The property of $\textsf{Atom}\mathrm{OIE}$).
Suppose there is an $\textsf{Atom}\mathrm{OIE}$ instance $\textsf{atom}\mathrm{Oie}$, it satisfies:
$\dim{\left(\mathcal{C}_{\textsf{atom}\mathrm{Oie}}\right)}=0$
$\land$ $\forall\mathrm{2tuple}\mathcal{T}_{i}\in\mathcal{F}_{\textsf{atom}\mathrm{Oie}}:\dim{\left(\mathrm{2tuple}\mathcal{T}_{i}\right)}=1$
Property 7 (The property of $\textsf{Comp}\mathrm{OIE}$).
Suppose there is an $\textsf{Comp}\mathrm{OIE}$ instance $\textsf{comp}\mathrm{Oie}$, it satisfies:
$\dim{\left(\mathcal{C}_{\textsf{comp}\mathrm{Oie}}\right)}>0$
$\land$ $\forall\mathrm{2tuple}\mathcal{T}_{i}\in\mathcal{F}_{\textsf{comp}\mathrm{Oie}}:$
$\qquad\dim{\left(\mathrm{2tuple}\mathcal{T}_{i}\right)}>1\ \land\ \dim{\left(\mathrm{2tuple}\mathcal{T}_{i}\right)}=\dim{\left(\mathcal{C}_{\textsf{comp}\mathrm{Oie}}\right)}$
3.2 Set of optional intervals events
In the previous subsection, we proposed $\mathrm{OIE}$. As we mentioned when analyzing events earlier, an event may be composed of several events. So in this subsection, we will analyze the properties of multiple $\mathrm{OIE}$ instances.
A set of sets of 2-tuples, that is, a set of $\mathrm{2Tuple}\mathcal{S}$ instances, is denoted as $\mathrm{2Tuple}\mathcal{SS}$, and an instance is denoted as $\mathrm{2tuple}\mathcal{SS}$, its ordered enumeration form is as follows:
$\mathrm{2tuple}\mathcal{SS}=\{\mathrm{2tuple}\mathcal{S}_{1},\mathrm{2tuple}\mathcal{S}_{2},\cdots,\mathrm{2tuple}\mathcal{S}_{n}\}$
For example,
[…]
In fact, this refers to the set formed by combining the third items of the 3 $\mathrm{OIE}$ instances ($\mathrm{oie}_{A}$, $\mathrm{oie}_{B}$ and $\mathrm{oie}_{C}$) in the example of the doctors’ submissions.
Back to the main topic, set composed of $n$ $\mathrm{OIE}$ instances with ordered enumeration form
$\{\mathrm{oie}_{1},\mathrm{oie}_{2},\cdots,\mathrm{oie}_{n}\}$
is called $\mathrm{OIE}\mathcal{S}$, and an instance is denoted as $\mathrm{oie}\mathcal{S}$, and $\mathrm{oie}\mathcal{S}$ can be $\emptyset$.
Here, we need to clarify 2 points.
The first is that, this paper only analyzes finite $\mathrm{OIE}\mathcal{S}$ instances. When the cardinality of an $\mathrm{OIE}\mathcal{S}$ instance is equal to $\infty$, we would have to deal with an infinite number of Event instances. That is an extremely special case, and this paper will not discuss such a situation.
The second is that, Although an $\mathrm{OIE}\mathcal{S}$ instance is denoted by an ordered enumeration for convenience, it represents an unordered set. The order of elements is irrelevant until an index tuple $\mathrm{idx}\mathcal{T}$ is specified to impose a sequence for operations. In the paper, other sets expressed through ordered enumeration forms also follow this notation convention.
Regarding $\mathrm{OIE}\mathcal{S}$ instance, this paper focuses on the set formed by the third items of each element. As two operations in the next chapter will utilize this set, we first provide its formal definition herein.
Definition 11 (Set of $\mathcal{I}$ of finite $\mathrm{OIE}\mathcal{S}$ instance).
For a finite $\mathrm{OIE}\mathcal{S}$ instance $\mathrm{oie}\mathcal{S}$ with the ordered enumeration form
$\mathrm{oie}\mathcal{S}=\{\mathrm{oie}_{1},\mathrm{oie}_{2},\cdots,\mathrm{oie}_{n}\}$
the set ordered enumeration form
$\{\mathcal{I}_{\mathrm{oie}_{1}},\mathcal{I}_{\mathrm{oie}_{2}},\cdots,\mathcal{I}_{\mathrm{oie}_{n}}\}$
is called “The set of $\mathcal{I}$ of $\mathrm{oie}\mathcal{S}$”, abbreviated as $\mathcal{IS}$ of $\mathrm{oie}\mathcal{S}$.
For the special case where an $\mathrm{OIE}\mathcal{S}$ instance is $\emptyset$, its $\mathcal{IS}$ is $\emptyset$.
For non-empty $\mathrm{OIE}\mathcal{S}$ instances, $\mathcal{IS}$ is an instance of the $\mathrm{2Tuple}\mathcal{SS}$ (set of sets of 2-tuples) type.
Take the aforementioned case of the 3 doctors submitting their manuscripts as an example here. We place $\mathrm{oie}_{Dr_{A}}$, $\mathrm{oie}_{Dr_{B}}$ and $\mathrm{oie}_{Dr_{C}}$ into a set to form an $\mathrm{OIE}\mathcal{S}$ instance $\mathrm{oie}\mathcal{S}_{Drs}$, with ordered enumeration form
$\mathrm{oie}\mathcal{S}_{Drs}=\{\mathrm{oie}_{Dr_{A}},\mathrm{oie}_{Dr_{B}},\mathrm{oie}_{Dr_{C}}\}$
Then, its $\mathcal{IS}$ is
$\mathcal{IS}_{\mathrm{oie}\mathcal{S}_{Drs}}$
$\quad=\{\ \mathcal{I}_{\mathrm{oie}_{Dr_{A}}},\ \mathcal{I}_{\mathrm{oie}_{Dr_{B}}},\ \mathcal{I}_{\mathrm{oie}_{Dr_{C}}}\ \}$
$\quad=$
$\qquad\{$
[…]
$\qquad\}$
3.3 Void optional intervals event
In this subsection, we need to supplement a type of $\mathrm{OIE}$ which can’t be mapped to any real-world event.
Definition 12 (Void $\mathrm{OIE}$).
An $\mathrm{OIE}$ instance whose $\mathcal{C}$(1st element) is $()$, $\mathcal{F}$(2nd element) is $\emptyset$, $\mathcal{I}$(3rd element) is $\emptyset$, and $\mathcal{A}$(4th element) is $\emptyset$, is a void $\mathrm{OIE}$ instance. We call its type $\mathrm{OIE}_{\mathrm{void}}$”. $\mathrm{OIE}_{\mathrm{void}}$ is also $\mathrm{OIE}$, and $\mathrm{OIE}_{\mathrm{void}}$ has only one instance, with form
$\mathrm{oie}_{\mathrm{void}}=\left(\ (),\ \emptyset,\ \emptyset,\ \emptyset\ \right)$
This object doesn’t have a subject, doesn’t have an intervals set to be selected and doesn’t have any details too, so it cannot be mapped to any real-world event. In the following context, we will propose the sequential operation of $\mathrm{OIE}$ to plan the execution of real-world events. When the result of the sequential operation is $\mathrm{oie}_{\mathrm{void}}$, it indicates that such a plan cannot be arranged in the real world.
4 Two sequential operations on the elements of a finite set of optional intervals events
In this section, we are going to start elaborating on another core of this paper, the “Sequential Operations”. We will propose two n-ary finitary operations on the set elements of an $\mathrm{OIE}\mathcal{S}$ instance. Through each operation, we will obtain a result that is still of the $\mathrm{OIE}$ type, this result can change the information about the optional intervals of all the $\mathrm{OIE}$ instances participating in the operation. That is, when these $\mathrm{OIE}$ instances participate in operations, their optional interval information may be reshaped. The application significance of this reshaping will be gradually discussed in the subsequent sections of this paper. The sequential operations we propose have strong abstraction capabilities. At the end of the paper, we will briefly describe their specific application prospects in the fields of physics, computer science and applied mathematics. In this section, we will focus on discussing the rules of these operations.
4.1 The set that is naturally isomorphic to the Cartesian product of the sets of interval 2-tuples of each member in a finite set of optional intervals events.
We first carry out the work of using the set indices to represent the operation order of the elements of a set.
Definition 13 (An index tuple of a finite index set).
Given a finite set $\mathrm{I}_{\mathcal{S}}$ with cardinality $n$ ($n\in\mathbb{N}^{*}$)
$\mathrm{I}_{\mathcal{S}}=\{\ 1,\ 2,\ 3,\ \cdots,\ n\ \}$
If a tuple $(i_{1},i_{2},i_{3},\cdots,i_{n})$ satisfies
$\forall\ k\in\mathrm{I}_{\mathcal{S}}:\ i_{k}\in\mathrm{I}_{\mathcal{S}}$
$\land$ $\forall\ m,n\in\mathrm{I}_{\mathcal{S}}(m\neq n):\ i_{m}\neq i_{n}$
Then the tuple $(i_{1},i_{2},i_{3},\cdots,i_{n})$ is an index tuple of the finite index set $\mathrm{I}_{\mathcal{S}}$, denoted as $\mathrm{Idx}\mathcal{T}$” with length $n$.
Property 8 (Permutation equivalence of two $\mathrm{Idx}\mathcal{T}$ instances).
For any two $\mathrm{Idx}\mathcal{T}$ instances with length $n(n\in\mathbb{N}^{+})$,
$\mathrm{idx}\mathcal{T}_{1}=\{\ i_{1},\ i_{2},\ i_{3},\ \cdots,\ i_{n}\ \},$
$\mathrm{idx}\mathcal{T}_{2}=\{\ j_{1},\ j_{2},\ j_{3},\ \cdots,\ j_{n}\ \},$
$\land$ $\mathrm{idx}\mathcal{T}_{1}\neq\mathrm{idx}\mathcal{T}_{2}$
there must be a permutation matrix $\mathcal{M}$ satisfying
$\mathrm{idx}\mathcal{T}_{1}\cdot\mathcal{M}=\mathrm{idx}\mathcal{T}_{2}$
That means for any two $\mathrm{Idx}\mathcal{T}$ instances $\mathrm{idx}\mathcal{T}_{1}$ and $\mathrm{idx}\mathcal{T}_{2}$ with length $n$, they are permutationally equivalent, denoted as
$\mathrm{idx}\mathcal{T}_{1}\stackrel{{\scriptstyle\mathcal{M}}}{{\sim}}\mathrm{idx}\mathcal{T}_{2}$
Here we emphasize that, in this paper, we will use Definition 13 to correspond to all sets in the form of ordered enumeration. Throughout the text of this paper, once the order of elements within an $\mathrm{Idx}\mathcal{T}$ instance can express the iteration order of elements in a set, it indicates that the length of this $\mathrm{Idx}\mathcal{T}$ instance is the same as the cardinality of the set.
Now, we need to design a function to get a set that is naturally isomorphic to the Cartesian product of all the elements in a finite $\mathrm{2Tuple}\mathcal{SS}$(set of sets of 2-tuples) instance, and this function serves as a helper function for refining the sequential operation rules later.
Definition 14 (Function to get the set that is naturally isomorphic to the Cartesian product of all members of a finite $\mathrm{2Tuple}\mathcal{SS}$ instance in an index order).
$\mathfrak{f}\mathrm{NatIso2CP}(\mathfrak{p}\mathrm{2tuple}\mathcal{SS},\ \mathfrak{p}\mathrm{idx}\mathcal{T})=$
$\quad The\ set\ that\ is\ naturally\ isomorphic\ to\ \prod_{i=1}^{n}\mathfrak{p}\mathrm{2tuple}\mathcal{S}_{\mathfrak{p}{idx_{i}}}$
We stipulate that, in the special case when $\mathfrak{p}\mathrm{2tuple}\mathcal{SS}=\emptyset$, the result is also $\emptyset$.
$\mathfrak{p}\mathrm{2tuple}\mathcal{SS}$ A finite set of sets of 2-tuples of size $n$, in the form of $\mathfrak{p}\mathrm{2tuple}\mathcal{SS}=\{\ \mathfrak{p}\mathrm{2tuple}\mathcal{S}_{1},\ \mathfrak{p}\mathrm{2tuple}\mathcal{S}_{2},\ \cdots,\ \mathfrak{p}\mathrm{2tuple}\mathcal{S}_{n}\ \}$
[…]
Return The set that is naturally isomorphic to the Cartesian product of all elements in $\mathfrak{p}\mathrm{2tuple}\mathcal{SS}$, with the order of operands constructed using $\mathfrak{p}\mathrm{idx}\mathcal{T}$ as the index, and the result is a set of tuples of 2-tuples
Here, we need to explain why we use the natural isomorphic of Cartesian products. “Natural isomorphism” is a term in category theory, whose purpose is to reveal the “intrinsicequivalenceintrinsic\ equivalence” between the constructions of different entities.
For Cartesian product operation
$\mathfrak{p}\mathrm{2tuple}\mathcal{S}_{\mathfrak{p}{idx_{1}}}\cdot\mathfrak{p}\mathrm{2tuple}\mathcal{S}_{\mathfrak{p}{idx_{2}}}\cdot\cdots\cdot\mathfrak{p}\mathrm{2tuple}\mathcal{S}_{\mathfrak{p}{idx_{n}}}$
Indeed, if we just calculate the Cartesian product, its rigorous formulation is
$(\cdots((\mathfrak{p}\mathrm{2tuple}\mathcal{S}_{\mathfrak{p}{idx_{1}}}\cdot\mathfrak{p}\mathrm{2tuple}\mathcal{S}_{\mathfrak{p}{idx_{2}}})\cdot\mathfrak{p}\mathrm{2tuple}\mathcal{S}_{\mathfrak{p}{idx_{3}}})\cdots)\cdot\mathfrak{p}\mathrm{2tuple}\mathcal{S}_{\mathfrak{p}{idx_{n}}}$
What we want is not the result restricted by parentheses in terms of operation order, but the flattened result of each element in this resulting set. This is why we use natural isomorphism.
Therefore, the set that is naturally isomorphic to the Cartesian product, as a set of $\mathrm{2Tuple}\mathcal{T}$ instances, can also be expressed in the ordered enumeration form of
$\mathrm{2tuple}\mathcal{TS}^{NatIso2CP}=\{\ \mathrm{2tuple}\mathcal{T}_{1},\ \mathrm{2tuple}\mathcal{T}_{2},\ \cdots,\ \mathrm{2tuple}\mathcal{T}_{k}\ \}$
According to Definition 14, the Cartesian product of all $\mathrm{2tuple}\mathcal{SS}$’ members with the index order $\mathrm{idx}\mathcal{T}$ as the operands, can be expressed as
$\mathrm{2tuple}\mathcal{TS}^{NatIso2CP}=\mathfrak{f}\mathrm{NatIso2CP}(\mathrm{2tuple}\mathcal{SS},\ \mathrm{idx}\mathcal{T})$
and its cardinality is
$\left|{\mathfrak{f}\mathrm{NatIso2CP}(\mathrm{2tuple}\mathcal{SS},\ \mathrm{idx}\mathcal{T})}\right|=\prod_{i=1}^{n}\left|{\mathrm{2tuple}\mathcal{S}_{\mathrm{idx}\mathcal{T}_{i}}}\right|$
We apply this helper function in Definition 14 to $\mathrm{OIE}\mathcal{S}$.
For a finite $\mathrm{OIE}\mathcal{S}$ instance $\mathrm{oie}\mathcal{S}$ with the ordered enumeration form
$\mathrm{oie}\mathcal{S}=\{\ \mathrm{oie}_{1},\ \mathrm{oie}_{2},\ \cdots,\ \mathrm{oie}_{n}\ \}$
The set that is naturally isomorphic to the Cartesian product of all the $\mathcal{I}$(3rd element) of each of $\mathrm{oie}\mathcal{S}$’ members with a certain $\mathrm{Idx}\mathcal{T}$ instance as the index order results in a $\mathrm{2Tuple}\mathcal{TS}$ instance. According to Definition 11 and Definition 14, this result can be expressed as
$\mathrm{2tuple}\mathcal{TS}^{NatIso2CP}=\mathfrak{f}\mathrm{NatIso2CP}(\mathcal{IS}_{\mathrm{oie}\mathcal{S}},\ \mathrm{idx}\mathcal{T})$
4.2 Infeasible and feasible subset of the set that is naturally isomorphic to the Cartesian product of all members of the $\mathrm{2Tuple}\mathcal{SS}$ instance of a finite $\mathrm{OIE}\mathcal{S}$ instance under an index order
The purpose of calculating set that is naturally isomorphic to the Cartesian product of all $\mathcal{I}$(3rd elements) of the members of an $\mathrm{OIE}\mathcal{S}$ instance, is to calculate the maximum combination as for subsequent use. However, since $\mathrm{OIE}$ instances are used to map Event instances, and Event instances are realized in our real world, which is a world with various limitations. As a result, for this $\mathrm{2Tuple}\mathcal{TS}$ instance, we also need to discuss some limiting factors. The content to be discussed in this subsection is that, considering this result as a set, we need to identify which elements of the set are infeasible, and then exclude them. The remaining elements of the set are the available ones.
4.2.1 Set of infeasible interval 2-tuples of members of a finite $\mathrm{OIE}\mathcal{S}$ instance under an index order
For an $\mathrm{OIE}$ instance $\mathrm{oie}$, when analyzing its execution either independently or in combination with other $\mathrm{OIE}$ instances, Whether each interval in the $\mathcal{I}$(3rd element) of $\mathrm{oie}$ can be implemented may change. That is, because there is independence and dependence between Event instances, $\mathrm{OIE}$ instances will also have the same properties.
Given that $\mathrm{oie}$ already has its own $\mathcal{I}$(3rd element), once the existence of other $\mathrm{OIE}$ instances is introduced, The feasibility of each interval in $\mathcal{I}_{\mathrm{oie}}$ may vary, that is, some intervals can no longer be implemented in reality.
Let’s take a simple example. A movie screenwriter is writing a script about the story of a father and a son.
Abstract the entire lives of the father and the son into two $\mathrm{OIE}$ instances: $\mathrm{oie}_{father}$ and $\mathrm{oie}_{son}$, which form an $\mathrm{OIE}\mathcal{S}$ instance named $\mathrm{oie}\mathcal{S}_{father-son}$. When only considering their own individual situations, the screenwriter first makes the following design for their $\mathcal{I}$:
[…]
Here, for the convenience of expression, instead of using moments, years are used, assuming it is 0 o’clock on January 1st of that year. If the script only features one person of the father-son pair and the other is not mentioned at all, the screenwriter can directly select an interval from $\mathcal{I}_{\mathrm{oie}_{father}}$ or from $\mathcal{I}_{\mathrm{oie}_{son}}$ to arrange the birth/death moments of the protagonist. However, if both the father and the son are present in the script, at least the following conditions must be met:
$A.$ The year of the father’s birth must be earlier than that of the son’s.
Even considering the development of history and biotechnology, there is
also a limit to the age difference between the two. The son must be born
after the father’s puberty.
$B.$ The year of the father’s death must be later than the year 10 months before
the son’s birth, unless there is biological assisted technology for sperm
preservation in the historical era.
$C.$ If it is not a science fiction film, each person’s lifespan should be
within a reasonable range.
According to these 3 restrictions, it is obvious that if the father’s lifespan is from 1830 to 1900, then it is unreasonable for the son’s lifespan to be $[1930,2010)$ or $[2077,2277)$.
According to Definition 11 and Definition 14, in the operand order of the father first and then the son, we take the set that is naturally isomorphic to the Cartesian product of both the third items of $\mathrm{oie}_{father}$ and $\mathrm{oie}_{son}$:
$\mathrm{2tuple}\mathcal{TS}^{NatIso2CP}=\mathfrak{f}\mathrm{NatIso2CP}(\mathcal{IS}_{\mathrm{oie}\mathcal{S}_{father-son}},\ (\mathrm{idx}_{father},\mathrm{idx}_{son}))$
Then we get the result:
$\mathrm{2tuple}\mathcal{TS}^{NatIso2CP}=\{$
[…]
This set has 9 elements, and each element is a $\mathrm{2Tuple}\mathcal{T}$ instance. The first item of each $\mathrm{2Tuple}\mathcal{T}$ instance is a lifespan corresponding to the father, and the second item is a lifespan corresponding to the son.
Considering Conditions A, B and C, and combining with the common knowledge corresponding to the known history of human technology, in this set, the screenwriter finally finds that there are only 3 reasonable settings:
Modern-era film: Father[1830, 1900)/Son[1860, 1930)
Contemporary film: Father[1910, 1990)/Son[1930, 2010)
Science-fiction film: Father[2050, 2140)/Son[2077, 2277)
That is to say, in the set composed of the other 6 elements
$\mathcal{S}_{infeasible}=\{$
[…]
The lifespan setup for each father and son cannot occur in reality. If we randomly select an element from the six elements of $\mathcal{S}_{infeasible}$, we cannot create a script about the father and the son. It is impossible for a son to be born in the 1930s while his father died in the 1900.
These impossibilities are not caused by the Cartesian product operation, but by the relationship between the two entities of the father and the son. This is determined by $\mathrm{oie}_{father}$ and $\mathrm{oie}_{son}$, and has nothing to do with the Cartesian product operation involved.
We can understand that the getting the set that is naturally isomorphic to the Cartesian product operation is a framework’s work. The framework can influence the result, and whether the operation result meets your expectations also depends on the special properties between the operands.
Therefore, we will proceed to make new definitions and analysis.
Definition 15 (Set of tuples of infeasible interval 2-tuples of all members of a finite $\mathrm{OIE}\mathcal{S}$ instance under an index order & helper function).
A finite $\mathrm{OIE}\mathcal{S}$ instance $\mathrm{oie}\mathcal{S}$ with a cardinality of $n$ has the ordered enumeration form
$\mathrm{oie}\mathcal{S}=\{\mathrm{oie}_{1},\mathrm{oie}_{2},\mathrm{oie}_{3},\cdots,\mathrm{oie}_{n}\}$
, and an $\mathrm{Idx}\mathcal{T}$ instance $\mathrm{idx}\mathcal{T}$ with form
$\mathrm{idx}\mathcal{T}=(\mathrm{idx}_{1},\mathrm{idx}_{2},\cdots,\mathrm{idx}_{n})$
Take the set that is naturally isomorphic to the Cartesian product of both the third items of members of $\mathrm{oie}\mathcal{S}$ with the index order $\mathrm{idx}\mathcal{T}$:
$\mathrm{2tuple}\mathcal{TS}^{NatIso2CP}=\mathfrak{f}\mathrm{NatIso2CP}(\mathcal{IS}_{\{\mathrm{oie}\mathcal{S}\}},\ \mathrm{idx}\mathcal{T})$
Define a predicate for any $\mathrm{2Tuple}\mathcal{T}$ instance $\mathrm{2tuple}\mathcal{T}$
$\mathbb{P}\mathrm{IsInf2tuple}\mathcal{T}(\mathrm{2tuple}\mathcal{T}):$
$\qquad\mathrm{2tuple}\mathcal{T}\in\mathrm{2tuple}\mathcal{TS}^{NatIso2CP}$
$\quad\land The\ allocation\ of$
$\qquad\qquad(\forall\ i\in\mathrm{idx}\mathcal{T}:\ the\ ith2-tuple\ in\mathrm{2tuple}\mathcal{T}\ is\ the$
$\qquad\qquad\quad interval\ of\ \mathrm{oie}\mathcal{S}_{i}^{\prime}s\ corresponding\ \textsf{Event}\ instance)$
$\qquad is\ infeasible$
Take the largest set consisting of all $\mathrm{2Tuple}\mathcal{T}$ instances that satisfy the predicate $\mathbb{P}\mathrm{IsInf2tuple}\mathcal{T}$. This set is called “The set of tuples of infeasible interval 2-tuples of all members of $\mathrm{oie}\mathcal{S}$ under the index order $\mathrm{idx}\mathcal{T}$”, abbreviated as $\mathrm{2Tuple}\mathcal{TS}^{Inf}$”.
$\mathfrak{f}\mathrm{Inf2tupleTS}(\mathfrak{p}\mathrm{oie}\mathcal{S},\mathfrak{p}\mathrm{idx}\mathcal{T})\rightarrow$
$\quad The\mathrm{2Tuple}\mathcal{TS}^{Inf}\ instance\ of\ \mathfrak{p}\mathrm{oie}\mathcal{S}\ with\ index\ order\ \mathrm{idx}\mathcal{T}$
$\mathfrak{p}\mathrm{oie}\mathcal{S}$ A finite set of $\mathrm{OIE}$ instances
$\mathfrak{p}\mathrm{idx}\mathcal{T}$ An index tuple
Return The set of tuples of infeasible interval 2-tuples of $\mathfrak{p}\mathrm{oie}\mathcal{S}$ with the index order $\mathfrak{p}\mathrm{idx}\mathcal{T}$
The $\mathrm{2Tuple}\mathcal{TS}^{Inf}$ instance of an $\mathrm{OIE}\mathcal{S}$ instance in an index order is a property among the individual elements within the $\mathrm{OIE}\mathcal{S}$ instance, and we have already emphasized this point earlier. How to calculate it depends on these $\mathrm{OIE}$ instances themselves.
Now, we have already used Definition 14 to calculate the set that is naturally isomorphic to the Cartesian product for all members of the $\mathcal{IS}$ instance of the $\mathrm{OIE}\mathcal{S}$ instance, and Definition 15 has provided the set to which the infeasible elements belong. So we can narrow down the scope, we only need to remove the infeasible elements from set that is naturally isomorphic to the Cartesian product to get the usable elements.
Definition 16 (Maximum feasible subset of the set that is naturally isomorphic to the Cartesian product of all members of the $\mathrm{2Tuple}\mathcal{SS}^{Intvl}$ instance of a finite $\mathrm{OIE}\mathcal{S}$ instance in an index order & helper function).
Suppose there are a finite $\mathrm{OIE}\mathcal{S}$ instance with ordered enumeration form
$\mathrm{oie}\mathcal{S}=\{\ \mathrm{oie}_{1},\ \mathrm{oie}_{2},\ \cdots,\ \mathrm{oie}_{n}\ \}$
, and an $\mathrm{Idx}\mathcal{T}$ instance with length $n$
$\mathrm{idx}\mathcal{T}=(idx_{1},idx_{2},\cdots,idx_{n})$
. Take the set that is naturally isomorphic to the Cartesian product of all members of the $\mathcal{IS}$ of $\mathrm{oie}\mathcal{S}$ in the index order specified by $\mathrm{idx}\mathcal{T}$
$\mathrm{2tuple}\mathcal{TS}^{NatIso2CP}=\mathfrak{f}\mathrm{NatIso2CP}(\mathcal{IS}_{\mathrm{oie}\mathcal{S}},\ \mathrm{idx}\mathcal{T})$
For any $\mathrm{2Tuple}\mathcal{T}$ instance with length $n$ and an index order of $\mathrm{idx}\mathcal{T}$
$\mathrm{2tuple}\mathcal{T}=(\ \mathrm{2tuple}_{idx_{1}},\ \mathrm{2tuple}_{idx_{2}},\ \cdots,\ \mathrm{2tuple}_{idx_{n}}\ )$
define the predicate:
$\mathbb{P}\mathrm{IsFeas2tupleT}(\mathrm{2tuple}\mathcal{T}):$
$\qquad\mathrm{2tuple}\mathcal{T}\in\mathrm{2tuple}\mathcal{TS}^{NatIso2CP}$
$\quad\land\mathrm{2tuple}\mathcal{T}\notin\mathfrak{f}\mathrm{Inf2tupleTS}(\mathrm{oie}\mathcal{S},\mathrm{idx}\mathcal{T})$
Take the largest set composed of all $\mathrm{2Tuple}\mathcal{T}$ instances that satisfy the predicate $\mathbb{P}\mathrm{IsFeas2tupleT}$
{ ${\{}\ \mathrm{2tuple}\mathcal{T}\mid\mathbb{P}\mathrm{IsFeas2tupleT}(\mathrm{2tuple}\mathcal{T})\ {\}}$
which is called “The maximum feasible subset of the set that is naturally isomorphic to the Cartesian product of all members of the $\mathrm{2Tuple}\mathcal{SS}^{Intvl}$ instance of $\mathrm{oie}\mathcal{S}$ in the index order specified by $\mathrm{idx}\mathcal{T}$”, denoted as $\mathrm{2Tuple}\mathcal{TS}^{feas}$”.
$\mathfrak{f}\mathrm{Feas2tuple}\mathcal{TS}(\mathfrak{p}\mathrm{oie}\mathcal{S},\mathfrak{p}\mathrm{idx}\mathcal{T})\rightarrow$
$\quad the\mathrm{2Tuple}\mathcal{TS}^{feas}\ instance\ of\mathrm{2tuple}\mathcal{TS}$
$\mathfrak{p}\mathrm{oie}\mathcal{S}$ a finite $\mathrm{OIE}\mathcal{S}$ instance
$\mathfrak{p}\mathrm{idx}\mathcal{T}$ An index tuple
Return The maximum feasible subset of the set that is naturally isomorphic to the Cartesian product of all elements of the $\mathrm{2Tuple}\mathcal{SS}^{Intvl}$ instance of $\mathfrak{p}\mathrm{oie}\mathcal{S}$ in the index order specified by $\mathfrak{p}\mathrm{idx}\mathcal{T}$
4.2.2 Isomorphism relationship of the set of tuples of infeasible interval 2-tuples of $\mathrm{oie}\mathcal{S}$ instances with same members
According to Definition 15
$\mathfrak{f}\mathrm{Inf2tupleTS}(\mathrm{oie}\mathcal{S}_{father\_son},(1,2))=$
$\quad\{$
[…]
$\quad\}$
We have already derived each item in this set one by one earlier. Following our derivation process, we can also arrive at the following conclusion.
$\mathfrak{f}\mathrm{Inf2tupleTS}(\mathrm{oie}\mathcal{S}_{father\_son},(2,1))=$
$\quad\{$
[…]
$\quad\}$
The difference between these two sets arises because the second parameter is $(1,2)$ in first case and $(2,1)$ in the other. In this part, we will discuss the relationship between these two sets.
$(1,2)$ and $(2,1)$ are $\mathrm{Idx}\mathcal{T}$ instances for a set with cardinality 2. For any $\mathrm{OIE}\mathcal{S}$ instance $\mathrm{oie}\mathcal{S}$ with cardinality n, if $\mathrm{idx}\mathcal{T}_{1}$ and $\mathrm{idx}\mathcal{T}_{2}$ are $\mathrm{Idx}\mathcal{T}$ instances for $\mathrm{oie}\mathcal{S}$, then $\mathrm{idx}\mathcal{T}_{1}$ and $\mathrm{idx}\mathcal{T}_{2}$ are permutationally equivalent. For example, let
$\mathrm{idx}\mathcal{T}_{1}=(1,2),\ \mathrm{idx}\mathcal{T}_{2}=(2,1)$
$\mathrm{idx}\mathcal{T}_{3}=(1,2,3),\ \mathrm{idx}\mathcal{T}_{4}=(2,3,1)$
$\mathcal{M}_{1}=\begin{pmatrix}0&1\\ 1&0\end{pmatrix},\ \mathcal{M}_{2}=\begin{pmatrix}0&0&1\\ 1&0&0\\ 0&1&0\end{pmatrix}$
They satisfy the following equations
[…]
denoted as
$\mathrm{idx}\mathcal{T}_{1}\stackrel{{\scriptstyle\mathcal{M}_{1}}}{{\sim}}\mathrm{idx}\mathcal{T}_{2},\ \mathrm{idx}\mathcal{T}_{3}\stackrel{{\scriptstyle\mathcal{M}_{2}}}{{\sim}}\mathrm{idx}\mathcal{T}_{4}$
The elements of two sets
$\mathfrak{f}\mathrm{Inf2tupleTS}(\mathrm{oie}\mathcal{S}_{father\_son},(1,2))$
$\mathfrak{f}\mathrm{Inf2tupleTS}(\mathrm{oie}\mathcal{S}_{father\_son},(2,1))$
also have the permutationally equivalent relationship.
 permutation matrix $\exists\text${ permutation matrix $\mathcal{M}$,
$\quad\forall tuple_{i}\in\mathfrak{f}\mathrm{Inf2tupleTS}(\mathrm{oie}\mathcal{S}_{father\_son},(1,2)),$
$\qquad\exists tuple_{j}\in\mathfrak{f}\mathrm{Inf2tupleTS}(\mathrm{oie}\mathcal{S}_{father\_son},(2,1)):$
$\quad\qquad tuple_{i}\stackrel{{\scriptstyle\mathcal{M}}}{{\sim}}tuple_{j}\ (\text{i.e. }tuple_{i}\cdot\mathcal{M}=tuple_{j})$
For example,
[…]
$\cdots$
So, We can conclude the property
Property 9 (Structure-preserving permutational isomorphism of $\mathrm{2Tuple}\mathcal{TS}^{Inf}$ instances).
Suppose there are an $\mathrm{OIE}\mathcal{S}$ instance $\mathrm{oie}\mathcal{S}$ with cardinality n, two $\mathrm{Idx}\mathcal{T}$ instances
$\mathrm{idx}\mathcal{T}_{1},\mathrm{idx}\mathcal{T}_{2}$
for $\mathrm{oie}\mathcal{S}$. Then
$\exists\ a\ permutation\ matrix\ \mathcal{M}:$
$\qquad\mathrm{idx}\mathcal{T}_{1}\cdot\mathcal{M}=\mathrm{idx}\mathcal{T}_{2}$
$\quad\land\mathfrak{f}\mathrm{Inf2tupleTS}(\mathrm{oie}\mathcal{S},\mathrm{idx}\mathcal{T}_{1})\ and\ \mathfrak{f}\mathrm{Inf2tupleTS}(\mathrm{oie}\mathcal{S},\mathrm{idx}\mathcal{T}_{2})$
$\qquad\quad are\ structure-preserving\ permutationally\ equivalent\$
$\qquad\quad with\ respect\ to\ matrix\ \mathcal{M}.$
denoted as
$\mathrm{idx}\mathcal{T}_{1}\stackrel{{\scriptstyle\mathcal{M}}}{{\sim}}\mathrm{idx}\mathcal{T}_{2}\rightarrow$
$\quad\mathfrak{f}\mathrm{Inf2tupleTS}(\mathrm{oie}\mathcal{S},\mathrm{idx}\mathcal{T}_{1})\stackrel{{\scriptstyle\mathcal{M}}}{{\cong}}\mathfrak{f}\mathrm{Inf2tupleTS}(\mathrm{oie}\mathcal{S},\mathrm{idx}\mathcal{T}_{2})$
In Subsection 4.1, we have discussed the set that is naturally isomorphic to the Cartesian product of the sets of interval 2-tuples of each member in a finite $\mathrm{OIE}\mathcal{S}$ instance. Indeed, two results with same $\mathrm{2Tuple}\mathcal{S}$ instances as operands but different orders of operations are structure-preserving permutationally equivalent. The proof follows the same approach as the one we used to discuss Property 9, and it is widely accepted.
Thus, whether we eliminate an $\mathrm{2Tuple}\mathcal{TS}^{Inf}$ instance from a Cartesian product or derive $\mathrm{2Tuple}\mathcal{TS}^{feas}$ directly using the same method, we arrive at the same property for $\mathrm{2Tuple}\mathcal{TS}^{feas}$.
Property 10 (Structure-preserving permutational isomorphism of $\mathrm{2Tuple}\mathcal{TS}^{feas}$ instances).
Suppose there are an $\mathrm{OIE}\mathcal{S}$ instance $\mathrm{oie}\mathcal{S}$ with cardinality n
$\mathrm{oie}\mathcal{S}=\{\mathrm{oie}_{1},\mathrm{oie}_{2},\cdots,\mathrm{oie}_{n}\}$
and two $\mathrm{Idx}\mathcal{T}$ instances with length n
$\mathrm{idx}\mathcal{T}_{1},\ \mathrm{idx}\mathcal{T}_{2}$
Then
$\exists\ a\ permutation\ matrix\ \mathcal{M}:$
$\qquad\mathrm{idx}\mathcal{T}_{1}\cdot\mathcal{M}=\mathrm{idx}\mathcal{T}_{2}$
$\quad\land\mathfrak{f}\mathrm{Feas2tuple}\mathcal{TS}(\mathrm{oie}\mathcal{S},\mathrm{idx}\mathcal{T}_{1})\ and\ \mathfrak{f}\mathrm{Feas2tuple}\mathcal{TS}(\mathrm{oie}\mathcal{S},\mathrm{idx}\mathcal{T}_{2})$
$\qquad\quad are\ structure-preserving\ permutationally\ equivalent\$
$\qquad\quad with\ respect\ to\ matrix\ \mathcal{M}.$
denoted as
$\mathrm{idx}\mathcal{T}_{1}\stackrel{{\scriptstyle\mathcal{M}}}{{\sim}}\mathrm{idx}\mathcal{T}_{2}\rightarrow$
$\quad\mathfrak{f}\mathrm{Feas2tuple}\mathcal{TS}(\mathrm{oie}\mathcal{S},\mathrm{idx}\mathcal{T}_{1})\stackrel{{\scriptstyle\mathcal{M}}}{{\cong}}\mathfrak{f}\mathrm{Feas2tuple}\mathcal{TS}(\mathrm{oie}\mathcal{S},\mathrm{idx}\mathcal{T}_{2})$
Here we clarify that, infeasibility constraints we have discussed in this subsection are inherent properties determined by the internal relationships among the participating $\mathrm{OIE}$ instances, these constraints reflect the physical, logical, or causal limitations between the corresponding real-world events. An innovative aspect of this paper lies in “quantifying event execution from the perspective of subjective planning”. During the process of subjective planning, it is essential to effectively screen out invalid results that violate physical laws, logical constraints or causal relationships. Therefore, this subsection serves as a crucial bridge connecting the purely formalized theory of this paper with practical applications.
4.2.3 Mutually independent $\mathrm{OIE}\mathcal{S}$
Definition 17 (Mutually independent $\mathrm{OIE}\mathcal{S}$).
Suppose there is an $\mathrm{OIE}\mathcal{S}$ instance $\mathrm{oie}\mathcal{S}$ with cardinality $n\in\mathbb{N}^{+}$, with form
$\mathrm{oie}\mathcal{S}=\{\mathrm{oie}_{1},\mathrm{oie}_{2},\cdots,\mathrm{oie}_{n}\}$
and every element therein must not be $\mathrm{oie}_{\mathrm{void}}$. Let
$\mathrm{idx}\mathcal{T}_{asc}=(1,2,3,\cdots,n)$
If
$\mathfrak{f}\mathrm{Inf2tupleTS}(\mathrm{oie}\mathcal{S},\mathrm{idx}\mathcal{T}_{asc})=\emptyset$
then $\mathrm{oie}\mathcal{S}$ is called mutually independent $\mathrm{oie}\mathcal{S}$. Such instances are called “Mutually Independent $\mathrm{OIE}\mathcal{S}$”, denoted as $\mathrm{OIE}\mathcal{S}^{\text{indep}}$.
Scenarios including 3 doctors submitting their manuscripts via their respective computers, multiple computing nodes in a supercomputing center executing tasks without mutual interference, and several athletes competing in their individual lanes without affecting one another, can all be characterized by Definition 17. And many subsequent examples in this paper will also be based on Definition 17.
4.3 Complete sequential addition
Building on the previous content, in this section, we will for the first time propose an operation for $\mathrm{OIE}$ instances. The result of this operation will have a significant impact on the intervals of the events that the participating $\mathrm{OIE}$ instances are mapped to. Since this impact is related to the execution order of the Event instances they are mapped to, we introduce the term “Sequential Operation”. Before delving into the details of the concept, we will first prepare several mathematical helper functions.
Definition 18 (Function for finding the minimum 1st item and the maximum 2nd item of a tuple of 2-tuples & permutation invariance).
1) definition & functions:
Suppose there is a $\mathrm{2Tuple}\mathcal{T}$ instance $\mathrm{2tuple}\mathcal{T}$ with length $n\geq 1$, with form
$(\ (x_{1},\ y_{1}),\ (x_{2},\ y_{2}),\ \cdots,\ (x_{n},\ y_{n})\ )$
We defined two extremal functions:
$\mathfrak{f}\mathrm{Min1of2tuple}\mathcal{T}(\mathrm{2tuple}\mathcal{T})\rightarrow min\{x_{i}|i\in\{1,2,\cdots,n\}\}$
$\mathrm{2tuple}\mathcal{T}$ A tuple of 2-tuples(a $\mathrm{2Tuple}\mathcal{T}$ instance)
Return The minimum 1st item among all elements of $\mathrm{2tuple}\mathcal{T}$
$\mathfrak{f}\mathrm{Max2of2tuple}\mathcal{T}(\mathrm{2tuple}\mathcal{T})\rightarrow\max\{y_{i}|i\in\{1,2,\cdots,n\}\}$
$\mathfrak{p}\mathrm{2tuple}\mathcal{T}$ A tuple of 2-tuples
Return The maximum 2nd item among all elements of $\mathrm{2tuple}\mathcal{T}$
2) permutation invariance:
Suppose there are two $\mathrm{2Tuple}\mathcal{T}$ instances $\mathrm{2tuple}\mathcal{T}$ and $\mathrm{2tuple}\mathcal{T}^{{}^{\prime}}$, with the same length $n$, and form
$\mathrm{2tuple}\mathcal{T}=(\ (x_{1},\ y_{1}),\ (x_{2},\ y_{2}),\ \cdots,\ (x_{n},\ y_{n})\ ),$
$\mathrm{2tuple}\mathcal{T}^{{}^{\prime}}=(\ (x_{1}^{{}^{\prime}},\ y_{1}^{{}^{\prime}}),\ (x_{2}^{{}^{\prime}},\ y_{2}^{{}^{\prime}}),\cdots,\ (x_{n}^{{}^{\prime}},\ y_{n}^{{}^{\prime}})\ )$
and there is a permutation matrix $\mathcal{M}$ satisfying
$\mathrm{2tuple}\mathcal{T}\stackrel{{\scriptstyle\mathcal{M}}}{{\sim}}\mathrm{2tuple}\mathcal{T}^{{}^{\prime}}$
Since the two $\mathrm{2Tuple}\mathcal{T}$ instances only change the order of their internal elements, without altering the existence or values of the elements, we conclude that:
$\mathrm{2tuple}\mathcal{T}\stackrel{{\scriptstyle\mathcal{M}}}{{\sim}}\mathrm{2tuple}\mathcal{T}^{{}^{\prime}}\rightarrow$
$\qquad\quad\mathfrak{f}\mathrm{Min1of2tuple}\mathcal{T}(\mathrm{2tuple}\mathcal{T})=\mathfrak{f}\mathrm{Min1of2tuple}\mathcal{T}(\mathrm{2tuple}\mathcal{T}^{{}^{\prime}})$
$\qquad\land\ \mathfrak{f}\mathrm{Max2of2tuple}\mathcal{T}(\mathrm{2tuple}\mathcal{T})=\mathfrak{f}\mathrm{Max2of2tuple}\mathcal{T}(\mathrm{2tuple}\mathcal{T}^{{}^{\prime}})$
Definition 19 (Domain-filtered subset of a non-empty set of tuples of 2-tuples under a domain-filtering 2-tuple & helper functions & permutation isomorphism invariance).
1) definition:
Consider a non-empty $\mathrm{2Tuple}\mathcal{TS}$ instance $\mathrm{2tuple}\mathcal{TS}$
$\mathrm{2tuple}\mathcal{TS}=\{\ \mathrm{2tuple}\mathcal{T}_{1},\ \mathrm{2tuple}\mathcal{T}_{2},\ \cdots,\ \mathrm{2tuple}\mathcal{T}_{m}\ \}$
with every element has the same length $0<n<+\infty$, and two real numbers $\alpha$ and $\beta$($\alpha<\beta$), we build a set
$\mathcal{S}=\{\ \mathrm{2tuple}\mathcal{T}\in\mathrm{2tuple}\mathcal{TS}\ |\ \forall(x,y)\in\mathrm{2tuple}\mathcal{T}:\ x\geq\alpha\land y\leq\beta\}$
Then the set
$\mathrm{2tuple}\mathcal{TS}_{dom}=$
[…]
is called “The domain-filtered subset of $\mathrm{2tuple}\mathcal{TS}$ under the domain-filtering 2-tuple $(\alpha,\beta)$”, denoted as $\mathrm{2Tuple}\mathcal{TS}_{\textbf{Dom}}|_{\alpha}^{\beta}$”. $(\alpha,\beta)$ is a domain-filtering 2-tuple of $\mathrm{2tuple}\mathcal{TS}$, written as $\mathrm{DomFlt2Tuple}$”.
2) helper function:
$\mathfrak{f}\mathrm{DomFlt}\mathrm{2tuple}\mathcal{TS}(\mathfrak{p}\mathrm{2tuple}\mathcal{TS},\mathfrak{p}\alpha,\mathfrak{p}\beta)\rightarrow A\mathrm{2Tuple}\mathcal{TS}_{\textbf{Dom}}|_{\mathfrak{p}\alpha}^{\mathfrak{p}\beta}\ instance$
$\mathfrak{p}\mathrm{2tuple}\mathcal{TS}$ A non-empty set of tuples of 2-tuples
$\mathfrak{p}\mathrm{A}$ The left boundary of the domain-filtered subset
$\mathfrak{p}\mathrm{B}$ The right boundary of the domain-filtered subset
Return The domain-filtered subset of $\mathfrak{p}\mathrm{2tuple}\mathcal{TS}$ with under $(\mathfrak{p}\mathrm{A},\mathfrak{p}\mathrm{B})$
3) permutation isomorphism invariance:
Keeping $\mathfrak{p}\alpha$ and $\mathfrak{p}\beta$ ($\mathfrak{p}\alpha<\mathfrak{p}\beta$) fixed, if there are two $\mathrm{2Tuple}\mathcal{TS}$ instances $\mathrm{2tuple}\mathcal{TS}_{1}$ and $\mathrm{2tuple}\mathcal{TS}_{2}$ with the same length $n$ and satisfying isomorphic via permutation matrix $\mathcal{M}$, then the two return value of this function are isomorphic via permutation matrix $\mathcal{M}$ too. i.e.
$\mathrm{2tuple}\mathcal{TS}_{1}\stackrel{{\scriptstyle\mathcal{M}}}{{\cong}}\mathrm{2tuple}\mathcal{TS}_{2}\rightarrow$
$\qquad\mathfrak{f}\mathrm{DomFlt}\mathrm{2tuple}\mathcal{TS}(\mathrm{2tuple}\mathcal{TS}_{1},\mathfrak{p}\alpha,\mathfrak{p}\beta)\stackrel{{\scriptstyle\mathcal{M}}}{{\cong}}\mathfrak{f}\mathrm{DomFlt}\mathrm{2tuple}\mathcal{TS}(\mathrm{2tuple}\mathcal{TS}_{2},\mathfrak{p}\alpha,\mathfrak{p}\beta)$
Proof of the permutation isomorphism invariance in Definition 19.
Suppose there are two non-empty $\mathrm{2Tuple}\mathcal{TS}$ instances $\mathrm{2tuple}\mathcal{TS}_{1}$ and $\mathrm{2tuple}\mathcal{TS}_{2}$ with the same length $n$ and satisfying isomorphic via permutation matrix $\mathcal{M}$
$\mathrm{2tuple}\mathcal{TS}_{1}\stackrel{{\scriptstyle\mathcal{M}}}{{\cong}}\mathrm{2tuple}\mathcal{TS}_{2}$
This means that there is a bijection
$\exists\ (\phi:\mathrm{2tuple}\mathcal{TS}_{1}\rightarrow\mathrm{2tuple}\mathcal{TS}_{2})\ :$
$\qquad\forall\mathrm{2tuple}\mathcal{T}\in\mathrm{2tuple}\mathcal{TS}_{1}:\phi(\mathrm{2tuple}\mathcal{T})\in\mathrm{2tuple}\mathcal{TS}_{2}$
where $\phi(\mathrm{2tuple}\mathcal{T})=\mathrm{2tuple}\mathcal{T}\cdot\mathcal{M}$.
For fixed $\alpha$ and $\beta$ ($\alpha<\beta$), we examine the computation processes of
$\mathfrak{f}\mathrm{DomFlt}\mathrm{2tuple}\mathcal{TS}(\mathrm{2tuple}\mathcal{TS}_{1},\alpha,\beta)$ and $\mathfrak{f}\mathrm{DomFlt}\mathrm{2tuple}\mathcal{TS}(\mathrm{2tuple}\mathcal{TS}_{2},\alpha,\beta)$ separately.
Step I. Construct intermediate sets $\mathcal{S}_{interm1}$ and $\mathcal{S}_{interm2}$.
$\mathcal{S}_{1}=\{\mathrm{2tuple}\mathcal{T}\in\mathrm{2tuple}\mathcal{TS}_{1}\ |\ \forall(x,y)\in\mathrm{2tuple}\mathcal{T}:\ x\geq\alpha\land y\leq\beta\ \},$
$\mathcal{S}_{2}=\{\mathrm{2tuple}\mathcal{T}\in\mathrm{2tuple}\mathcal{TS}_{2}\ |\ \forall(x,y)\in\mathrm{2tuple}\mathcal{T}:\ x\geq\alpha\land y\leq\beta\ \}$
For any
$\mathrm{2tuple}\mathcal{T}_{\lambda}\in\mathcal{S}_{1}$
let
$\mathrm{2tuple}\mathcal{T}_{\mu}=\phi(\mathrm{2tuple}\mathcal{T}_{\lambda})=\mathrm{2tuple}\mathcal{T}_{\lambda}\cdot\mathcal{M}$
Since $\mathcal{M}$ only permutes the 2-tuples inside $\mathrm{2tuple}\mathcal{T}_{\lambda}$ structure without altering the set of these 2-tuples, we have:
$\forall(x_{\lambda},y_{\lambda})\in\mathrm{2tuple}\mathcal{T}_{\lambda}\Longleftrightarrow$
$\qquad\exists(x_{\mu},y_{\mu})\in\mathrm{2tuple}\mathcal{T}_{\mu}:x_{\lambda}=x_{\mu}\land y_{\lambda}=y_{\mu}$
Thus, the condition
$\forall(x_{\lambda},y_{\lambda})\in\mathrm{2tuple}\mathcal{T}_{\lambda}:x_{\lambda}\geq\alpha\land y_{\lambda}\leq\beta$
holds if and only if the condition
$\forall(x_{\mu},y_{\mu})\in\mathrm{2tuple}\mathcal{T}_{\mu}:x_{\mu}\geq\alpha\land y_{\mu}\leq\beta$
holds.
This indicates that
$\mathrm{2tuple}\mathcal{T}_{\lambda}\in\mathcal{S}_{1}$
if and only if
$\mathrm{2tuple}\mathcal{T}_{\mu}\in\mathcal{S}_{2}$
Thus, at the set level, this relationship can be expressed as
$\mathcal{S}_{2}=\phi(\mathcal{S}_{1})$
This precisely means that $\mathcal{S}_{1}$ and $\mathcal{S}_{2}$ are isomorphic via permutation matrix $\mathcal{M}$.
Step II. permutation invariance of boundary conditions:
According to Definition 18, permutation invariance holds for both $\mathfrak{f}\mathrm{Min1of2tuple}\mathcal{T}$ and $\mathfrak{f}\mathrm{Max2of2tuple}\mathcal{T}$.
$\mathfrak{f}\mathrm{Min1of2tuple}\mathcal{T}(\mathrm{2tuple}\mathcal{T}_{\mu})$ $=\mathfrak{f}\mathrm{Min1of2tuple}\mathcal{T}(\phi(\mathrm{2tuple}\mathcal{T}_{\lambda}))$
$\mathfrak{f}\mathrm{Max2of2tuple}\mathcal{T}(\mathrm{2tuple}\mathcal{T}_{\mu})$ $=\mathfrak{f}\mathrm{Max2of2tuple}\mathcal{T}(\phi(\mathrm{2tuple}\mathcal{T}_{\lambda}))$
Since we have got $\mathcal{S}_{2}=\phi(\mathcal{S}_{1})$, so the following equivalence holds:
$\forall k\in[1,n],\exists\mathrm{2tuple}\mathcal{T}\in\mathcal{S}_{1}:the1st\ item\ of\ the\ \text{k-th}\ elem\ of\mathrm{2tuple}\mathcal{T}=\alpha\Longleftrightarrow$
$\qquad\forall k\in[1,n],\exists\phi(\mathrm{2tuple}\mathcal{T})\in\mathcal{S}_{2}:the1st\ item\ of\ the\ \text{k-th}\ elem\ of\ \phi(\mathrm{2tuple}\mathcal{T})=\alpha$
$\forall k\in[1,n],\exists\mathrm{2tuple}\mathcal{T}\in\mathcal{S}_{1}:the2nd\ item\ of\ the\ \text{k-th}\ elem\ of\mathrm{2tuple}\mathcal{T}=\beta\Longleftrightarrow$
$\qquad\forall k\in[1,n],\exists\phi(\mathrm{2tuple}\mathcal{T})\in\mathcal{S}_{2}:the2nd\ item\ of\ the\ \text{k-th}\ elem\ of\ \phi(\mathrm{2tuple}\mathcal{T})=\beta$
$\exists k\in[1,n],\nexists\mathrm{2tuple}\mathcal{T}\in\mathcal{S}_{1}:the1st\ item\ of\ the\ \text{k-th}\ elem\ of\mathrm{2tuple}\mathcal{T}=\alpha\Longleftrightarrow$
$\qquad\exists k\in[1,n],\nexists\phi(\mathrm{2tuple}\mathcal{T})\in\mathcal{S}_{2}:the1st\ item\ of\ the\ \text{k-th}\ elem\ of\ \phi(\mathrm{2tuple}\mathcal{T})=\alpha$
$\exists k\in[1,n],\nexists\mathrm{2tuple}\mathcal{T}\in\mathcal{S}_{1}:the2nd\ item\ of\ the\ \text{k-th}\ elem\ of\mathrm{2tuple}\mathcal{T}=\beta\Longleftrightarrow$
$\qquad\exists k\in[1,n],\nexists\phi(\mathrm{2tuple}\mathcal{T})\in\mathcal{S}_{2}:the2nd\ item\ of\ the\ \text{k-th}\ elem\ of\ \phi(\mathrm{2tuple}\mathcal{T})=\beta$
Step III. conclusion From Step II, it follows that, $\mathcal{S}_{interm1}$ satisfies the condition of being non-empty under the boundary rule if and only if $\mathcal{S}_{interm2}$ satisfies the condition of being non-empty under the boundary rule.
According to Definition 19, if both $\mathcal{S}_{interm1}$ and $\mathcal{S}_{interm2}$ satisfy the condition and are non-empty
$\mathfrak{f}\mathrm{DomFlt}\mathrm{2tuple}\mathcal{TS}(\mathrm{2tuple}\mathcal{TS}_{1},\alpha,\beta)=\mathcal{S}_{interm1},$
$\mathfrak{f}\mathrm{DomFlt}\mathrm{2tuple}\mathcal{TS}(\mathrm{2tuple}\mathcal{TS}_{2},\alpha,\beta)=\mathcal{S}_{interm2}=\phi(\mathcal{S}_{interm1})$
Since $\phi$ is precisely the permutation mapping defined by $\mathcal{M}$, the two results are isomorphic via permutation matrix $\mathcal{M}$.
Otherwise, if they both don’t satisfy the condition, they are both $\emptyset$, $\emptyset$ is isomorphic to itself via any permutation matrix.
To conclude, we have proven the permutation isomorphism invariance in Definition 19.
According to Definition 19, the domain-filtered subset of a $\mathrm{2Tuple}\mathcal{TS}$ instance $\mathrm{2tuple}\mathcal{TS}$ with respect to the domain-filtering 2-tuple $(\alpha,\beta)$ can be expressed as
$\mathrm{2tuple}\mathcal{TS}_{dom}|_{\alpha}^{\beta}=\mathfrak{f}\mathrm{DomFlt}\mathrm{2tuple}\mathcal{TS}(\mathrm{2tuple}\mathcal{TS},\alpha,\beta)$
When $\mathrm{2tuple}\mathcal{TS}_{dom}|_{\alpha}^{\beta}=\emptyset$, it means that there is no domain-filtered subset of $\mathrm{2tuple}\mathcal{TS}$ under $(\alpha,\beta)$.
The function of this method is that it can use a range, a domain-filtering 2-tuple, to obtain a certain type of subset from a $\mathrm{2Tuple}\mathcal{TS}$ instance. The characteristic of this subset is that each element of the subset is restricted within the filtering domain. Moreover, there must be an element in this subset such that for a certain 2-tuple within this element, its first item is equal to the left boundary of the subset range; and there must also be an element in this subset such that for a certain 2-tuple within this element, its second item is equal to the right boundary of the subset range.
To utilize this subset, we further define several variable types and related helper functions.
Using the helper function in Definition 7, we can take the $\mathrm{2Tuple}\mathcal{TS}$ instance obtained from Definition 19 and calculate a bound 2-tuple. These will be used when performing sequential operations on $\mathrm{OIE}\mathcal{S}$ to generate a new $\mathrm{Intvl2Tuple}\mathcal{S}$ instance for the operation result.
Now, we propose a sequential operation for $\mathrm{OIE}\mathcal{S}$, the complete sequential addition of members of a finite $\mathrm{OIE}\mathcal{S}$ under a domain-filtering 2-tuple in an index order. All members within an $\mathrm{OIE}\mathcal{S}$ instance participate in this operation, and the result of the operation is a new $\mathrm{OIE}$ instance.
Definition 20 (Complete sequential addition of members of a finite $\mathrm{OIE}\mathcal{S}$ instance under a domain-filtering 2-tuple in an index order).
There is a finite $\mathrm{OIE}\mathcal{S}$ instance $\mathrm{oie}\mathcal{S}$ containing $n$ ($n\neq\infty$) elements in the ordered enumeration form
$\mathrm{oie}\mathcal{S}=\{\mathrm{oie}_{1},\mathrm{oie}_{2},\mathrm{oie}_{3},\cdots,\mathrm{oie}_{n}\}$
, an $\mathrm{Idx}\mathcal{T}$ instance with length $n$
$\mathrm{idx}\mathcal{T}=(\mathrm{idx}_{1},\mathrm{idx}_{2},\mathrm{idx}_{3},\cdots,\mathrm{idx}_{n})$
and a $\mathrm{DomFlt2Tuple}$ instance
$\mathrm{domFlt2tuple}=(\alpha,\beta)$
And for $\mathrm{oie}\mathcal{S}$ and $\mathrm{Idx}\mathcal{T}$ instance
$\mathrm{idx}\mathcal{T}_{asc}=(1,\ 2,\ 3,\ \cdots,\ n)$
The $\mathrm{2Tuple}\mathcal{TS}^{Inf}$ instance of $\mathrm{oie}\mathcal{S}$ in the index order $\mathrm{idx}\mathcal{T}_{asc}$ is
$\mathrm{2tuple}\mathcal{TS}^{inf}_{\mathrm{oie}\mathcal{S}}=\mathfrak{f}\mathrm{Inf2tupleTS}(\mathrm{oie}\mathcal{S},\mathrm{idx}\mathcal{T}_{asc})$
Define a finitary operation $\oplus_{n}|_{\alpha}^{\beta}$}”, its operational rules are as follows:
Step 1. If
$\bigcap_{i=1}^{n}\mathcal{A}_{\mathrm{oie}_{\mathrm{idx}_{i}}}\neq\emptyset$
$\lor$ $\exists\ k\ \in\ [1,n]:\ \mathrm{oie}_{\mathrm{idx}_{k}}=\mathrm{oie}_{\mathrm{void}}$
then, the result is directly set to the $\mathrm{OIE}_{\mathrm{void}}$ instance, with form
$\mathrm{oie}_{\mathrm{void}}={{(}}\ (),\ \emptyset,\ \emptyset,\ \emptyset\ {{)}}$
and the operation will terminate. Otherwise, initialize a new $\mathrm{OIE}$ instance $\mathrm{oie}_{result}$ as the result. Temporarily set the $\mathcal{C}$(1st element) of $\mathrm{oie}_{result}$ to the result of tuple flattening for all the operands into a tuple by the order of $\mathrm{idx}\mathcal{T}$
$\mathcal{C}_{\mathrm{oie}_{result}}=(\mathrm{oie}_{\mathrm{idx}_{1}},\ \mathrm{oie}_{\mathrm{idx}_{2}},\ \mathrm{oie}_{\mathrm{idx}_{3}},\ \cdots,\ \mathrm{oie}_{\mathrm{idx}_{n}})$
, and temporarily set the $\mathcal{A}$(4th element) of $\mathrm{oie}_{result}$ to set
$\mathcal{A}_{\mathrm{oie}_{result}}=\bigcup_{i=1}^{n}\mathcal{A}_{\mathrm{oie}_{\mathrm{idx}_{i}}}$
then proceed to Step 2.
Step 2. Using Definition 16, take the feasible subset of the set that is naturally isomorphic to the Cartesian product of all members of the $\mathrm{2Tuple}\mathcal{SS}^{Intvl}$ instance of $\mathrm{oie}\mathcal{S}$ in the index order $\mathrm{idx}\mathcal{T}$, denoted as $\mathrm{2tuple}\mathcal{TS}^{feas}$
$\mathrm{2tuple}\mathcal{TS}^{feas}=\mathfrak{f}\mathrm{Feas2tuple}\mathcal{TS}(\mathrm{oie}\mathcal{S},\mathrm{idx}\mathcal{T})$
If
$\mathrm{2tuple}\mathcal{TS}^{feas}=\emptyset$
then $\mathrm{oie}_{result}$ is instead set to $\mathrm{oie}_{\mathrm{void}}$, with form
$\mathrm{oie}_{\mathrm{void}}={{(}}\ (),\ \emptyset,\ \emptyset,\ \emptyset\ {{)}}$
and the operation ends without performing the subsequent steps. Otherwise, proceed to Step 3.
Step 3. Using Definition 19, get the corresponding $\mathrm{2Tuple}\mathcal{TS}_{\textbf{Dom}}\big|_{\alpha}^{\beta}$ instance
$\mathrm{2tuple}\mathcal{TS}_{dom}\big|_{\alpha}^{\beta}$ $=\mathfrak{f}\mathrm{DomFlt}\mathrm{2tuple}\mathcal{TS}(\mathrm{2tuple}\mathcal{TS}^{feas},\alpha,\beta)$
$=\mathfrak{f}\mathrm{DomFlt}\mathrm{2tuple}\mathcal{TS}(\mathfrak{f}\mathrm{Feas2tuple}\mathcal{TS}(\mathrm{oie}\mathcal{S},\mathrm{idx}\mathcal{T}),\alpha,\beta)$
with form
$\mathrm{2tuple}\mathcal{TS}_{dom}\big|_{\alpha}^{\beta}=\{\mathrm{2tuple}\mathcal{T}_{1},\mathrm{2tuple}\mathcal{T}_{2},\cdots,\mathrm{2tuple}\mathcal{T}_{\lambda}\}$
If
$\mathrm{2tuple}\mathcal{TS}_{dom}\big|_{\alpha}^{\beta}\neq\emptyset$
then the $\mathcal{F}$(2nd element) of $\mathrm{oie}_{result}$ is assigned to $\mathrm{2tuple}\mathcal{TS}_{dom}\big|_{\alpha}^{\beta}$
$\mathcal{F}_{\mathrm{oie}_{result}}=\mathrm{2tuple}\mathcal{TS}_{dom}\big|_{\alpha}^{\beta}$
and then Step 4 is executed. Otherwise, if
$\mathrm{2tuple}\mathcal{TS}_{dom}\big|_{\alpha}^{\beta}=\emptyset$
$\mathrm{oie}_{result}$ is instead set to $\mathrm{oie}_{\mathrm{void}}$, with form
$\mathrm{oie}_{\mathrm{void}}={{(}}\ (),\ \emptyset,\ \emptyset,\ \emptyset\ {{)}}$
And the operation ends without performing the subsequent steps.
Step 4. Using Definition 5, get the $\mathcal{I}$(3rd element) of $\mathrm{oie}_{result}$
$\mathcal{I}_{\mathrm{oie}_{result}}$ $=\bigg\{\ \mathfrak{f}\mathrm{Bound2tuple}(\mathrm{2tuple}\mathcal{T}_{i})\ \bigg|\mathrm{2tuple}\mathcal{T}_{i}\in\mathrm{2tuple}\mathcal{TS}_{dom}\big|_{\alpha}^{\beta}\ \bigg\}$
$=\bigg\{\ \mathfrak{f}\mathrm{Bound2tuple}(\mathrm{2tuple}\mathcal{T}_{i})\ \bigg|\$
$\qquad\quad\mathrm{2tuple}\mathcal{T}_{i}\in\mathfrak{f}\mathrm{DomFlt}\mathrm{2tuple}\mathcal{TS}(\mathfrak{f}\mathrm{Feas2tuple}\mathcal{TS}(\mathrm{oie}\mathcal{S},\mathrm{idx}\mathcal{T}),\alpha,\beta)\ \bigg\}$
At this juncture, the operation is completed.
Finally, the form of $\mathrm{oie}_{result}$ is
$\mathrm{oie}_{result}={{(}}\ (),\ \emptyset,\ \emptyset,\ \emptyset\ {{)}}$
or
$\mathrm{oie}_{result}=$
$\quad{{(}}$
$\qquad(\mathrm{oie}_{\mathrm{idx}_{1}},\ \mathrm{oie}_{\mathrm{idx}_{2}},\ \mathrm{oie}_{\mathrm{idx}_{3}},\ \cdots,\ \mathrm{oie}_{\mathrm{idx}_{n}}),$
$\qquad\mathfrak{f}\mathrm{DomFlt}\mathrm{2tuple}\mathcal{TS}(\mathfrak{f}\mathrm{Feas2tuple}\mathcal{TS}(\mathrm{oie}\mathcal{S},\mathrm{idx}\mathcal{T}),\alpha,\beta),$
$\qquad\bigg\{\ \mathfrak{f}\mathrm{Bound2tuple}(\mathrm{2tuple}\mathcal{T}_{i})\ \bigg|\$
$\qquad\qquad\mathrm{2tuple}\mathcal{T}_{i}\in\mathfrak{f}\mathrm{DomFlt}\mathrm{2tuple}\mathcal{TS}(\mathfrak{f}\mathrm{Feas2tuple}\mathcal{TS}(\mathrm{oie}\mathcal{S},\mathrm{idx}\mathcal{T}),\alpha,\beta)\ \bigg\},$
$\qquad\bigcup_{i=1}^{n}\mathcal{A}_{\mathrm{oie}_{\mathrm{idx}_{i}}},$
$\quad{{)}}$
After executing the above-mentioned process, including the process where the result is $\mathrm{oie}_{\mathrm{void}}$, it is called “The complete sequential addition of members of a finite $\mathrm{OIE}\mathcal{S}$ instance under domain-filtering 2-tuple $(\alpha,\beta)$ in index order $\mathrm{idx}\mathcal{T}$”, denoted as
$\overset{n}{\underset{\begin{subarray}{c}i=1\\ {\oplus|_{\alpha}^{\beta}}\end{subarray}}{\sum}}\mathrm{oie}_{\mathrm{idx}_{i}}$
or
${\oplus|_{\alpha}^{\beta}}(\mathrm{oie}_{\mathrm{idx}_{1}},\ \mathrm{oie}_{\mathrm{idx}_{2}},\ \cdots,\ \mathrm{oie}_{\mathrm{idx}_{n}})$
The meanings of the four steps of the complete sequential addition are as follows.
Step 1: “ Calculate the 1st and 4th elements of the result, excluding cases where the same $\textsf{A}\mathrm{Event}^{*}$ instances are involved or $\mathrm{oie}_{\mathrm{void}}$ is involved”.
Cases where the same $\textsf{A}\mathrm{Event}^{*}$ instances are involved in operations leads to duplicate planning of real-world events, are invalid under our operation rule. Additionally, if any operand is $\mathrm{oie}_{\mathrm{void}}$, this means $\emptyset$ will participate in the Cartesian product operation of the sets- and the result will still be $\emptyset$. This will directly result in the complete sequential addition being $\mathrm{oie}_{\mathrm{void}}$ under our operation rule. Intuitively, operating of an entity that cannot plan real-world events with other plan-capable entities renders the integrated plan invalid. Therefore, both in terms of operational logic and practical requirements, we directly stipulate the results of the complete sequential additions in these two cases are void.
For other valid conditions, additional constraints will apply in subsequent steps, if triggered, the operation’s result may be instead set to $\mathrm{oie}_{\mathrm{void}}$. Thus, after temporarily assigning the $\mathcal{C}$(the 1st element) and $\mathcal{A}$(the 4th element) of $\mathrm{oie}_{result}$, proceed to Step 2.
Step 2: “Obtain all feasible combinations of the set that is naturally isomorphic to the Cartesian product”. This is achieved by performing the Cartesian product, getting the set that is naturally isomorphic to it, and then selecting the feasible subset. For detailed procedures, refer to function $\mathfrak{f}\mathrm{FeasInt2tuple}\mathcal{S}$ in Definition 16. If the feasible subset is $\emptyset$, it indicates that there are no possible intervals for event to execute in the real world, then the result of the operation is directly determined as $\mathrm{oie}_{\mathrm{void}}$.
Step 3: “Filter all feasible combinations within the domain”. The filtering rule is characterized by defining a range and selecting all combinations that falls within this range, with some combinations necessarily adhering to the left and right boundaries of the defined range. If the subset after filtering is $\emptyset$, it means that there is no subset that can fulfill the requirements of the complete sequential addition, then the operation result $\mathrm{oie}_{result}$ is directly set to an $\mathrm{OIE}_{\mathrm{void}}$ instance. Otherwise, the $\mathcal{F}$(the 2nd element) of $\mathrm{oie}_{result}$ is assigned to the subset after filtering.
Step 4: “Summarize the feature set of the subset remaining after the filtering in the previous steps”. This feature set is precisely the $\mathcal{I}$(the 3rd element) of $\mathrm{oie}_{result}$. It’s a set of 2-tuples, stored the all possible intervals for the event mapped by the result of complete sequential addition. This step implements Property 5
Intuitively, this is a certain level of “equal opportunity”.
Imagine a 100 metres race, all the athletes have the opportunity to start running after the effective natural reaction time following the gunshot. Although within the scope of classical physics, the physical phenomenon of “all athletes starting at exactly the same moment” has not been observed and recorded to date. Technically, as of 2025, humans are still unable to perfectly observe a single moment without any error. Even sophisticated instruments have errors, and it is almost completely impossible for several athletes, relying on their ears to hear the sound and without any coordination among themselves, to start running at exactly the same moment. However, they all have this opportunity, whether they can achieve it depends on themselves. This achieves a certain level of opportunity fairness.
4.4 Complete sequential multiplication
The addition operation proposed in subsection 4.3 is characterized by the fact that after screening out valid data for each operand participating in the operation, each participant has an interval, with the opportunity to start at the same moment and also has an interval to end at the same moment. The emphasis of the complete sequential addition is on “Fairness”, as all participants have the opportunity.
In this subsection, we propose another type of sequential operation. Its concept is contrary to the fairness of the complete sequential addition, from a certain perspective, it emphasizes “unfairness”. The operation rules directly constrain the chronological order of the starting timestamp and ending timestamp of all participants. After the sequential operation, a certain event will always occur later than another event. We call this “Complete Sequential Multiplication”.
Before giving the full definition, let’s first present some mathematical helper functions.
Definition 21 (Ascending ordered filtered subset of the set of tuples of 2-tuples & helper function & Need not satisfy permutation isomorphism invariance).
1) definition:
Consider a non-empty finite $\mathrm{2Tuple}\mathcal{TS}$ instance $\mathrm{2tuple}\mathcal{TS}$ with length $n>0$, with form
$\mathrm{2tuple}\mathcal{TS}=\{\ \mathrm{2tuple}\mathcal{T}_{1},\ \mathrm{2tuple}\mathcal{T}_{2},\ \cdots,\ \mathrm{2tuple}\mathcal{T}_{n}\ \}$
Then the set
$\mathrm{2tuple}\mathcal{TS}_{asc}=\left\{\ \mathrm{2tuple}\mathcal{T}\in\mathrm{2tuple}\mathcal{TS}\Biggr|\noindent\hbox{}{{\hbox{$\begin{array}[c]{@{}l@{}}\forall\ \mathrm{2tuple}_{i},\ \mathrm{2tuple}_{j}\in\mathrm{2tuple}\mathcal{T}\ \land\ i<j:\\ \ \ \ the\ 2nd\ item\ of\ \mathrm{2tuple}_{i}\leq the\ 1st\ item\ of\ \mathrm{2tuple}_{j}\end{array}$}}}\hfill\hbox{}\right\}$
is called “The ascending ordered filtered subset of $\mathrm{2tuple}\mathcal{TS}$”, denoted as $\mathrm{AscOrderFlt}\mathrm{2Tuple}\mathcal{TS}$.
2) helper function:
$\mathfrak{f}\mathrm{AscOrderFlt2tupleTS}(\mathfrak{p}\mathrm{2tuple}\mathcal{TS})\rightarrow$
$\qquad The\ \mathrm{AscOrderFlt}\mathrm{2Tuple}\mathcal{TS}\ instance\ of\ \mathfrak{p}\mathrm{2tuple}\mathcal{TS}$
$\mathfrak{p}\mathrm{2tuple}\mathcal{TS}$ A $\mathrm{2Tuple}\mathcal{TS}$ instance
Return The $\mathrm{AscOrderFlt}\mathrm{2Tuple}\mathcal{TS}$ instance of $\mathfrak{p}\mathrm{2tuple}\mathcal{TS}$
3) need not satisfy permutation isomorphism invariance:
If there are two $\mathrm{2Tuple}\mathcal{TS}$ instances $\mathrm{2tuple}\mathcal{TS}_{1}$ and $\mathrm{2tuple}\mathcal{TS}_{2}$ with the same length $n$ and satisfying isomorphic via permutation matrix $\mathcal{M}$, then the two return value of this function need not satisfy isomorphic via permutation matrix $\mathcal{M}$.
Proof of need not satisfy permutation isomorphism invariance in Definition 21.
Suppose there are two non-empty $\mathrm{2Tuple}\mathcal{TS}$ instances $\mathrm{2tuple}\mathcal{TS}_{1}$ and $\mathrm{2tuple}\mathcal{TS}_{2}$ with the same length $n$ and satisfying isomorphic via permutation matrix $\mathcal{M}$
$\mathrm{2tuple}\mathcal{TS}_{1}\stackrel{{\scriptstyle\mathcal{M}}}{{\cong}}\mathrm{2tuple}\mathcal{TS}_{2}$
This means that there is a bijection
$\exists\ (\phi:\mathrm{2tuple}\mathcal{TS}_{1}\rightarrow\mathrm{2tuple}\mathcal{TS}_{2})\ :$
$\qquad\forall\mathrm{2tuple}\mathcal{T}\in\mathrm{2tuple}\mathcal{TS}_{1}:\phi(\mathrm{2tuple}\mathcal{T})\in\mathrm{2tuple}\mathcal{TS}_{2}$
where $\phi(\mathrm{2tuple}\mathcal{T})=\mathrm{2tuple}\mathcal{T}\cdot\mathcal{M}$.
(1) The condition of $\mathrm{2tuple}\mathcal{TS}_{1}=\mathrm{2tuple}\mathcal{TS}_{2}$
In this situation, the permutation matrix $\mathcal{M}$ is an identity permutation matrix. Then we use an ascending ordered filter to get a subset from same $\mathrm{2Tuple}\mathcal{TS}$ instance. They must be the same. Thus, this condition is satisfied.
(2) The condition of $\mathrm{2tuple}\mathcal{TS}_{1}\neq\mathrm{2tuple}\mathcal{TS}_{2}$
In this situation, the permutation matrix $\mathcal{M}$ isn’t an identity permutation matrix. Thus
$\mathrm{2tuple}\mathcal{TS}_{1}\stackrel{{\scriptstyle\mathcal{M}}}{{\cong}}\mathrm{2tuple}\mathcal{TS}_{2}$
means that for any $\mathrm{2Tuple}\mathcal{T}$ pairs in a bijection
$\mathrm{2tuple}\mathcal{T}_{\lambda}\in\mathrm{2tuple}\mathcal{TS}_{1},\mathrm{2tuple}\mathcal{T}_{\mu}\in\mathrm{2tuple}\mathcal{TS}_{2}$
$\land$ $\mathrm{2tuple}\mathcal{T}_{\lambda}\stackrel{{\scriptstyle\mathcal{M}}}{{\sim}}\mathrm{2tuple}\mathcal{T}_{\mu}$
there must be two 2-tuples
$(x_{i},y_{i}),(x_{j},y_{j})\in\mathrm{2tuple}\mathcal{T}_{\lambda}$
$\land$ $(x_{i},y_{i}),(x_{j},y_{j})\in\mathrm{2tuple}\mathcal{T}_{\mu}$
$\land$ $x_{i}<y_{i},x_{j}<y_{j}$
their orders in $\mathrm{2tuple}\mathcal{T}_{\lambda}$ and $\mathrm{2tuple}\mathcal{T}_{\mu}$ are swapped, with form
$\mathrm{2tuple}\mathcal{T}_{\lambda}=(\ \cdots,\ (x_{i},y_{i}),\ \cdots,\ (x_{j},y_{j}),\ \cdots\ )$
$\mathrm{2tuple}\mathcal{T}_{\mu}=(\ \cdots,\ (x_{j},y_{j}),\ \cdots,\ (x_{i},y_{i}),\ \cdots\ )$
If
$x_{i}<y_{i}\leq x_{j}<y_{j}$
then according to Definition 21, $\mathrm{2tuple}\mathcal{T}_{\lambda}$ satisfies the rule, so it remains in
$\mathfrak{f}\mathrm{AscOrderFlt2tupleTS}(\mathrm{2tuple}\mathcal{TS}_{1})$. But $\mathrm{2tuple}\mathcal{T}_{\mu}$ doesn’t satisfy the rule, so it doesn’t remain in $\mathfrak{f}\mathrm{AscOrderFlt2tupleTS}(\mathrm{2tuple}\mathcal{TS}_{2})$. Not even a single instance where $(x_{j},y_{j})$
precedes $(x_{i},y_{i})$ can exist in $\mathfrak{f}\mathrm{AscOrderFlt2tupleTS}(\mathrm{2tuple}\mathcal{TS}_{2})$; thus, no element in that set is permutationally equivalent to $\mathrm{2tuple}\mathcal{T}_{\lambda}$.
Otherwise if
$x_{j}<y_{j}\leq x_{i}<y_{i}$
This is consistent with the analytical approach just employed.
We can conclude that in this situation, $\mathrm{2tuple}\mathcal{TS}_{1}$ and $\mathrm{2tuple}\mathcal{TS}_{2}$ don’t satisfy isomorphism via permutation matrix $\mathcal{M}$.
Combining the conclusions of (1) and (2), we complete the proof.
Now, we propose another sequential operation for $\mathrm{OIE}\mathcal{S}$, the elements participating in the operation are all the elements within an $\mathrm{OIE}\mathcal{S}$ instance, and the result of the operation is a new $\mathrm{OIE}$ instance.
Definition 22 (Complete sequential multiplication of members of a finite $\mathrm{OIE}\mathcal{S}$ instance in an index order).
There is a finite $\mathrm{OIE}\mathcal{S}$ instance $\mathrm{oie}\mathcal{S}$ containing $n$ ($n\neq\infty$) elements in the ordered enumeration form
$\mathrm{oie}\mathcal{S}=\{\mathrm{oie}_{1},\mathrm{oie}_{2},\mathrm{oie}_{3},\cdots,\mathrm{oie}_{n}\}$
, an $\mathrm{Idx}\mathcal{T}$ instance with length $n$
$\mathrm{idx}\mathcal{T}=(\mathrm{idx}_{1},\mathrm{idx}_{2},\mathrm{idx}_{3},\cdots,\mathrm{idx}_{n})$
And for $\mathrm{oie}\mathcal{S}$ and $\mathrm{Idx}\mathcal{T}$ instance
$\mathrm{idx}\mathcal{T}_{asc}=(1,\ 2,\ 3,\ \cdots,\ n)$
The $\mathrm{2Tuple}\mathcal{TS}^{Inf}$ instance of $\mathrm{oie}\mathcal{S}$ in the index order $\mathrm{idx}\mathcal{T}_{asc}$ is
$\mathrm{2tuple}\mathcal{TS}^{inf}_{\mathrm{oie}\mathcal{S}}=\mathfrak{f}\mathrm{Inf2tupleTS}(\mathrm{oie}\mathcal{S},\mathrm{idx}\mathcal{T}_{asc})$
Replace finitary operation $\otimes$, its operational rules are as follows:
Step 1. If
$\bigcap_{i=1}^{n}\mathcal{A}_{\mathrm{oie}_{\mathrm{idx}_{i}}}\neq\emptyset$
$\lor$ $\exists\ k\ \in\ [1,n]:\ \mathrm{oie}_{\mathrm{idx}_{k}}=\mathrm{oie}_{\mathrm{void}}$
then, the result is directly set to the $\mathrm{OIE}_{\mathrm{void}}$ instance, with form
$\mathrm{oie}_{\mathrm{void}}={{(}}\ (),\ \emptyset,\ \emptyset,\ \emptyset\ {{)}}$
and the operation will terminate. Otherwise, initialize a new $\mathrm{OIE}$ instance $\mathrm{oie}_{result}$ as the result. Temporarily set the $\mathcal{C}$(1st element) of $\mathrm{oie}_{result}$ to tuple
$\mathcal{C}_{\mathrm{oie}_{result}}=(\mathrm{oie}_{\mathrm{idx}_{1}},\ \mathrm{oie}_{\mathrm{idx}_{2}},\ \mathrm{oie}_{\mathrm{idx}_{3}},\ \cdots,\ \mathrm{oie}_{\mathrm{idx}_{n}})$
, and temporarily set the $\mathcal{A}$(4th element) of $\mathrm{oie}_{result}$ to set
$\mathcal{A}_{\mathrm{oie}_{result}}=\bigcup_{i=1}^{n}\mathcal{A}_{\mathrm{oie}_{\mathrm{idx}_{i}}}$
then proceed to Step 2.
Step 2. Using Definition 16, take the feasible subset of the set that is naturally isomorphic to the Cartesian product of all members of the $\mathrm{2Tuple}\mathcal{SS}^{Intvl}$ instance of $\mathrm{oie}\mathcal{S}$ in the index order $\mathrm{idx}\mathcal{T}$, denoted as $\mathrm{2tuple}\mathcal{TS}^{feas}$
$\mathrm{2tuple}\mathcal{TS}^{feas}=$
$\quad\mathfrak{f}\mathrm{Feas2tuple}\mathcal{TS}(\mathrm{oie}\mathcal{S},\mathrm{idx}\mathcal{T})$
If
$\mathrm{2tuple}\mathcal{TS}^{feas}=\emptyset$
then $\mathrm{oie}_{result}$ is instead set to $\mathrm{oie}_{\mathrm{void}}$, with form
$\mathrm{oie}_{\mathrm{void}}={{(}}\ (),\ \emptyset,\ \emptyset,\ \emptyset\ {{)}}$
and the operation ends without performing the subsequent steps. Otherwise, proceed to Step 3.
Step 3. Using Definition 21, get the corresponding $\mathrm{AscOrderFlt}\mathrm{2Tuple}\mathcal{TS}$ instance from $\mathrm{2tuple}\mathcal{TS}^{feas}$, and use $ascOrderFlt\mathrm{2tuple}\mathcal{TS}$ to represent it
$ascOrderFlt\mathrm{2tuple}\mathcal{TS}$ $=\mathfrak{f}\mathrm{AscOrderFlt2tupleTS}(\mathrm{2tuple}\mathcal{TS}^{feas})$
$=\mathfrak{f}\mathrm{AscOrderFlt2tupleTS}(\mathfrak{f}\mathrm{Feas2tuple}\mathcal{TS}(\mathrm{oie}\mathcal{S},\mathrm{idx}\mathcal{T}))$
with form is
$ascOrderFlt\mathrm{2tuple}\mathcal{TS}=\{\mathrm{2tuple}\mathcal{T}_{1},\mathrm{2tuple}\mathcal{T}_{2},\cdots,\mathrm{2tuple}\mathcal{T}_{\lambda}\}$
If
$ascOrderFlt\mathrm{2tuple}\mathcal{TS}\neq\emptyset$
then the $\mathcal{F}$(2nd element) of $\mathrm{oie}_{result}$ is assigned to $ascOrderFlt\mathrm{2tuple}\mathcal{TS}$
$\mathcal{F}_{\mathrm{oie}_{result}}=ascOrderFlt\mathrm{2tuple}\mathcal{TS}$
and then Step 4 is executed. Otherwise, if
$ascOrderFlt\mathrm{2tuple}\mathcal{TS}=\emptyset$
$\mathrm{oie}_{result}$ is instead set to $\mathrm{oie}_{\mathrm{void}}$, with form
$\mathrm{oie}_{\mathrm{void}}={{(}}\ (),\ \emptyset,\ \emptyset,\ \emptyset\ {{)}}$
And the operation ends without performing the subsequent steps.
Step 4. Using Definition 5, get the $\mathcal{I}$(3rd element) of $\mathrm{oie}_{result}$
$\mathcal{I}_{\mathrm{oie}_{result}}$ $=\bigg\{\ \mathfrak{f}\mathrm{Bound2tuple}(\mathrm{2tuple}\mathcal{T}_{i})\ \bigg|\mathrm{2tuple}\mathcal{T}_{i}\in ascOrderFlt\mathrm{2tuple}\mathcal{TS}\ \bigg\}$
$=\bigg\{\ \mathfrak{f}\mathrm{Bound2tuple}(\mathrm{2tuple}\mathcal{T}_{i})\ \bigg|\$
$\qquad\quad\mathrm{2tuple}\mathcal{T}_{i}\in\mathfrak{f}\mathrm{AscOrderFlt2tupleTS}(\mathfrak{f}\mathrm{Feas2tuple}\mathcal{TS}(\mathrm{oie}\mathcal{S},\mathrm{idx}\mathcal{T}))\ \bigg\}$
At this juncture, the operation is completed.
Finally, the form of $\mathrm{oie}_{result}$ is
$\mathrm{oie}_{result}={{(}}\ (),\ \emptyset,\ \emptyset,\ \emptyset\ {{)}}$
or
$\mathrm{oie}_{result}=$
$\quad{{(}}$
$\qquad(\mathrm{oie}_{\mathrm{idx}_{1}},\ \mathrm{oie}_{\mathrm{idx}_{2}},\ \mathrm{oie}_{\mathrm{idx}_{3}},\ \cdots,\ \mathrm{oie}_{\mathrm{idx}_{n}}),$
$\qquad\mathfrak{f}\mathrm{AscOrderFlt2tupleTS}(\mathfrak{f}\mathrm{Feas2tuple}\mathcal{TS}(\mathrm{oie}\mathcal{S},\mathrm{idx}\mathcal{T})),$
$\qquad\bigg\{\ \mathfrak{f}\mathrm{Bound2tuple}(\mathrm{2tuple}\mathcal{T}_{i})\ \bigg|\$
$\qquad\qquad\mathrm{2tuple}\mathcal{T}_{i}\in\mathfrak{f}\mathrm{AscOrderFlt2tupleTS}(\mathfrak{f}\mathrm{Feas2tuple}\mathcal{TS}(\mathrm{oie}\mathcal{S},\mathrm{idx}\mathcal{T}))\ \bigg\},$
$\qquad\bigcup_{i=1}^{n}\mathcal{A}_{\mathrm{oie}_{\mathrm{idx}_{i}}},$
$\quad{{)}}$
After the execution of the above-mentioned process, including the process where the result is $\mathrm{oie}_{\mathrm{void}}$, it is called the complete sequential multiplication of members of a finite $\mathrm{OIE}\mathcal{S}$ in the index order $\mathrm{idx}\mathcal{T}$, denoted as
$\overset{n}{\underset{\begin{subarray}{c}i=1\\ {\otimes}\end{subarray}}{\prod}}\mathrm{oie}_{\mathrm{idx}_{i}}$
or
${\otimes}(\mathrm{oie}_{\mathrm{idx}_{1}},\ \mathrm{oie}_{\mathrm{idx}_{2}},\ \cdots,\ \mathrm{oie}_{\mathrm{idx}_{n}})$
Similar to the addition operation (Definition 20), the condition for successfully completing this type of operation (where the result is not an $\mathrm{OIE}_{\mathrm{void}}$ instance) is that the specified condition restrictions must be met in Step 1, Step 2 and Step 3.
The significance of the restriction in Step 1 and Step 2 is same as Step 1 and Step 2 of complete sequential addition.
The significance of the restriction in Step 3 is that after the screening, there are valid optional intervals left. But different from the complete sequential addition, for each element in the $\mathcal{F}$(2nd element) of the $\mathrm{oie}_{result}$, each interval segment follows the order of $\mathrm{idx}\mathcal{T}$, and the starting timestamp of the subsequent interval segment is later than or equal to the ending timestamp of the previous interval segment. That is to say, the characteristic of multiplication lies in sorting, rather than emphasizing fairness within a domain like addition.
Intuitively, this is a kind of unfairness in terms of sequence. In the Winter Olympics, there is a sport called downhill skiing. It requires athletes to compete alone on the track in a certain order, and those with shorter times are ranked higher. This sport is one of the few racing competitions where athletes have exclusive use of the track. The order of who goes first and who goes later is not random, it is determined by the preliminary round results. Skiers with better preliminary round results compete later. This order represents a form of unfairness, but before the finals, each skier already knows their starting position.
In a downhill skiing race, athletes start the competition after the referee gives the command, and it is also a racing competition. Let $n$ skiers participate in the final, and $Skier_{i}$ is the $\mathrm{OIE}$ instance corresponding to the event that the $i$-th skier participates in the final.
The best skiing time of the $i$-th skier is $Tmin_{i}$, and the slowest skiing time is $Tmax_{i}$. The starting timestamp of the first skier in the final (that is, the start time of the whole competition) is $Ts_{\alpha}$, and the latest ending timestamp of the last skier (that is, the time when the whole competition must end) is $Ts_{\beta}$.
Therefore, before the results of the preliminary round are announced and when the starting order of the $i$-th skier in the final is uncertain, the 3rd item of $Skier_{i}$ is
$\left\{(\mathrm{ts}_{i},\mathrm{te}_{i})\,\middle|\,\begin{aligned} &\mathrm{te}_{i}-\mathrm{ts}_{i}\geq Tmin_{i}\\ \land\ &\mathrm{te}_{i}-\mathrm{ts}_{i}\leq Tmax_{i}\\ \land\ &\mathrm{ts}_{i}\geq Ts_{\alpha}\\ \land\ &\mathrm{te}_{i}\leq Ts_{\beta}\\ \end{aligned}\right\}$
When the preliminary round is over, the order of each contestant in the final is also determined. Use
$\mathrm{idx}\mathcal{T}=(\mathrm{idx}_{1},\mathrm{idx}_{2},\mathrm{idx}_{3},\cdots,\mathrm{idx}_{n})$
to represent the starting order. Then, according to the Definition 22, this final can be expressed as
$\mathrm{oie}_{final}=\overset{n}{\underset{\begin{subarray}{c}i=1\\ {\otimes}\end{subarray}}{\prod}}Skier_{\mathrm{idx}_{i}}$
The 2nd item of $\mathrm{oie}_{final}$ contains all possible combinations of the scores of all skiers. Skiers in different combinations have different exact appearance timestamp, while their order of appearance in the final is same. Finally only one combination happened in reality.
4.5 Void result of sequential operations
According to Definition 20 and Definition 22, in “Complete Sequential Addition” and “Complete Sequential Multiplication” of $\mathrm{OIE}\mathcal{S}$, if the specified conditions are not met at certain steps during operations, the result is the only instance of $\mathrm{OIE}_{\mathrm{void}}$, with form
$\mathrm{oie}_{\mathrm{void}}={{(}}\ (),\ \emptyset,\ \emptyset,\ \emptyset\ {{)}}$
In this subsection, we will discuss several typical cases where the result of sequential operations is this special instance.
Property 11 (Sequential operations involving identical $\mathrm{OIE}$ instances result in an $\mathrm{OIE}_{\mathrm{void}}$ instance).
Property 11 serves as a more intuitive scenario for the condition
$\bigcap_{i=1}^{n}\mathcal{A}_{\mathrm{oie}_{\mathrm{idx}_{i}}}\neq\emptyset$
which triggers $\mathrm{oie}_{\mathrm{void}}$ result in Step 1 of both complete sequential addition and multiplication. From a philosophical perspective, an Event instance event has its own entity, an entity cannot have an identical copy existing in the same status within the same time interval. For example, an athlete cannot simultaneously compete against himself in a sprint race on two different tracks. And the sentence $Dr_{A}$ and $Dr_{A}$ submitted a paper during a certain interval” is also an obvious ungrammatical statement. We believe that $\mathrm{OIE}$ also has the similar property.
To extend Property 11: Once an event instance becomes a component of two other complex event instance, it is also irrational to repeatedly plan these two complex event. The check for whether there is an intersection of the $\mathcal{A}$ of $\mathrm{OIE}$ instances in Step1 of two sequential operations is a formalized expression of this line of thought.
Property 12 (The sequential operations involving $\mathrm{OIE}_{\mathrm{void}}$ instances result $\mathrm{oie}_{\mathrm{void}}$).
When an $\mathrm{OIE}_{\mathrm{void}}$ instance $\mathrm{oie}_{\mathrm{void}}$ participates in the two types of sequential operations proposed in this paper, its $\mathcal{I}$(3rd item) will perform the Cartesian product operation with other $\mathrm{OIE}$ instances’ $\mathcal{I}$. Cause it is $\emptyset$, so the Cartesian product result is still $\emptyset$. No matter what filter is applied to $\emptyset$, the result is always $\emptyset$. Therefore, regardless of the type of sequential operators used, the $\mathcal{F}$(2nd item) of result will be $\emptyset$, which in turn causes the $\mathcal{F}$(3rd item) of the result to also be $\emptyset$. Consequently, the final result will inevitably be an $\mathrm{OIE}_{\mathrm{void}}$ instance.
4.6 Permutationally equivalent relation of $\mathrm{OIE}$ instances and results of sequential operations
In this subsection, we discuss the permutationally equivalent relation of $\mathrm{OIE}$ instances.
4.6.1 Permutationally equivalent relation of $\mathrm{OIE}$ instances
Definition 23 (Permutationally equivalent relationship of Non-$\mathrm{oie}_{\mathrm{void}}$ $\mathrm{OIE}$ instances).
Suppose there are two non-$\mathrm{oie}_{\mathrm{void}}$ $\mathrm{OIE}$ instances $\mathrm{oie}_{1}$ and $\mathrm{oie}_{2}$.
If there exists a permutation matrix $\mathcal{M}$ satisfying that
$\mathcal{C}_{\mathrm{oie}_{1}}\cdot\mathcal{M}=\mathcal{C}_{\mathrm{oie}_{2}},$
$\land$ $\mathcal{F}_{\mathrm{oie}_{1}}\stackrel{{\scriptstyle\mathcal{M}}}{{\cong}}\mathcal{F}_{\mathrm{oie}_{2}},$
$\quad(The2nd\ element\ of\ \mathrm{oie}_{1}\ and\ the2nd\ element$
$\quad of\ \mathrm{oie}_{2}\ are\ isomorphic$
$\quad via\ the\ permutation\ matrix\ \mathcal{M})$
$\land$ $\mathcal{I}_{\mathrm{oie}_{1}}=\mathcal{I}_{\mathrm{oie}_{2}},$
$\land$ $\mathcal{A}_{\mathrm{oie}_{1}}=\mathcal{A}_{\mathrm{oie}_{2}}$
then we say that $\mathrm{oie}_{1}$ and $\mathrm{oie}_{2}$ are permutationally equivalent with respect to $\mathcal{M}$”, denoted as
$\mathrm{oie}_{1}\stackrel{{\scriptstyle\mathcal{M}}}{{\sim}}\mathrm{oie}_{2}$
Here we need to explain the practical significance of Definition 23.
$\mathcal{F}_{\mathrm{oie}_{1}}\cong_{\mathcal{M}}\mathcal{F}_{\mathrm{oie}_{2}}$
means that
$\forall\mathrm{2tuple}\mathcal{T}_{i}\in\mathcal{F}_{\mathrm{oie}_{1}},\ \exists\mathrm{2tuple}\mathcal{T}_{j}\in\mathcal{F}_{\mathrm{oie}_{2}}:\mathrm{2tuple}\mathcal{T}_{i}\stackrel{{\scriptstyle\mathcal{M}}}{{\sim}}\mathrm{2tuple}\mathcal{T}_{j}$
$\land$ $\mathcal{F}_{\mathrm{oie}_{1}}\xrightarrow{\sim}\mathcal{F}_{\mathrm{oie}_{2}}\ form\ a\ bijection$
It is a very common format to describe N events as follows:
$\qquad The\ first\ event\ is\ like\ this\ and\ that,\ the\ second\ event\ is\ like\ this\ and\ that$.
$The\ Nth\ event\ is\ like\ this\ and\ that.$
In such way, there must be a 1st one presented first, a 2nd one that appears afterward, and a last one appears at the end. But in reality, the order in which these N events are presented does not affect their inherent nature.
In the corresponding $\mathrm{OIE}\mathcal{S}$, this means that the difference between two $\mathrm{OIE}$ instances satisfy the permutationally equivalent relation, is “The two instances differ in expression form due to the different presentation orders of their meta $\mathrm{OIE}$ instances”. They can be transformed into each other’s expression from simply by performing a permutation operation. When it comes to the real-world events mapped by each $\mathrm{OIE}$ instance in the $\mathcal{C}$, the intervals and combinations of there real-world events are not affected by the permutation operation.
If the $\mathcal{C}$(1st element) of an $\mathrm{OIE}$ instance has only one element, this $\mathrm{OIE}$ instance is $\textsf{Atom}\mathrm{OIE}$ type, it has a special permutationally equivalent relation. Then for the $\textsf{Atom}\mathrm{OIE}$ instances, we summarize the following property.
Property 13 (Permutationally equivalent relationship of $\textsf{Atom}\mathrm{OIE}$ instances implies their equality).
Suppose there are two $\textsf{Atom}\mathrm{OIE}$ instances $\textsf{atom}\mathrm{Oie}_{1}$ and $\textsf{atom}\mathrm{Oie}_{2}$. If they are permutationally equivalent with respect to $\mathcal{M_{I}}$
$\textsf{atom}\mathrm{Oie}_{1}\stackrel{{\scriptstyle\mathcal{M_{I}}}}{{\sim}}\textsf{atom}\mathrm{Oie}_{2}$
Then
$\textsf{atom}\mathrm{Oie}_{1}=\textsf{atom}\mathrm{Oie}_{2},$
$\land$ $\mathcal{M_{I}}\ is\ an\ identity\ permutation\ matrix$
For $\mathrm{OIE}_{\mathrm{void}}$, it only has one instance, with form
$\mathrm{oie}_{\mathrm{void}}={{(}}\ (),\ \emptyset,\ \emptyset,\ \emptyset\ {{)}}$
we hold that it is permutationally equivalent to itself.
Property 14 (Permutationally equivalent relationship of $\mathrm{OIE}_{\mathrm{void}}$ instance to itself).
4.6.2 Permutationally equivalent relation of results of complete sequential addition
In this subsubsection, we will prove that there exists a permutationally equivalent relation between the results of the same operands arranged in different orders under complete sequential addition.
Property 15 (the permutation equivalent relation between the results of the same operands arranged in different orders under complete sequential addition.).
Suppose there are a non-empty finite $\mathrm{OIE}\mathcal{S}$ instance
$\mathrm{oie}\mathcal{S}=\{\mathrm{oie}_{1},\ \mathrm{oie}_{2},\ \cdots,\ \mathrm{oie}_{n}\}$
two $\mathrm{Idx}\mathcal{T}$ instances with length $n$
$\mathrm{idx}\mathcal{T}_{1}=\{i_{1},\ i_{2},\ i_{3},\ \cdots,\ i_{n}\}$
$\mathrm{idx}\mathcal{T}_{2}=\{j_{1},\ j_{2},\ j_{3},\ \cdots,\ j_{n}\}$
and a $\mathrm{DomFlt2Tuple}$ instance $(\alpha,\beta)$.
Do two complete sequential additions
$\mathrm{oie}_{i}=\overset{n}{\underset{\begin{subarray}{c}k=1\\ {\oplus|_{\alpha}^{\beta}}\end{subarray}}{\sum}}\mathrm{oie}_{i_{k}},$
$\mathrm{oie}_{j}=\overset{n}{\underset{\begin{subarray}{c}k=1\\ {\oplus|_{\alpha}^{\beta}}\end{subarray}}{\sum}}\mathrm{oie}_{j_{k}}$
then there exists a permutation matrix $\mathcal{M}$ such that $\mathrm{oie}_{i}$ and $\mathrm{oie}_{j}$ are permutationally equivalent via $\mathcal{M}$, denoted as
$\mathrm{oie}_{i}\stackrel{{\scriptstyle\mathcal{M}}}{{\sim}}\mathrm{oie}_{j}$
Proof of the Property 15.
(1) The situation $\mathrm{OIE}_{\mathrm{void}}$ instance exists in $\mathrm{oie}\mathcal{S}$ or the intersection of all $\mathcal{A}$ across all operands isn’t $\emptyset$.
Assume that the $\mathrm{OIE}\mathcal{S}$ instance $\mathrm{oie}\mathcal{S}$ contains an $\mathrm{OIE}_{\mathrm{void}}$ instance or the intersection of the $\mathcal{A}$(4th elements) of all operands is non-empty. Then, by Step 1 of Definition 20, the result of the complete sequential addition for any index order is set to $\mathrm{oie}_{\mathrm{void}}$. Since $\mathrm{oie}_{\mathrm{void}}$ is the unique instance of type $\mathrm{OIE}_{\mathrm{void}}$, both operations yield the same result, namely $\mathrm{oie}_{\mathrm{void}}$. By Property 14, $\mathrm{oie}_{\mathrm{void}}$ is permutationally equivalent via any permutation matrix to itself. Therefore, in this case, the two results of complete sequential addition are permutationally equivalent.
This completes the proof for these cases.
(2) The situation $\mathrm{OIE}_{\mathrm{void}}$ instance doesn’t exist in $\mathrm{oie}\mathcal{S}$ and the intersection of all $\mathcal{A}$ across all operands is $\emptyset$.
separately for each of the four elements of these two $\mathrm{OIE}$ instances.
Step I. Proof for $\mathcal{C}$(1st element) of $\mathrm{oie}_{i}$ and $\mathrm{oie}_{j}$
In Step 1, there are two $\mathcal{C}$(1st element) of $\mathrm{oie}_{i}$ and $\mathrm{oie}_{j}$ respectively. According to Property 8, there must be a permutation matrix $\mathcal{M}$ satisfying that
$\mathrm{idx}\mathcal{T}_{1}\stackrel{{\scriptstyle\mathcal{M}}}{{\sim}}\mathrm{idx}\mathcal{T}_{2}$
and two complete sequential addition operations are exactly the ones that apply the order of these two $\mathrm{Idx}\mathcal{T}$ instances directly, according to Definition 20
$\mathcal{C}_{\mathrm{oie}_{i}}=(\mathrm{oie}_{i_{1}},\ \mathrm{oie}_{i_{2}},\ \mathrm{oie}_{i_{3}},\ \cdots,\ \mathrm{oie}_{i_{n}})$
$\mathcal{C}_{\mathrm{oie}_{j}}=(\mathrm{oie}_{j_{1}},\ \mathrm{oie}_{j_{2}},\ \mathrm{oie}_{j_{3}},\ \cdots,\ \mathrm{oie}_{j_{n}})$
then, the two $\mathcal{C}$(1st element) of $\mathrm{oie}_{i}$ and $\mathrm{oie}_{j}$ must be permutationally equivalent based on $\mathcal{M}$ too, i.e.
$\mathcal{C}_{\mathrm{oie}_{i}}\cdot\mathcal{M}=\mathcal{C}_{\mathrm{oie}_{j}}$
This equation satisfies the first condition in Definition 23, the proof for $\mathcal{C}$ of $\mathrm{OIE}$ is complete.
Step II. Proof for $\mathcal{F}$(2nd element) of $\mathrm{oie}_{i}$ and $\mathrm{oie}_{j}$
In Step 2, According to Property 10
$\mathfrak{f}\mathrm{Feas2tuple}\mathcal{TS}(\mathrm{oie}\mathcal{S},\mathrm{idx}\mathcal{T}_{1})\stackrel{{\scriptstyle\mathcal{M}}}{{\cong}}\mathfrak{f}\mathrm{Feas2tuple}\mathcal{TS}(\mathrm{oie}\mathcal{S},\mathrm{idx}\mathcal{T}_{2})$
If
$\mathfrak{f}\mathrm{Feas2tuple}\mathcal{TS}(\mathrm{oie}\mathcal{S},\mathrm{idx}\mathcal{T}_{1})=\emptyset$
$\land$ $\mathfrak{f}\mathrm{Feas2tuple}\mathcal{TS}(\mathrm{oie}\mathcal{S},\mathrm{idx}\mathcal{T}_{2})=\emptyset$
then according to Definition 20
$\mathrm{oie}_{i}=\mathrm{oie}_{j}=\mathrm{oie}_{\mathrm{void}}$
according to Property 14
$\mathrm{oie}_{i}\stackrel{{\scriptstyle\mathcal{M}}}{{\sim}}\mathrm{oie}_{j}$
this completes the proof for this conditional branch.
Otherwise, they are both not $\emptyset$, process goes to step 3, we use a $\mathrm{DomFlt2Tuple}$ instance $(\alpha,\beta)$ to filter the isomorphic two $\mathrm{2Tuple}\mathcal{TS}^{feas}$ instances, get two $\mathrm{2Tuple}\mathcal{TS}^{feas}$ instances, then assign them to $\mathcal{F}_{\mathrm{oie}_{i}}$ and $\mathcal{F}_{\mathrm{oie}_{j}}$
$\mathcal{F}_{\mathrm{oie}_{i}}=\mathfrak{f}\mathrm{DomFlt}\mathrm{2tuple}\mathcal{TS}(\mathfrak{f}\mathrm{Feas2tuple}\mathcal{TS}(\mathrm{oie}\mathcal{S},\mathrm{idx}\mathcal{T}_{1}),\alpha,\beta),$
$\mathcal{F}_{\mathrm{oie}_{j}}=\mathfrak{f}\mathrm{DomFlt}\mathrm{2tuple}\mathcal{TS}(\mathfrak{f}\mathrm{Feas2tuple}\mathcal{TS}(\mathrm{oie}\mathcal{S},\mathrm{idx}\mathcal{T}_{2}),\alpha,\beta)$
If
$\mathcal{F}_{\mathrm{oie}_{i}}=\emptyset\land\ \mathcal{F}_{\mathrm{oie}_{j}}=\emptyset$
then according to Definition 20
$\mathrm{oie}_{i}=\mathrm{oie}_{j}=\mathrm{oie}_{\mathrm{void}}$
according to Property 14
$\mathrm{oie}_{i}\stackrel{{\scriptstyle\mathcal{M}}}{{\sim}}\mathrm{oie}_{j}$
this completes the proof for this conditional branch.
Otherwise, both $\mathcal{F}_{\mathrm{oie}_{i}}$ and $\mathcal{F}_{\mathrm{oie}_{j}}$ are not $\emptyset$, according to the permutation isomorphism invariance in Definition 19, it can be concluded that
$\mathfrak{f}\mathrm{Feas2tuple}\mathcal{TS}(\mathrm{oie}\mathcal{S},\mathrm{idx}\mathcal{T}_{1})\stackrel{{\scriptstyle\mathcal{M}}}{{\cong}}\mathfrak{f}\mathrm{Feas2tuple}\mathcal{TS}(\mathrm{oie}\mathcal{S},\mathrm{idx}\mathcal{T}_{2})\rightarrow$
$\qquad\mathcal{F}_{\mathrm{oie}_{i}}\stackrel{{\scriptstyle\mathcal{M}}}{{\cong}}\mathcal{F}_{\mathrm{oie}_{j}}$
The proof for $\mathcal{F}$(2nd element) of $\mathrm{OIE}$ is complete.
Step III. Proof for $\mathcal{I}$(3rd element) of $\mathrm{oie}_{i}$ and $\mathrm{oie}_{j}$
The form of $\mathcal{I}$(3rd element) of $\mathrm{oie}_{i}$ and $\mathrm{oie}_{j}$ are as following
$\mathcal{I}_{\mathrm{oie}_{i}}=\bigg\{\ \mathfrak{f}\mathrm{Bound2tuple}(\mathrm{2tuple}\mathcal{T})\ \bigg|\mathrm{2tuple}\mathcal{T}\in\mathcal{F}_{\mathrm{oie}_{i}}\ \bigg\},$
$\mathcal{I}_{\mathrm{oie}_{j}}=\bigg\{\ \mathfrak{f}\mathrm{Bound2tuple}(\mathrm{2tuple}\mathcal{T})\ \bigg|\mathrm{2tuple}\mathcal{T}\in\mathcal{F}_{\mathrm{oie}_{j}}\ \bigg\}$
We have already proved that $\mathcal{F}_{\mathrm{oie}_{i}}$ and $\mathcal{F}_{\mathrm{oie}_{j}}$ are isomorphic via a permutation matrix $\mathcal{M}$, thus, there exists a bijection of permutational equivalence between their members. That means
$\forall\mathrm{2tuple}\mathcal{T}_{\lambda}\in\mathcal{F}_{\mathrm{oie}_{i}},\ \exists\mathrm{2tuple}\mathcal{T}_{\mu}\in\mathcal{F}_{\mathrm{oie}_{j}}:$
$\qquad\mathrm{2tuple}\mathcal{T}_{\lambda}\stackrel{{\scriptstyle\mathcal{M}}}{{\sim}}\mathrm{2tuple}\mathcal{T}_{\mu}$
and this correspondence is bijective.
According to the permutation invariance in Definition 7, for every corresponding pair $\mathrm{2tuple}\mathcal{T}_{\lambda}$ and $\mathrm{2tuple}\mathcal{T}_{\mu}$ of $\mathcal{F}_{\mathrm{oie}_{i}}$ and $\mathcal{F}_{\mathrm{oie}_{j}}$,
$\mathrm{2tuple}\mathcal{T}_{\lambda}\stackrel{{\scriptstyle\mathcal{M}}}{{\sim}}\mathrm{2tuple}\mathcal{T}_{\mu}\rightarrow$
$\qquad\mathfrak{f}\mathrm{Bound2tuple}(\mathrm{2tuple}\mathcal{T}_{\lambda})=\mathfrak{f}\mathrm{Bound2tuple}(\mathrm{2tuple}\mathcal{T}_{\mu})$
Then we can conclude that
$\bigg\{\ \mathfrak{f}\mathrm{Bound2tuple}(\mathrm{2tuple}\mathcal{T})\ \bigg|\mathrm{2tuple}\mathcal{T}\in\mathcal{F}_{\mathrm{oie}_{i}}\ \bigg\}=$
$\qquad\bigg\{\ \mathfrak{f}\mathrm{Bound2tuple}(\mathrm{2tuple}\mathcal{T})\ \bigg|\mathrm{2tuple}\mathcal{T}\in\mathcal{F}_{\mathrm{oie}_{j}}\ \bigg\}$
that is $\mathcal{I}_{\mathrm{oie}_{i}}=\mathcal{I}_{\mathrm{oie}_{j}}$, the proof for $\mathcal{I}$(3rd element) of $\mathrm{OIE}$ is complete.
Step IV. Proof for $\mathcal{A}$(4th element) of $\mathrm{oie}_{i}$ and $\mathrm{oie}_{j}$
In Step 1 of Definition 20, the $\mathcal{A}$(4th element) of the result $\mathrm{OIE}$ instance is set to the union of all $\mathcal{A}$(4th elements) of the operands:
$\mathcal{A}_{\mathrm{oie}_{result}}=\bigcup_{k=1}^{n}\mathcal{A}_{\mathrm{oie}_{k}}$
For the two complete sequential addition operations with different index orders, we have:
$\mathcal{A}_{\mathrm{oie}_{i}}=\bigcup_{k=1}^{n}\mathcal{A}_{\mathrm{oie}_{i_{k}}},\ \mathcal{A}_{\mathrm{oie}_{j}}=\bigcup_{k=1}^{n}\mathcal{A}_{\mathrm{oie}_{j_{k}}}$
For finite $\mathrm{OIE}\mathcal{S}$ instance
$\mathrm{oie}\mathcal{S}=\{\ \mathrm{oie}_{1},\ \mathrm{oie}_{2},\ \cdots,\ \mathrm{oie}_{n}\ \}$
since $\mathrm{idx}\mathcal{T}_{1}=\{i_{1},i_{2},\cdots,i_{n}\}$ and $\mathrm{idx}\mathcal{T}_{2}=\{j_{1},j_{2},\cdots,j_{n}\}$ are permutationally equivalent, multiset $\{\mathrm{oie}_{i_{1}},\mathrm{oie}_{i_{2}},\cdots,\mathrm{oie}_{i_{n}}\}$ and multiset $\{\mathrm{oie}_{j_{1}},\mathrm{oie}_{j_{2}},\cdots,\mathrm{oie}_{j_{n}}\}$ are permutationally equivalent, as they both contain exactly the same elements from $\mathrm{oie}\mathcal{S}$(just may be in different orders). we have:
$\mathcal{A}_{\mathrm{oie}_{i}}=\bigcup_{k=1}^{n}\mathcal{A}_{\mathrm{oie}_{i_{k}}},\ \mathcal{A}_{\mathrm{oie}_{j}}=\bigcup_{k=1}^{n}\mathcal{A}_{\mathrm{oie}_{j_{k}}}$
Since set union is a commutative and idempotent operation that is independent of the order of its operands, we have:
$\bigcup_{k=1}^{n}\mathcal{A}_{\mathrm{oie}_{i_{k}}}=\bigcup_{k=1}^{n}\mathcal{A}_{\mathrm{oie}_{j_{k}}}=\bigcup_{k=1}^{n}\mathcal{A}_{\mathrm{oie}_{k}}$
Therefore, we can conclude that
$\mathcal{A}_{\mathrm{oie}_{i}}=\mathcal{A}_{\mathrm{oie}_{j}}$
the proof for $\mathcal{A}$(4th element) of $\mathrm{OIE}$ is complete.
Now, the proof for case (2) is complete.
We have completed the proofs for cases (1) and (2) above. Since (1) and (2) cover all possible cases, we have completed the proof.
4.6.3 Permutationally equivalent relation of results of complete sequential multiplication
In this subsubsection, we will prove that there need not exist a permutationally equivalent relation between the results of the same operands arranged in different orders under complete sequential addition.
Property 16 (the permutation equivalent relation between the results of the same operands arranged in different orders under complete sequential multiplication.).
Suppose there are a non-empty finite $\mathrm{OIE}\mathcal{S}$ instance
$\mathrm{oie}\mathcal{S}=\{\mathrm{oie}_{1},\ \mathrm{oie}_{2},\ \cdots,\ \mathrm{oie}_{n}\}$
two $\mathrm{Idx}\mathcal{T}$ instances with length $n$
$\mathrm{idx}\mathcal{T}_{1}=\{i_{1},\ i_{2},\ i_{3},\ \cdots,\ i_{n}\}$
$\mathrm{idx}\mathcal{T}_{2}=\{j_{1},\ j_{2},\ j_{3},\ \cdots,\ j_{n}\}$
Do two complete sequential multiplications
$\mathrm{oie}_{i}=\overset{n}{\underset{\begin{subarray}{c}k=1\\ {\otimes}\end{subarray}}{\prod}}\mathrm{oie}_{i_{k}},$
$\mathrm{oie}_{j}=\overset{n}{\underset{\begin{subarray}{c}k=1\\ {\otimes}\end{subarray}}{\prod}}\mathrm{oie}_{j_{k}}$
Then it is not always true that $\mathrm{oie}_{i}$ and $\mathrm{oie}_{j}$ are permutationally equivalent.
Proof of the Property 16.
We proceed by considering two cases.
(1) The situation $\mathrm{OIE}_{\mathrm{void}}$ instance exists in $\mathrm{oie}\mathcal{S}$ or the intersection of all $\mathcal{A}$ across all operands isn’t $\emptyset$.
Assume that the $\mathrm{OIE}\mathcal{S}$ instance $\mathrm{oie}\mathcal{S}$ contains an $\mathrm{OIE}_{\mathrm{void}}$ instance or the intersection of the $\mathcal{A}$(4th elements) of all operands is non-empty. Then, by Step 1 of Definition 20, the result of the complete sequential addition for any index order is set to $\mathrm{oie}_{\mathrm{void}}$. Since $\mathrm{oie}_{\mathrm{void}}$ is the unique instance of type $\mathrm{OIE}_{\mathrm{void}}$, both operations yield the same result, namely $\mathrm{oie}_{\mathrm{void}}$. By Property 14, $\mathrm{oie}_{\mathrm{void}}$ is permutationally equivalent via any permutation matrix to itself. Therefore, in this case, the two results of complete sequential addition are permutationally equivalent.
This proof that the two results of complete sequential multiplication are permutationally equivalent in these two conditions.
(2) The situation $\mathrm{OIE}_{\mathrm{void}}$ instance doesn’t exist in $\mathrm{oie}\mathcal{S}$ and the intersection of all $\mathcal{A}$ across all operands is $\emptyset$.
In Step 2, According to Property 10
$\mathfrak{f}\mathrm{Feas2tuple}\mathcal{TS}(\mathrm{oie}\mathcal{S},\mathrm{idx}\mathcal{T}_{1})\stackrel{{\scriptstyle\mathcal{M}}}{{\cong}}\mathfrak{f}\mathrm{Feas2tuple}\mathcal{TS}(\mathrm{oie}\mathcal{S},\mathrm{idx}\mathcal{T}_{2})$
If
$\mathfrak{f}\mathrm{Feas2tuple}\mathcal{TS}(\mathrm{oie}\mathcal{S},\mathrm{idx}\mathcal{T}_{1})=\emptyset$
$\land$ $\mathfrak{f}\mathrm{Feas2tuple}\mathcal{TS}(\mathrm{oie}\mathcal{S},\mathrm{idx}\mathcal{T}_{2})=\emptyset$
then according to Definition 20
$\mathrm{oie}_{i}=\mathrm{oie}_{j}=\mathrm{oie}_{\mathrm{void}}$
according to Property 14
$\mathrm{oie}_{i}\stackrel{{\scriptstyle\mathcal{M}}}{{\sim}}\mathrm{oie}_{j}$
this completes the proof for this conditional branch.
Otherwise, they are both not $\emptyset$, process goes to step 3,
Get two $\mathrm{AscOrderFlt}\mathrm{2Tuple}\mathcal{TS}$ instances of this two $\mathrm{2Tuple}\mathcal{TS}^{feas}$ instances, then assign them to $\mathcal{F}_{\mathrm{oie}_{i}}$ and $\mathcal{F}_{\mathrm{oie}_{j}}$
$\mathcal{F}_{\mathrm{oie}_{i}}=\mathfrak{f}\mathrm{AscOrderFlt2tupleTS}(\mathfrak{f}\mathrm{Feas2tuple}\mathcal{TS}(\mathrm{oie}\mathcal{S},\mathrm{idx}\mathcal{T}_{1})),$
$\mathcal{F}_{\mathrm{oie}_{j}}=\mathfrak{f}\mathrm{AscOrderFlt2tupleTS}(\mathfrak{f}\mathrm{Feas2tuple}\mathcal{TS}(\mathrm{oie}\mathcal{S},\mathrm{idx}\mathcal{T}_{2}))$
According to Definition 21, the permutation isomorphism between $\mathfrak{f}\mathrm{Feas2tuple}\mathcal{TS}(\mathrm{oie}\mathcal{S},\mathrm{idx}\mathcal{T}_{1})$ and $\mathfrak{f}\mathrm{Feas2tuple}\mathcal{TS}(\mathrm{oie}\mathcal{S},\mathrm{idx}\mathcal{T}_{2})$ can not make $\mathcal{F}_{\mathrm{oie}_{i}}$ and $\mathcal{F}_{\mathrm{oie}_{j}}$ must be permutationally equivalent. Thus, in situation (2), Permutational equivalence need not hold.
Combining the conclusions of (1) and (2), we complete the proof.
4.7 Natural Complete sequential addition
Definition 24 (Natural complete sequential addition of members of a finite $\mathrm{OIE}\mathcal{S}$ instance in an index order).
There is a finite $\mathrm{OIE}\mathcal{S}$ instance $\mathrm{oie}\mathcal{S}$ containing $n(n\in\mathbb{N}^{+})$ elements in the ordered enumeration form
$\mathrm{oie}\mathcal{S}=\{\mathrm{oie}_{1},\mathrm{oie}_{2},\mathrm{oie}_{3},\cdots,\mathrm{oie}_{n}\}$
and an $\mathrm{Idx}\mathcal{T}$ instance with length $n$
$\mathrm{idx}\mathcal{T}=(\mathrm{idx}_{1},\mathrm{idx}_{2},\mathrm{idx}_{3},\cdots,\mathrm{idx}_{n})$
For any $\mathrm{OIE}\mathcal{S}$ instance $\mathrm{oie}_{\lambda}$, let
$\pi_{1}^{\min}(\mathrm{oie}_{\lambda})=\min\left\{\pi_{1}(t)\mid t\in\mathcal{I}_{\mathrm{oie}_{\lambda}}\right\}$
where $\pi_{1}(t)$ denotes the first projection(get the 1st element) of 2-tuple $t$, and
$\pi_{2}^{\max}(\mathrm{oie}_{\lambda})=\max\left\{\pi_{2}(t)\mid t\in\mathcal{I}_{\mathrm{oie}_{\lambda}}\right\}$
where $\pi_{2}(t)$ denotes the second projection(get the 2nd element) of 2-tuple $t$.
If
$\exists(A,B),\forall i\in[1,n]:$
$\qquad\mathrm{oie}_{i}\neq\mathrm{oie}_{\mathrm{void}}\land\pi_{1}^{\min}(\mathrm{oie}_{i})=A\land\pi_{2}^{\max}(\mathrm{oie}_{i})=B$
then by Definition 20, the following complete sequential addition
${\oplus|_{A}^{B}}(\mathrm{oie}_{\mathrm{idx}_{1}},\ \mathrm{oie}_{\mathrm{idx}_{2}},\ \cdots,\ \mathrm{oie}_{\mathrm{idx}_{n}})$
is called “Natural Complete Sequential Addition”, and the expression can be simplified to
$\oplus(\mathrm{oie}_{\mathrm{idx}_{1}},\ \mathrm{oie}_{\mathrm{idx}_{2}},\ \cdots,\ \mathrm{oie}_{\mathrm{idx}_{n}})$
Natural complete sequential addition serves as a natural special case of complete sequential addition free from subjective intervention, which eliminates the subjectivity arising from the artificial specification of the domain-filtering 2-tuple. It directly takes the common and natural time boundaries of all $\mathrm{OIE}$ instances participating in the operation, namely the unified minimum left boundary A and the unified maximum right boundary B, as the filtering domain of the operation, thus ensuring the objectivity and natural adaptability of the operation.
It fully inherits the core semantics of complete sequential addition and is still applied to model concurrent events with equal opportunity within a shared time domain. It only imposes a naturalized and standardized qualification on the filtering domain without altering the core rules of the operation. Meanwhile, it provides a concise form for the standardized modeling of real-world scenarios, fitting well with concurrent scenarios that naturally share a unified time window such as the 100 metres race. For such scenarios, event modeling can be completed directly through the simplified expression without additional specification of the filtering domain.
From a formal perspective, the removal of the filtering domain in Natural Complete Sequential Addition results in a notational form identical to that of Complete Sequential Multiplication. This unification endows the two sequential operations with a high degree of mathematical elegance.
5 Algebraic properties of sequential operations
In the existing research on the algebraic properties of event order, one approach is lattice theory(see Birkhoff (1940)). Lattice theory analyzes the order of events by means of several symmetric methods and structures related to group theory.
The method in this paper is not based on lattice theory but directly based on group theory. This is mainly because $\mathrm{OIE}$(a 4-tuple) and the two sequential operations we proposed introduce the idea of subjective arrangement. With an $\mathrm{OIE}$ instance, before the corresponding Event instance occurs, we already know or assume all the possible values of its interval. By using it, we manipulate the occurrence of events, and we are no longer passive observers but take the initiative, which is quite different from the traditional thinking in classical physics. We will further discuss this point later.
The content of this section is based on Section 3 and Section 4. We will conduct an in-depth exploration of the algebraic properties of the two sequential operations respectively.
5.1 Closure
In this subsection, we discuss the closure properties of the two sequential operations and some related derived content.
Property 17 (Complete Sequential Addition and Complete Sequential Multiplication both satisfy closure).
Proof of Property 17.
According to Definition 20 and Definition 22, complete sequential addition and complete sequential multiplication both return an $\mathrm{OIE}$ instance. It is either $\mathrm{oie}_{\mathrm{void}}$, with form
$($$)\left$(\ (),\ $\emptyset$,\ $\emptyset$,\ $\emptyset\ \right$)
or a non-$\mathrm{oie}_{\mathrm{void}}$ $\mathrm{OIE}$ instance with form
$\left(\mathcal{C},\mathcal{F},\mathcal{I},\mathcal{A}\right)$
in which
$the\ length\ of\ \mathcal{C}\ >0$
$\land$ $\left|{\mathcal{F}}\right|>0$
$\land$ $\left|{\mathcal{I}}\right|>0$
$\land$ $\left|{\mathcal{A}}\right|>0$
Since $\mathrm{oie}_{\mathrm{void}}$is also an instance of $\mathrm{OIE}$, it follows directly from these two definitions that they satisfy the closure property.
5.2 Commutative
In this subsection, we discuss the commutativity of the two sequential operations.
Property 18 (Complete sequential addition and complete sequential multiplication are non-commutative in general.).
Proof of the Property 18.
Suppose there are an $\mathrm{OIE}\mathcal{S}$ instance $\mathrm{oie}\mathcal{S}$ with cardinality $n>0$
$\mathrm{oie}\mathcal{S}=\{\ \mathrm{oie}_{1},\ \mathrm{oie}_{2},\ \mathrm{oie}_{3},\cdots,\mathrm{oie}_{n}\}$
two $\mathrm{Idx}\mathcal{T}$ instances with length $n$
$\mathrm{idx}\mathcal{T}_{i}=(\ i_{1},\ i_{2},\ i_{3},\ \cdots,\ i_{n}\ ),$
$\mathrm{idx}\mathcal{T}_{j}=(\ j_{1},\ j_{2},\ j_{3},\ \cdots,\ j_{n}\ )$
satisfying
$\mathrm{idx}\mathcal{T}_{i}\neq\mathrm{idx}\mathcal{T}_{j}$
$\land$ $\exists\ a\ permutation\ matrix\ \mathcal{M}:$
$\qquad\mathrm{idx}\mathcal{T}_{i}\stackrel{{\scriptstyle\mathcal{M}}}{{\sim}}\mathrm{idx}\mathcal{T}_{j}\$
and a $\mathrm{DomFlt2Tuple}$ instance $(\alpha,\beta)$
(1) Complete sequential addition
Let
$\mathrm{oie}_{i}={\oplus|_{\alpha}^{\beta}}(\mathrm{oie}_{i_{1}},\ \mathrm{oie}_{i_{2}},\ \cdots,\ \mathrm{oie}_{i_{n}}),$
$\mathrm{oie}_{j}={\oplus|_{\alpha}^{\beta}}(\mathrm{oie}_{j_{1}},\ \mathrm{oie}_{j_{2}},\ \cdots,\ \mathrm{oie}_{j_{n}})$
According to Definition 15, exists a permutation matrix $\mathcal{M}$
$\mathrm{oie}_{i}\stackrel{{\scriptstyle\mathcal{M}}}{{\sim}}\mathrm{oie}_{j}$
If $\mathrm{oie}_{i}=\mathrm{oie}_{\mathrm{void}}$, then $\mathrm{oie}_{j}=\mathrm{oie}_{\mathrm{void}}$ as well. So in this condition
$\mathrm{oie}_{i}=\mathrm{oie}_{j}$
Otherwise $\mathrm{oie}_{i}\neq\mathrm{oie}_{\mathrm{void}}$, then $\mathrm{oie}_{j}\neq\mathrm{oie}_{\mathrm{void}}$ as well. We have imposed the condition that $\mathrm{idx}\mathcal{T}_{i}\neq\mathrm{idx}\mathcal{T}_{j}$, from which it follows that $\mathcal{C}_{oie_{i}}\neq\mathcal{C}_{oie_{j}}$. According to Definition 8, we can conclude that
$\mathrm{oie}_{i}\neq\mathrm{oie}_{j}$
Thus, we can conclude that complete sequential addition is non-commutative in general.
(2) Complete sequential multiplication
Let
$\mathrm{oie}_{i}={\otimes}(\mathrm{oie}_{i_{1}},\ \mathrm{oie}_{i_{2}},\ \cdots,\ \mathrm{oie}_{i_{n}}),$
$\mathrm{oie}_{j}={\otimes}(\mathrm{oie}_{j_{1}},\ \mathrm{oie}_{j_{2}},\ \cdots,\ \mathrm{oie}_{j_{n}})$
According to Definition 16,
$\mathrm{idx}\mathcal{T}_{i}\stackrel{{\scriptstyle\mathcal{M}}}{{\sim}}\mathrm{idx}\mathcal{T}_{j}$
can’t guarantee that $\mathrm{oie}_{i}$ and $\mathrm{oie}_{j}$ are permutation equivalent, it follows that their equality cannot be guaranteed either. Thus, we can conclude that complete sequential multiplication is non-commutative in general.
Combining the conclusions of (1) and (2), we complete the proof.
5.3 Orbit space of complete sequential addition and complete sequential multiplication
In this subsection, we will discuss the orbit space of complete sequential addition and complete sequential multiplication.
5.3.1 Orbit space of complete sequential addition
Definition 25 (Orbit space of complete sequential addition).
Suppose there is an $\mathrm{OIE}\mathcal{S}$ instance with cardinality $0<n<\infty$, with form
$\mathrm{oie}\mathcal{S}=\{\ \mathrm{oie}_{1},\ \mathrm{oie}_{2},\ \mathrm{oie}_{3},\cdots,\mathrm{oie}_{n}\}$
an $\mathrm{idx}\mathcal{T}\mathcal{S}$ be the set of all $\mathrm{Idx}\mathcal{T}$ instances of length n with form
$\mathrm{idx}\mathcal{T}\mathcal{S}=\left\{(\ i_{1},\ i_{2},\ i_{3},\ \cdots,\ i_{n}\ )\Big|\{i_{1},\ i_{2},\ i_{3},\ \cdots,\ i_{n}\}=\{1,\ 2,\ 3,\ \cdots,\ n\}\right\}$
and a $\mathrm{DomFlt2Tuple}$ instance $(\alpha,\beta)$. $\oplus|_{\alpha}^{\beta}$} denote the Complete Sequential Addition operation (Definition 20) with domain-filtering 2-tuple $\alpha$ and $\beta$.
[…]
$\mathcal{O}\left(oieS,\ {\oplus|_{\alpha}^{\beta}}\right)=\left\{[\mathrm{oie}_{res}]_{\sim}\Big|\exists\mathrm{idx}\mathcal{T}\in\mathrm{idx}\mathcal{T}\mathcal{S}:\mathrm{oie}_{res}={\oplus|_{\alpha}^{\beta}}(\mathrm{oie}_{i_{1}},\ \mathrm{oie}_{i_{2}},\ \cdots,\ \mathrm{oie}_{i_{n}})\right\}$
where $\mathrm{oie}]_{\sim}$ denotes the equivalence class of $\mathrm{OIE}$ under permutational equivalence.
Property 19 (Orbit space of complete sequential addition has only one element).
Proof of Property 19.
Let $\mathrm{oie}\mathcal{S}$ be a finite $\mathrm{OIE}\mathcal{S}$instance with n elements $(0<n<\infty)$ and let $(\alpha,\beta)$ be a $\mathrm{DomFlt2Tuple}$ instance.
Denote by $\mathrm{idx}\mathcal{T}\mathcal{S}$ the set of all $\mathrm{Idx}\mathcal{T}$ instances of length n that are permutations of $\{1,2,3,\cdots,n\}$.
$\forall\mathrm{idx}\mathcal{T}_{1},\mathrm{idx}\mathcal{T}_{2}(\mathrm{idx}\mathcal{T}_{1}\neq\mathrm{idx}\mathcal{T}_{2})\ \in\mathrm{idx}\mathcal{T}\mathcal{S},\exists\ a\ permutation\ matrix\ \mathcal{M}:$
$\qquad\mathrm{idx}\mathcal{T}_{1}\stackrel{{\scriptstyle\mathcal{M}}}{{\sim}}\mathrm{idx}\mathcal{T}_{2}$
Denote $\mathrm{idx}\mathcal{T}_{1}$ and $\mathrm{idx}\mathcal{T}_{2}$ as
$\mathrm{idx}\mathcal{T}_{1}=(i_{1},i_{2},\cdots,i_{n}),$
$\mathrm{idx}\mathcal{T}_{2}=(j_{1},j_{2},\cdots,j_{n})$
And let
$\mathrm{oie}_{res_{1}}={\oplus|_{\alpha}^{\beta}}(\mathrm{oie}_{i_{1}},\ \mathrm{oie}_{i_{2}},\ \cdots,\ \mathrm{oie}_{i_{n}}),$
$\mathrm{oie}_{res_{2}}={\oplus|_{\alpha}^{\beta}}(\mathrm{oie}_{j_{1}},\ \mathrm{oie}_{j_{2}},\ \cdots,\ \mathrm{oie}_{j_{n}})$
Then, Property 15 guarantees that
$\mathrm{oie}{res_{1}}\stackrel{{\scriptstyle\mathcal{M}}}{{\sim}}\mathrm{oie}{res_{2}}$
[…]
This complete the proof. ∎
5.3.2 Orbit space of complete sequential multiplication
Definition 26 (Orbit space of complete sequential multiplication).
Suppose there is an $\mathrm{OIE}\mathcal{S}$ instance with cardinality $0<n<\infty$, with form
$\mathrm{oie}\mathcal{S}=\{\ \mathrm{oie}_{1},\ \mathrm{oie}_{2},\ \mathrm{oie}_{3},\cdots,\mathrm{oie}_{n}\}$
an $\mathrm{idx}\mathcal{T}\mathcal{S}$ be the set of all $\mathrm{Idx}\mathcal{T}$ instances of length n,
$\mathrm{idx}\mathcal{T}\mathcal{S}=\left\{(\ i_{1},\ i_{2},\ i_{3},\ \cdots,\ i_{n}\ )\Big|\{i_{1},\ i_{2},\ i_{3},\ \cdots,\ i_{n}\}=\{1,\ 2,\ 3,\ \cdots,\ n\}\right\}$
$\otimes$} denote the Complete Sequential Multiplication operation (Definition 22).
[…]
$\mathcal{O}\left(oieS,\ {\otimes}\right)=\left\{[\mathrm{oie}_{res}]_{\sim}\Big|\exists\mathrm{idx}\mathcal{T}\in\mathrm{idx}\mathcal{T}\mathcal{S}:\mathrm{oie}_{res}={\otimes}(\mathrm{oie}_{i_{1}},\ \mathrm{oie}_{i_{2}},\ \cdots,\ \mathrm{oie}_{i_{n}})\right\}$
where $\mathrm{oie}]_{\sim}$ denotes the equivalence class of $\mathrm{OIE}$ under permutational equivalence.
Property 20 (Orbit space of complete sequential multiplication has at least one element).
Proof of Property 20.
Let $\mathrm{oie}\mathcal{S}$ be a finite $\mathrm{OIE}\mathcal{S}$instance with n elements $(0<n<\infty)$.
Denote by $\mathrm{idx}\mathcal{T}\mathcal{S}$ the set of all $\mathrm{Idx}\mathcal{T}$ instances of length n that are permutations of $\{1,2,3,\cdots,n\}$.
$\forall\mathrm{idx}\mathcal{T}_{1},\mathrm{idx}\mathcal{T}_{2}(\mathrm{idx}\mathcal{T}_{1}\neq\mathrm{idx}\mathcal{T}_{2})\in\mathrm{idx}\mathcal{T}\mathcal{S},\exists\ a\ permutation\ matrix\ \mathcal{M}:$
$\qquad\mathrm{idx}\mathcal{T}_{1}\stackrel{{\scriptstyle\mathcal{M}}}{{\sim}}\mathrm{idx}\mathcal{T}_{2}$
Denote $\mathrm{idx}\mathcal{T}_{1}$ and $\mathrm{idx}\mathcal{T}_{2}$ as
$\mathrm{idx}\mathcal{T}_{1}=(i_{1},i_{2},\cdots,i_{n}),$
$\mathrm{idx}\mathcal{T}_{2}=(j_{1},j_{2},\cdots,j_{n})$
And let
$\mathrm{oie}_{res_{1}}={\otimes}(\mathrm{oie}_{i_{1}},\ \mathrm{oie}_{i_{2}},\ \cdots,\ \mathrm{oie}_{i_{n}}),$
$\mathrm{oie}_{res_{2}}={\otimes}(\mathrm{oie}_{j_{1}},\ \mathrm{oie}_{j_{2}},\ \cdots,\ \mathrm{oie}_{j_{n}})$
Then, Property 16 guarantees that
$\mathrm{oie}{res_{1}}\ and\ \mathrm{oie}{res_{2}}\ need\ not\ be\ permutationally\ equivalent.$
[…]
This complete the proof. ∎
6 The implement of optional intervals event
In traditional classical physics, we observe the world, draw conclusions, and conduct experiments to verify them. From the perspective of the ideas presented in this paper, this approach is rather passive, everything ultimately stems from observation, and we are totally led by the “observed phenomena”.
For instance, phenomenon A occurs first, followed by phenomenon B, and this sequence repeats 1000 times. Once it is observed by humans, various methods of traditional physics can be employed to analyze and verify it.
However, what if, on the 1001st occasion, phenomenon B occurs before phenomenon A, but by then, the conclusion that A comes before B has already been established? What should we do in such a situation? This is precisely the drawback of giving priority to observation over conclusion, and it is also a topic that many philosophers and physicists have discussed.
Now, with $\mathrm{OIE}$ and sequential operations, we can introduce subjectivity. In this chapter, we will discuss how to use $\mathrm{OIE}$ and sequential operations to determine the execution order of events, as well as what limitations exist.
6.1 The implementation of $\mathrm{OIE}$ instance
As defined in Definition 6, an $\mathrm{OIE}$ instance maps to an Event instance. An Event instance is concrete and executable, which is what we commonly regard as an event. In contrast, an $\mathrm{OIE}$ instance is an abstract object that stores information about several available intervals of the Event instance it maps to(except $\mathrm{oie}_{\mathrm{void}}$).
Therefore, one use of an $\mathrm{OIE}$ instance is to guide when an Event instance should be executed(in status running). In this subsection, we will discuss the guiding methods, and we will propose two types of them.
First type, we use the $\mathcal{I}$ of an $\mathrm{OIE}$ instance to determine the interval of the Event instance it maps to.
Definition 27 (The 1st type of implementation of an $\mathrm{OIE}$ instance).
Given a non-$\mathrm{oie}_{\mathrm{void}}$ $\mathrm{OIE}$ instance $\mathrm{oie}$ and an Event instance event, and
$\mathrm{oie}\rightarrow\textsf{event}$
From $\mathcal{I}_{\mathrm{oie}}$(3rd element of $\mathrm{oie}$), select an interval 2-tuple to serve as the running status interval of event. This is called “The first type of implementation of an $\mathrm{OIE}$ instance”.
The characteristic of the first type of implementation is that it only provides the running status interval of event from the starting timestamp to the ending timestamp, without giving the specific details within this interval.
In order to further determine the specific details within the running status interval of event, we propose another type of implementation for the $\mathrm{OIE}$ instance.
Definition 28 (The 2nd type of implementation of an $\mathrm{OIE}$ instance).
Given a non-empty $\mathrm{OIE}\mathcal{S}$ instance $\mathrm{oie}\mathcal{S}$ in the form of
$\mathrm{oie}\mathcal{S}=\{\ \mathrm{oie}_{1},\ \mathrm{oie}_{2},\ \mathrm{oie}_{3},\ \cdots,\ \mathrm{oie}_{n}\ \}$
and an $\mathrm{Idx}\mathcal{T}$ instance
$\mathrm{idx}\mathcal{T}=\{i_{1},i_{2},i_{3},\cdots,i_{n}\}$
[…]
$\mathrm{oie}_{res}=\bm{*}(\mathrm{oie}_{i_{1}},\mathrm{oie}_{i_{2}},\mathrm{oie}_{i_{3}},\cdots,\mathrm{oie}_{i_{n}})$
Also, there is an Event instance event satisfied
$\mathrm{oie}_{res}\rightarrow\textsf{event}$
From $\mathcal{F}_{\mathrm{oie}_{res}}$(2nd element of $\mathrm{oie}_{res}$), select a $\mathrm{2Tuple}\mathcal{T}$ instance $\mathrm{2tuple}\mathcal{T}$, with form
$\mathrm{2tuple}\mathcal{T}=(\ (x_{i_{1}},y_{i_{1}}),\ (x_{i_{2}},y_{i_{2}}),\ \cdots,\ (x_{i_{n}},y_{i_{n}})\ )$
For each sub Event instance $\textsf{event}_{i_{k}}$ that $\mathrm{oie}_{i_{k}}$ maps to, use $x_{i_{k}},y_{i_{k}})$ to serve as the running status interval of $\textsf{event}_{i_{k}}$. That means, set the starting timestamp of $\textsf{event}_{i_{k}}$ as the first item of $x_{i_{k}}$, and set the ending timestamp as the second item of $y_{i_{k}}$.
This is called the 2nd type of implementation of an $\mathrm{OIE}$ instance.
The 2nd type of implementation not only plans the overall starting timestamp and ending timestamp of the Event instance $\mathrm{oie}_{res}$ maps to, but also plans the running status of all the constituent events of this event, which makes the planning more specific.
For example, for a composed event, the overall time is from 1:00 PM to 5:00 PM on a certain day. It is composed of three sub-events, and their respective intervals are
[…]
It can be observed that the second type of implementation (Definition 28) can express that there is no sub-event execution in the interval $3:00\text{ PM},4:00\text{ PM})$. If the first type of implementation (Definition 27) is used, this information cannot be expressed.
We believe that the first type of implementation is suitable for the $\textsf{Atom}\mathrm{OIE}$ instance, which maps to an atomic and indivisible Event instance. The second type of implementation is more suitable for the $\textsf{Comp}\mathrm{OIE}$ instance, which maps to a composed event.
6.2 The relationship between 2 types of sequential operations and general n-ary finitary operations
Definition 29 (Projection based on ending timestamp of the $\mathcal{F}$ of an $\mathrm{OIE}$ instances obtained by a certain sequential operation for the set of n-ray finitary operations).
Let the following objects be given:
  1. 1.
    A set of operands $\mathrm{operand}\mathcal{S}=\{opr_{1},opr_{2},\dots,opr_{n}\}$ with $n\in\mathbb{N}^{+}$.
  2. 2.
    A set of events $\mathrm{event}\mathcal{S}=\{\textsf{event}_{1},\textsf{event}_{2},\dots,\textsf{event}_{n}\}$ together with a mapping $\mathrm{event}\mathcal{S}\to\mathrm{operand}\mathcal{S}$.
  3. 3.
    A set of $\mathrm{OIE}$ instances $\mathrm{oie}\mathcal{S}=\{\mathrm{oie}_{1},\mathrm{oie}_{2},\dots,\mathrm{oie}_{n}\}$ together with a mapping $\mathrm{oie}\mathcal{S}\to\mathrm{event}\mathcal{S}$.
  4. 4.
    An index tuple $\mathrm{idx}\mathcal{T}=(i_{1},i_{2},\dots,i_{n})$ where $\{i_{1},\dots,i_{n}\}=\{1,\dots,n\}$.
  5. 5.
    A sequential operation $\in\{\oplus|_{\alpha(n)}^{\beta},\ \otimes\}$. Applying * to $\mathrm{oie}\mathcal{S}$ and $\mathrm{idx}\mathcal{T}$ yields
    $\mathrm{oie}_{res}=*(\mathrm{oie}_{i_{1}},\mathrm{oie}_{i_{2}},\dots,\mathrm{oie}_{i_{n}}),$
    and we require $\mathrm{oie}_{res}\neq\mathrm{oie}_{\mathrm{void}}$, if it is not, then the projection does not hold.
  6. 6.
    An $n$-ary finitary operation symbol $\odot$ used to build expressions.
Construction of the projection
  1. 1.
    Initialize $\mathrm{operation}\mathcal{S}=\emptyset$.
  2. 2.
    For each $\mathrm{2tuple}\mathcal{T}\in\mathcal{F}_{\mathrm{oie}_{res}}$ having the form
    $\mathrm{2tuple}\mathcal{T}=((x_{1},y_{1}),(x_{2},y_{2}),\dots,(x_{n},y_{n})),$
    where $y_{k}$ is an ending timestamp of $\textsf{event}_{i_{k}}$ ($k=1,\dots,n$), proceed as follows:
    • Group equal ending timestamps. Let the distinct values among $\{y_{1},\dots,y_{n}\}$ be $v_{1}<v_{2}<\dots<v_{t}$. For each $v_{s}$ ($s=1,\dots,t$), let $I_{s}=\{k\mid y_{k}=v_{s}\}$ be the set of indices with that value, and let $m_{s}=|I_{s}|$. Denote by $O_{s}=\{opr_{k}\mid k\in I_{s}\}$ the corresponding multiset of operands.
    • Generate all intra‑group permutations. For each group $s$ let $\Pi_{s}$ be the set of all $m_{s}!$ permutations of the elements of $O_{s}$.
    • Form all global permutations respecting the group order. The set of admissible global permutations is the Cartesian product
      $\Pi=\Pi_{1}\times\Pi_{2}\times\dots\times\Pi_{t},$
      where each tuple $(\pi_{1},\dots,\pi_{t})\in\Pi$ corresponds to the concatenated permutation
      $\pi=\pi_{1}\|\pi_{2}\|\dots\|\pi_{t}$
      (the operands of group $O_{1}$ appear first in the order given by $\pi_{1}$, then those of $O_{2}$ in the order given by $\pi_{2}$, etc.).
    • Generate expressions for every global permutation. For each $\pi\in\Pi$ build the $n$-ary expression
      $\odot\!\bigl(\pi(1),\pi(2),\dots,\pi(n)\bigr),$
      and insert it into $\mathrm{operation}\mathcal{S}$.
  3. 3.
    Remove duplicate expressions from $\mathrm{operation}\mathcal{S}$.
The construction described above is called the projection based on the ending timestamps of the $\mathcal{F}$ element of $\mathrm{oie}_{res}$ obtained by the sequential operation *. It is denoted by
[…]
Property 21 (The number of operation orders determined by projection(Definition 29) with complete sequential addition for the set of n-ray finitary operations).
Let the following objects be given:
  1. 1.
    A set of operands $\mathrm{operand}\mathcal{S}=\{opr_{1},opr_{2},\dots,opr_{n}\}$ with $n\in\mathbb{N}^{+}$.
  2. 2.
    A set of events $\mathrm{event}\mathcal{S}=\{\textsf{event}_{1},\textsf{event}_{2},\dots,\textsf{event}_{n}\}$ together with a mapping $\mathrm{event}\mathcal{S}\to\mathrm{operand}\mathcal{S}$.
  3. 3.
    A set of $\mathrm{OIE}$ instances $\mathrm{oie}\mathcal{S}=\{\mathrm{oie}_{1},\mathrm{oie}_{2},\dots,\mathrm{oie}_{n}\}$ together with a mapping $\mathrm{oie}\mathcal{S}\to\mathrm{event}\mathcal{S}$.
  4. 4.
    An index tuple $\mathrm{idx}\mathcal{T}=(i_{1},i_{2},\dots,i_{n})$ where $\{i_{1},\dots,i_{n}\}=\{1,\dots,n\}$.
  5. 5.
    Complete sequential operation $\oplus|_{\alpha}^{\beta}$}. Applying $\oplus|_{\alpha}^{\beta}$} to $\mathrm{oie}\mathcal{S}$ and $\mathrm{idx}\mathcal{T}$ yields
    $\mathrm{oie}_{res}={\oplus|_{\alpha}^{\beta}}(\mathrm{oie}_{i_{1}},\mathrm{oie}_{i_{2}},\dots,\mathrm{oie}_{i_{n}}),$
    and we require $\mathrm{oie}_{res}\neq\mathrm{oie}_{\mathrm{void}}$, if it is not, then the projection does not hold.
  6. 6.
    An $n$-ary finitary operation symbol $\odot$ used to build expressions.
Using Definition 29, the following projection holds:
${\oplus|_{\alpha}^{\beta}}(\mathrm{oie}\mathcal{S},\ \mathrm{idx}\mathcal{T})\;\xrightarrow[\mathrm{event}\mathcal{S},\ \mathrm{operand}\mathcal{S},\ \odot]{\text{ascending order of TSend}}\;\mathrm{operation}\mathcal{S}.$
Then the range of the number of operand orders of the n-ary finitary operation is
$[n,n!]$
Proof of Property 21.
We prove the two bounds separately.
Lower bound ($\geq n$):
  • By Definition 19(domain-filtered subset), for each index $k\in\{1,\dots,n\}$, there exists a $2\mathrm{tuple}\mathcal{T}^{(k)}\in\mathcal{F}_{\text{oie}_{res}}$ such that the second component (the ending timestamp) of the $k$-th element in $2\mathrm{tuple}\mathcal{T}^{(k)}$ equals $\beta$.
  • In the projection construction (Definition 29), for each such $2\mathrm{tuple}\mathcal{T}^{(k)}$, the distinct ending timestamps are grouped. Since the $k$-th element has ending timestamp $\beta$, and $\beta$ is the maximum possible ending timestamp in any $2\mathrm{tuple}\mathcal{T}$ (by the domain-filtering condition), the group containing $\beta$ is the last group.
  • Within that last group, the operand $opr_{k}$ (corresponding to the $k$-th element) appears. By permuting the operands within the last group, we can obtain an operation order where $opr_{k}$ is placed last (since the last group is placed at the end of the operand sequence).
  • Therefore, for each $k$, there exists at least one operation order in $\text{operation}\mathcal{S}$ with $opr_{k}$ as the last operand. These $n$ operation orders are distinct because their last operands are different.
  • Hence, $|\text{operation}\mathcal{S}|\geq n$.
Upper bound ($\leq n!$):
  • Each element of $\text{operation}\mathcal{S}$ is an $n$-ary expression $\odot(\pi(1),\dots,\pi(n))$, where $\pi$ is a permutation of the $n$ operands.
  • There are at most $n!$ distinct permutations of $n$ distinct operands.
  • Therefore, $|\text{operation}\mathcal{S}|\leq n!$.
Combining both bounds, we have:
$n\leq|\text{operation}\mathcal{S}|\leq n!.$
This completes the proof. ∎
Property 22 (The number of operation orders determined by projection(Definition 29) with complete sequential multiplication for the set of n-ray finitary operations ).
Let the following objects be given:
  1. 1.
    A set of operands $\mathrm{operand}\mathcal{S}=\{opr_{1},opr_{2},\dots,opr_{n}\}$ with $n\in\mathbb{N}^{+}$.
  2. 2.
    A set of events $\mathrm{event}\mathcal{S}=\{\textsf{event}_{1},\textsf{event}_{2},\dots,\textsf{event}_{n}\}$ together with a mapping $\mathrm{event}\mathcal{S}\to\mathrm{operand}\mathcal{S}$.
  3. 3.
    A set of $\mathrm{OIE}$ instances $\mathrm{oie}\mathcal{S}=\{\mathrm{oie}_{1},\mathrm{oie}_{2},\dots,\mathrm{oie}_{n}\}$ together with a mapping $\mathrm{oie}\mathcal{S}\to\mathrm{event}\mathcal{S}$.
  4. 4.
    An index tuple $\mathrm{idx}\mathcal{T}=(i_{1},i_{2},\dots,i_{n})$ where $\{i_{1},\dots,i_{n}\}=\{1,\dots,n\}$.
  5. 5.
    Complete sequential multiplication $\otimes$. Applying $\otimes$ to $\mathrm{oie}\mathcal{S}$ and $\mathrm{idx}\mathcal{T}$ yields
    $\mathrm{oie}_{res}=\otimes(\mathrm{oie}_{i_{1}},\mathrm{oie}_{i_{2}},\dots,\mathrm{oie}_{i_{n}}),$
    and we require $\mathrm{oie}_{res}\neq\mathrm{oie}_{\mathrm{void}}$, if it is not, then the projection does not hold.
  6. 6.
    An $n$-ary finitary operation symbol $\odot$ used to build expressions.
Using Definition 29, the following projection holds:
$\otimes(\mathrm{oie}\mathcal{S},\ \mathrm{idx}\mathcal{T})\;\xrightarrow[\mathrm{event}\mathcal{S},\ \mathrm{operand}\mathcal{S},\ \odot]{\text{ascending order of TSend}}\;\mathrm{operation}\mathcal{S}.$
Then the number of operand orders of the n-ary finitary operation is 1
Proof of Property 22.
according to Definition 22 and Definition 21, condition
$\begin{array}[c]{@{}l@{}}\forall\ \mathrm{2tuple}_{i},\ \mathrm{2tuple}_{j}\in\mathrm{2tuple}\mathcal{T}\ \land\ i<j:\\ \ \ \ the\ 2nd\ item\ of\ \mathrm{2tuple}_{i}\leq the\ 1st\ item\ of\ \mathrm{2tuple}_{j}\end{array}$
Therefore, their 2nd item of interval 2tuple must be arranged in increasing order, which determines that the n Event instances mapped by $\mathrm{oie}_{res}$ must be arranged in the order of the $\mathrm{OIE}$ instances participating in the complete sequential multiplication, so there is only one permutation, the number is 1. ∎
Corollary 1 (Uniqueness of sets of n-ray finitary operations within an equivalence class in orbit space).
Given an operand set $\mathrm{operand}\mathcal{S}$ with cardinality $0<n<\infty$, an $\mathrm{Event}\mathcal{S}$ instance $\mathrm{event}\mathcal{S}$ and an $\mathrm{OIE}\mathcal{S}$ instance $\mathrm{oie}\mathcal{S}$. Exists $\mathrm{Idx}\mathcal{T}$ instances, for any $\mathrm{Idx}\mathcal{T}$ instance $\mathrm{idx}\mathcal{T}$ in them, there is an operation set $\mathrm{operation}\mathcal{S}$ with cardinality $0<m<\infty$ for n-ary finitary operation $\odot$, they satisfy
[…]
where * denotes either the complete sequential addition $\oplus|_{\alpha}^{\beta}$ or the complete sequential multiplication $\otimes$.
Let $\mathcal{O}(\mathrm{oie}\mathcal{S},*)$ be the orbit space of the sequential operation * on $\mathrm{oie}\mathcal{S}$ (see Definition 25 and Definition 26). Then, for each equivalence class $\mathrm{oie}_{res}]_{\sim}$ in $\mathcal{O}(\mathrm{oie}\mathcal{S},*)$, all $\mathrm{OIE}$ instance in $\mathrm{oie}_{res}]_{\sim}$ yield the same set of n-ary finitary operation expression via the projection.
That is, for any $\mathrm{oie}_{res_{1}},\mathrm{oie}_{res_{2}}\in[\mathrm{oie}_{res}]_{\sim}$, the $\mathrm{Operation}\mathcal{S}$ instances are the same.
Proof of Cor. 1.
Assume $\mathrm{oie}_{res_{1}}$ and $\mathrm{oie}_{res_{2}}$ belong to the same equivalence class in $\mathcal{O}(\mathrm{oie}\mathcal{S},\ast)$. By the definition of the orbit space, $\mathrm{oie}_{res_{1}}$ and $\mathrm{oie}_{res_{2}}$ are permutationally equivalent (Definition 3.15). Hence there exists a permutation matrix $\mathcal{M}$ such that
$\mathcal{C}_{\mathrm{oie}_{res_{1}}}\cdot\mathcal{M}=\mathcal{C}_{\mathrm{oie}_{res_{2}}},\qquad\mathcal{F}_{\mathrm{oie}_{res_{1}}}\stackrel{{\scriptstyle\mathcal{M}}}{{\cong}}\mathcal{F}_{\mathrm{oie}_{res_{2}}},\qquad\mathcal{I}_{\mathrm{oie}_{res_{1}}}=\mathcal{I}_{\mathrm{oie}_{res_{2}}},\qquad\mathcal{A}_{\mathrm{oie}_{res_{1}}}=\mathcal{A}_{\mathrm{oie}_{res_{2}}}.$
The permutation isomorphism $\mathcal{F}_{\mathrm{oie}_{res_{1}}}\stackrel{{\scriptstyle\mathcal{M}}}{{\cong}}\mathcal{F}_{\mathrm{oie}_{res_{2}}}$ means that there is a bijection $\phi:\mathcal{F}_{\mathrm{oie}_{res_{1}}}\to\mathcal{F}_{\mathrm{oie}_{res_{2}}}$ with $\phi(2\mathrm{tuple}\mathcal{T})=2\mathrm{tuple}\mathcal{T}\cdot\mathcal{M}$ for every $2\mathrm{tuple}\mathcal{T}\in\mathcal{F}_{\mathrm{oie}_{res_{1}}}$.
Now consider the projection defined in Definition 5.3. For each $2\mathrm{tuple}\mathcal{T}\in\mathcal{F}_{\mathrm{oie}_{res_{1}}}$, the projection constructs a set of $n$-ary expressions $\odot(\pi(1),\dots,\pi(n))$ based on the ending timestamps (the second components) of the $2$-tuples in $2\mathrm{tuple}\mathcal{T}$. Let us denote this set by $\mathsf{OpSet}(2\mathrm{tuple}\mathcal{T})$. The final operation set $\mathsf{operation}\mathcal{S}_{1}$ for $\mathrm{oie}_{res_{1}}$ is the union of all $\mathsf{OpSet}(2\mathrm{tuple}\mathcal{T})$ for $2\mathrm{tuple}\mathcal{T}\in\mathcal{F}_{\mathrm{oie}_{res_{1}}}$ (with duplicates removed). Similarly, for $\mathrm{oie}_{res_{2}}$ we obtain $\mathsf{operation}\mathcal{S}_{2}$ from $\mathcal{F}_{\mathrm{oie}_{res_{2}}}$.
Take an arbitrary $2\mathrm{tuple}\mathcal{T}\in\mathcal{F}_{\mathrm{oie}_{res_{1}}}$ and write it as
$2\mathrm{tuple}\mathcal{T}=\big((x_{1},y_{1}),(x_{2},y_{2}),\dots,(x_{n},y_{n})\big).$
Its image under $\phi$ is
$2\mathrm{tuple}\mathcal{T}^{\prime}=\phi(2\mathrm{tuple}\mathcal{T})=2\mathrm{tuple}\mathcal{T}\cdot\mathcal{M}=\big((x_{M(1)},y_{M(1)}),(x_{M(2)},y_{M(2)}),\dots,(x_{M(n)},y_{M(n)})\big),$
where $M$ is the permutation corresponding to $\mathcal{M}$. The ending timestamps in $2\mathrm{tuple}\mathcal{T}$ are the multiset $\{y_{1},\dots,y_{n}\}$, which is the same as in $2\mathrm{tuple}\mathcal{T}^{\prime}$ because $\mathcal{M}$ merely reorders the pairs. Consequently, the grouping of indices by equal ending timestamps is preserved up to the permutation: if for a value $v$ the set of indices with $y_{k}=v$ in $2\mathrm{tuple}\mathcal{T}$ is $I$, then in $2\mathrm{tuple}\mathcal{T}^{\prime}$ the set of indices with the same ending timestamp $v$ is $M(I)=\{M(k)\mid k\in I\}$.
Recall that the operands are associated with the OIE instances via the fixed mappings $\mathrm{oie}\mathcal{S}\to\mathrm{event}\mathcal{S}\to\mathsf{operand}\mathcal{S}$. For $\mathrm{oie}_{res_{1}}$, the tuple $\mathcal{C}_{\mathrm{oie}_{res_{1}}}=(\mathrm{oie}_{i_{1}},\dots,\mathrm{oie}_{i_{n}})$ determines that the $k$-th component of any $2\mathrm{tuple}\mathcal{T}\in\mathcal{F}_{\mathrm{oie}_{res_{1}}}$ corresponds to operand $opr_{i_{k}}$. For $\mathrm{oie}_{res_{2}}$, we have $\mathcal{C}_{\mathrm{oie}_{res_{2}}}=(\mathrm{oie}_{j_{1}},\dots,\mathrm{oie}_{j_{n}})$ with $(j_{1},\dots,j_{n})=(i_{M(1)},\dots,i_{M(n)})$ because $\mathcal{C}_{\mathrm{oie}_{res_{1}}}\cdot\mathcal{M}=\mathcal{C}_{\mathrm{oie}_{res_{2}}}$. Hence the operand corresponding to the $k$-th component of $2\mathrm{tuple}\mathcal{T}^{\prime}$ is $opr_{j_{k}}=opr_{i_{M(k)}}$.
Now examine the construction of $\mathsf{OpSet}(2\mathrm{tuple}\mathcal{T})$ and $\mathsf{OpSet}(2\mathrm{tuple}\mathcal{T}^{\prime})$. In both cases, the operands are partitioned into groups according to the distinct ending timestamps $v_{1}<v_{2}<\dots<v_{t}$. For $2\mathrm{tuple}\mathcal{T}$, the group for $v_{s}$ contains the operands $\{opr_{i_{k}}\mid k\in I_{s}\}$. For $2\mathrm{tuple}\mathcal{T}^{\prime}$, the group for $v_{s}$ contains the operands $\{opr_{j_{k}}\mid k\in M(I_{s})\}=\{opr_{i_{M(k)}}\mid k\in M(I_{s})\}$. Since $M$ is a bijection, $\{opr_{i_{M(k)}}\mid k\in M(I_{s})\}=\{opr_{i_{k}}\mid k\in I_{s}\}$. Thus each group consists of exactly the same set of operands.
The projection then generates all permutations of the operands that respect the group order: the operands of the first group (with the smallest ending timestamp) appear first in any order, followed by the operands of the second group, etc. Because the groups are identical as sets, the collection of all such permutations is the same for $2\mathrm{tuple}\mathcal{T}$ and $2\mathrm{tuple}\mathcal{T}^{\prime}$. Therefore $\mathsf{OpSet}(2\mathrm{tuple}\mathcal{T})=\mathsf{OpSet}(2\mathrm{tuple}\mathcal{T}^{\prime})$.
Since $\phi$ is a bijection between $\mathcal{F}_{\mathrm{oie}_{res_{1}}}$ and $\mathcal{F}_{\mathrm{oie}_{res_{2}}}$, we have
$\bigcup_{2\mathrm{tuple}\mathcal{T}\in\mathcal{F}_{\mathrm{oie}_{res_{1}}}}\mathsf{OpSet}(2\mathrm{tuple}\mathcal{T})=\bigcup_{2\mathrm{tuple}\mathcal{T}^{\prime}\in\mathcal{F}_{\mathrm{oie}_{res_{2}}}}\mathsf{OpSet}(2\mathrm{tuple}\mathcal{T}^{\prime}).$
Removing duplicate expressions (as required in Definition 5.3) yields the same set of operation expressions. Hence $\mathsf{operation}\mathcal{S}_{1}=\mathsf{operation}\mathcal{S}_{2}$.
This shows that any two permutationally equivalent OIE instances produce the same set of $n$-ary finitary operation expressions through the projection. Therefore, within an equivalence class of the orbit space, all OIE instances yield the same $\mathsf{operation}\mathcal{S}$. ∎
6.3 The restriction rules between the implementation of $\mathrm{OIE}$ and the binary algebraic system
In this subsection, we will present more general conclusions that can be applied in practice for complete sequential addition and complete sequential multiplication.
Corollary 2 (The Relationship between the projection based on ending timestamp of the $\mathcal{F}$ of an $\mathrm{OIE}$ instances obtained by complete sequential addition $\oplus|_{\alpha}^{\beta}$ and the n-ary finitary operation set ).
Given an $\mathrm{OIE}\mathcal{S}$ instance $\mathrm{oie}\mathcal{S}$ with cardinality $0<n<+\infty$
$\mathrm{oie}\mathcal{S}=\{\mathrm{oie}_{1},\mathrm{oie}_{2},\cdots,\mathrm{oie}_{n}\}$
an $\mathrm{Event}\mathcal{S}$ instance $\mathrm{event}\mathcal{S}$ with cardinality $0<n<+\infty$, an $\mathrm{Operand}\mathcal{S}$ instance $\mathrm{operand}\mathcal{S}$ satisfying
$\mathrm{oie}\mathcal{S}\rightarrow\mathrm{event}\mathcal{S}\rightarrow\mathrm{operation}\mathcal{S}$
with form
$\mathrm{operand}\mathcal{S}=\{F_{1}(\mathrm{oie}_{1}),F_{2}(\mathrm{oie}_{2}),\cdots,F_{n}(\mathrm{oie}_{n})\}$
and an $\mathrm{Operation}\mathcal{S}$ instance $\mathrm{operation}\mathcal{S}$ with n-ray finitary operation $\odot$. And They form a projection based on ending timestamp of the $\mathcal{F}$ of an $\mathrm{OIE}$ instances obtained by complete sequential addition $\oplus|_{\alpha}^{\beta}$ under the $\mathrm{Idx}\mathcal{T}$ instance $\mathrm{idx}\mathcal{T}$ (Definition 29)
$\oplus|_{\alpha}^{\beta}(\mathrm{oie}\mathcal{S},\ \mathrm{idx}\mathcal{T})\;\xrightarrow[\mathrm{event}\mathcal{S},\ \mathrm{operand}\mathcal{S},\ \odot]{\text{ascending order of TSend}}\;\mathrm{operation}\mathcal{S}.$
If $\odot$” is an n-ary finitary operation that satisfies permutation equivalence
$\forall\ (opr_{i_{1}},opr_{i_{2}},\cdots,opr_{i_{n}})\stackrel{{\scriptstyle\mathcal{M}}}{{\sim}}(opr_{j_{1}},opr_{j_{2}},\cdots,opr_{j_{n}}):$
$\quad\odot(opr_{i_{1}},opr_{i_{2}},\cdots,opr_{i_{n}})=\odot(opr_{j_{1}},opr_{j_{2}},\cdots,opr_{j_{n}})$
then
$\forall\ \odot(opr_{i_{1}},opr_{i_{2}},\cdots,opr_{i_{n}}),\odot(opr_{j_{1}},opr_{j_{2}},\cdots,opr_{j_{n}})\in\mathrm{operation}\mathcal{S}:$
$\odot(opr_{i_{1}},opr_{i_{2}},\cdots,opr_{i_{n}})=\odot(opr_{j_{1}},opr_{j_{2}},\cdots,opr_{j_{n}})$
Otherwise, if $\odot$” is an n-ary finitary operation that does not satisfy permutation equivalence, then the above relationship is not necessarily satisfied.
Corollary 3 (The Relationship between the projection based on ending timestamp of the $\mathcal{F}$ of an $\mathrm{OIE}$ instances obtained by a complete sequential multiplication and the n-ary finitary operation set ).
Given an $\mathrm{OIE}\mathcal{S}$ instance $\mathrm{oie}\mathcal{S}$ with cardinality $0<n<+\infty$, an $\mathrm{Event}\mathcal{S}$ instance $\mathrm{event}\mathcal{S}$ with cardinality $0<n<+\infty$, an $\mathrm{Operand}\mathcal{S}$ instance $\mathrm{operand}\mathcal{S}$ satisfying
$\mathrm{oie}\mathcal{S}\rightarrow\mathrm{event}\mathcal{S}\rightarrow\mathrm{operation}\mathcal{S}$
with form
$\mathrm{operand}\mathcal{S}=\{F_{1}(\mathrm{oie}_{1}),F_{2}(\mathrm{oie}_{2}),\cdots,F_{n}(\mathrm{oie}_{n})\}$
and an $\mathrm{Operation}\mathcal{S}$ instance $\mathrm{operation}\mathcal{S}$ with n-ray finitary operation $\odot$. And They form a projection based on ending timestamp of the $\mathcal{F}$ of an $\mathrm{OIE}$ instances obtained by complete sequential multiplication $\otimes$ under the $\mathrm{Idx}\mathcal{T}$ instance $\mathrm{idx}\mathcal{T}$ (Definition 29)
$\otimes(\mathrm{oie}\mathcal{S},\ \mathrm{idx}\mathcal{T})\;\xrightarrow[\mathrm{event}\mathcal{S},\ \mathrm{operand}\mathcal{S},\ \odot]{\text{ascending order of TSend}}\;\mathrm{operation}\mathcal{S}$
where
$\mathrm{idx}\mathcal{T}=(i_{1},i_{2},\cdots,i_{n})$
Then, the $\mathrm{operation}\mathcal{S}$ has only one element operation
$\odot(F_{i_{1}}(\mathrm{oie}_{i_{1}}),F_{i_{2}}(\mathrm{oie}_{i_{2}}),\cdots,F_{i_{n}}(\mathrm{oie}_{i_{n}}))$
That means no matter whether $\odot$” is an n-ary finitary operation that satisfies permutation equivalence
$\forall\ \odot(opr_{i_{1}},opr_{i_{2}},\cdots,opr_{i_{n}}),\odot(opr_{j_{1}},opr_{j_{2}},\cdots,opr_{j_{n}})\in\mathrm{operation}\mathcal{S}:$
$\odot(opr_{i_{1}},opr_{i_{2}},\cdots,opr_{i_{n}})=\odot(opr_{j_{1}},opr_{j_{2}},\cdots,opr_{j_{n}})$
Corollary 4 (Uniqueness of the operation set for each equivalence class in the orbit space).
Let $\mathrm{oie}\mathcal{S}$ be a finite $\mathrm{OIE}$ instance and let $\in\{\oplus|_{\alpha(n)}^{\beta},\otimes\}$ be a sequential operation. Denote by $\mathcal{O}(\mathrm{oie}\mathcal{S},*)$ the corresponding orbit space (see Definitions 25 and Definition 26). For every equivalence class $\mathrm{oie}_{res}]_{\sim}\in\mathcal{O}(\mathrm{oie}\mathcal{S},*)$, there exists a unique set of n-ary finitary operation expressions, denoted $\mathrm{operation}\mathcal{S}_{[\mathrm{oie}_{res}]_{\sim}}$, such that for each $\mathrm{oie}_{res}\in[\mathrm{oie}_{res}]_{\sim}$ the projection defined in Definition 29 yields exactly $\mathrm{operation}\mathcal{S}_{[\mathrm{oie}_{res}]_{\sim}}$.
Moreover, the following structural properties hold:
  1. 1.
    For complete sequential addition ($\oplus|_{\alpha}^{\beta}$):
    • The orbit space $\mathcal{O}(\mathrm{oie}\mathcal{S},\oplus|_{\alpha}^{\beta})$ consists of a single equivalence class (Property 19). Consequently, the whole orbit is associated with a unique set $\mathrm{operation}\mathcal{S}$.
    • According to Corollary 2, If the n-ary finitary operation $\odot$ satisfies permutation equivalence (i.e., its result is invariant under any reordering of its operands), then every expression in $\mathrm{operation}\mathcal{S}$ denotes the same operation. Otherwise, the expressions in $\mathrm{operation}\mathcal{S}$ may denote distinct operations.
  2. 2.
    For complete sequential multiplication ($\otimes$):
    • The orbit space $\mathcal{O}(\mathrm{oie}\mathcal{S},\otimes)$ may contain several equivalence classes (Property 20).
    • According to Corollary 3, for each equivalence class $\mathrm{oie}_{res}]_{\sim}$, the set $\mathrm{operation}\mathcal{S}_{[\mathrm{oie}_{res}]_{\sim}}$ contains exactly one expression. This expression is uniquely determined by the order of operands specified by any index tuple that produces a representative of the class.
Proof of Corollary 5.9.
We prove the three parts of the corollary in order.
1. Uniqueness within an equivalence class. Let $\mathrm{oie}_{res}]_{\sim}$ be an equivalence class in $\mathcal{O}(\mathrm{oie}\mathcal{S},*)$. Take any two OIE instances $\mathrm{oie}_{res_{1}},\mathrm{oie}_{res_{2}}\in[\mathrm{oie}_{res}]_{\sim}$. By definition, they are permutationally equivalent. Applying Corollary 1 to $\mathrm{oie}_{res_{1}}$ and $\mathrm{oie}_{res_{2}}$ shows that the projection (Definition 29) yields the same set of $n$-ary finitary operation expressions for both. Hence, all members of the equivalence class produce the same set, which we denote by $\mathrm{operation}\mathcal{S}_{[\mathrm{oie}_{res}]_{\sim}}$.
2. Properties for complete sequential addition. Property 19 states that the orbit space of complete sequential addition contains exactly one equivalence class. Therefore, the whole orbit corresponds to a single set $\mathrm{operation}\mathcal{S}$. Now assume that the operation $\odot$ satisfies permutation equivalence. Corollary 2 then implies that every expression in $\mathrm{operation}\mathcal{S}$ (which are precisely the permutations of operands that respect the grouping induced by ending timestamps) evaluates to the same result. If $\odot$ does not satisfy permutation equivalence, the expressions in $\mathrm{operation}\mathcal{S}$ may differ.
3. Properties for complete sequential multiplication. By Property 20, the orbit space of complete sequential multiplication may have more than one equivalence class. For an arbitrary class $\mathrm{oie}_{res}]_{\sim}$, choose any representative $\mathrm{oie}_{res}$. Corollary 3 asserts that the projection for $\mathrm{oie}_{res}$ produces a set $\mathrm{operation}\mathcal{S}_{[\mathrm{oie}_{res}]_{\sim}}$ with exactly one element. This is because the complete sequential multiplication forces the ending timestamps to be strictly ordered (each event must finish before the next one starts), leaving only one admissible ordering of the operands in the projection—namely the order prescribed by the index tuple that defines the operation. Hence, $\mathrm{operation}\mathcal{S}_{[\mathrm{oie}_{res}]_{\sim}}$ is a singleton.
The proof is now complete. ∎
6.4 Algebraic Logic of Speed-based Competitions
We now employ $\mathrm{OIE}$ and two sequential operations to explain why both a 100-meter dash and a downhill skiing race can determine a winner.
Consider two scenarios: a 100-meter dash with 8 athletes, and a downhill skiing race with 8 skiers. The 100-meter dash adopts a starting format in which all athletes line up on the track simultaneously, with the firing of the starting gun serving as the unified official start time. In contrast, downhill skiing requires that only one skier be on the course at any given time; hence the eight competitors must ski one after another, each having exclusive use of the course. For both competitions, the participant who records the shortest time is declared the champion.
We abstract a binary operation #\# on positive real numbers:
$x\,\#\,y=\min(x,y).$
This operation #\# satisfies closure, associativity, and commutativity. According to universal algebra, applying #\# to $n$ positive real numbers is an $n$-ary finitary operation $\min$” that satisfies permutation equivalence:
$\min(x_{1},x_{2},\dots,x_{n}).$
6.4.1 100-meter dash
Assume the starting gun is fired at 10:00:00 Beijing Time on January 7, 2026, corresponding to the Unix timestamp $\mathit{ts}_{\text{start}}=1736253600$. While there is no mandatory time limit for finishing a 100‑meter race, for our theoretical model we suppose that the race must be completed within $20$ seconds; hence the latest allowed finish timestamp is $\mathit{te}_{\max}=1736253620$.
We define 8 atomic Event instances representing each athlete’s performance:
$\{\,\textsf{atom}\textsf{Event}_{\text{athlete}_{1}},\,\textsf{atom}\textsf{Event}_{\text{athlete}_{2}},\,\dots,\,\textsf{atom}\textsf{Event}_{\text{athlete}_{8}}\,\}.$
Abstracting the interval component yields eight Event instances:
$\{\,\text{event}^{*}_{\text{athlete}_{1}},\,\text{event}^{*}_{\text{athlete}_{2}},\,\dots,\,\text{event}^{*}_{\text{athlete}_{8}}\,\}.$
Define an index tuple representing the initial lane assignment:
$\mathrm{idx}\mathcal{T}_{\text{asc}}=(1,2,3,4,5,6,7,8),$
meaning that $athlete_{i}$ is assigned to lane $i$. Assuming that the lane allocation does not affect any athlete’s performance, so thet are independent.
Following the competition rules we construct for each athlete’s OIE instance the third component $\mathcal{I}$ (the set of feasible intervals):
$\mathcal{I}=\{\,(\mathit{ts},\mathit{te})\mid 1736253600\leq\mathit{ts}<\mathit{te}\leq 1736253620\ \land\ 9.4\leq(\mathit{te}-\mathit{ts})\leq 20\,\}.$
The key feature here is that we affirm every athlete has the opportunity to challenge the assumed human physiological limit of $9.4$ seconds, to finish the race safely, and, most importantly, to enjoy “fairness”—all athletes have the chance to start at the same instant. Whether this can be achieved depends on the athletes’ own performances.
If one wishes to incorporate individual ability differences or potential lane effects, $\mathcal{F}$ could be adjusted accordingly; however, the lower and upper bounds remain unchanged because they correspond to the firing of the starting gun and the latest permissible finish time.
According to Definition 17, we build the mutually independent set of $\mathrm{OIE}\mathcal{S}$ instances
$\mathrm{oie}\mathcal{S}^{indep}=\{\mathrm{oie}_{1},\mathrm{oie}_{2},\dots,\mathrm{oie}_{8}\}$
where each
$\mathrm{oie}_{i}=\Bigl((),\{\,((\mathit{ts},\mathit{te}))\mid(\mathit{ts},\mathit{te})\in\mathcal{I}\,\},\mathcal{I},\{\,\textsf{event}^{*}_{i}\,\}\Bigr).$
Now we form a complete sequential addition operation:
$\mathrm{oie}_{100\text{dash}}=\oplus|^{1736253620}_{1736253600}(\mathrm{oie}_{1},\mathrm{oie}_{2},\dots,\mathrm{oie}_{8})$
(According to Definition 24, it can also be
$\mathrm{oie}_{100\text{dash}}=\oplus(\mathrm{oie}_{1},\mathrm{oie}_{2},\dots,\mathrm{oie}_{8})$
According to Definition 29 we construct the projection
$\oplus|^{1736253620}_{1736253600}(\mathrm{oie}\mathcal{S},\,\mathrm{idx}\mathcal{T}_{\text{asc}})\xrightarrow[\mathrm{event}\mathcal{S},\ \mathrm{operand}\mathcal{S},\ \min]{\text{ascending order of TSend}}\mathrm{operation}\mathcal{S}.$
The final result of the 100‑meter dash is obtained by ignoring individual reaction delays and directly subtracting the starting‑gun timestamp from the finish‑line timestamp. Let $F_{i}(\mathrm{oie}_{i})$ denote the result of athlete $i$; then the operation that determines the champion is the $n$-ary finitary operation
$\min\bigl(F_{1}(\mathrm{oie}_{1}),F_{2}(\mathrm{oie}_{2}),\dots,F_{8}(\mathrm{oie}_{8})\bigr).$
By Corollary 2, because $\min$ satisfies permutation equivalence, regardless of the order in which the individual finish timestamps occur, the resulting projected $n$-ary operation is always equivalent to the expression above.
Equivalence under different lane assignments.
The addition operation above corresponds to “athletes 1 to 8 are assigned to lanes 1 to 8, respectively”. If we wish to place athlete 1 in lane 2 and athlete 2 in lane 1, we would perform
$\mathrm{oie}_{100\text{dash}_{21}}=\oplus|^{1736253620}_{1736253600}(\mathrm{oie}_{2},\mathrm{oie}_{1},\mathrm{oie}_{3},\dots,\mathrm{oie}_{8}).$
If we reverse the order completely, we would perform
$\mathrm{oie}_{100\text{dash}_{\text{reverse}}}=\oplus|^{1736253620}_{1736253600}(\mathrm{oie}_{8},\mathrm{oie}_{7},\dots,\mathrm{oie}_{1}).$
These operations merely permute the order of the eight operands; hence they are permutationally equivalent
$\mathrm{oie}_{100\text{dash}}\stackrel{{\scriptstyle\mathcal{M}}}{{\sim}}\mathrm{oie}_{100\text{dash}_{21}}\stackrel{{\scriptstyle\mathcal{M}^{{}^{\prime}}}}{{\sim}}\mathrm{oie}_{100\text{dash}_{\text{reverse}}}$
Consequently, all of them belong to the same equivalence class in the orbit space $\mathcal{O}(\mathrm{oie}\mathcal{S},\oplus|^{1736253620}_{1736253600})$. By Corollary 4, because $\min$ satisfies permutation equivalence, no matter what order the individual finish timestamps take, the projected $n$-ary operation remains equivalent to
$\min\bigl(F_{1}(\mathrm{oie}_{1}),F_{2}(\mathrm{oie}_{2}),\dots,F_{8}(\mathrm{oie}_{8})\bigr).$
Thus, the specific lane assignment has no effect on the operational expression that determines the champion (and likewise on the ranking of each athlete).
6.4.2 Downhill Skiing
Downhill skiing is a winter sport where skiers start one after another in a predetermined order, and only one skier can be on the track at any given time. The competition rules require each athlete to ski from the start to the finish, and the one with the shortest time is the champion.
Assume the competition starts at 10:00:00 Beijing Time on January 7, 2026, corresponding to the Unix timestamp $ts_{\text{start}}=1736253600$. The competition has a total time limit: the last skier must finish within 20 minutes, so the latest allowed finish timestamp is $te_{\max}=ts_{\text{start}}+1200=1736254800$.
Each skier’s skiing time has physiological and technical lower and upper bounds. Let the fastest possible time for athlete $i$ be $T_{\min}^{(i)}=90$ seconds (1.5 minutes) and the slowest possible time be $T_{\max}^{(i)}=120$ seconds (2 minutes). For simplicity, we assume that all athletes have the same feasible time intervals, but the model can easily be extended to personalized intervals.
We define eight AtomEvent instances representing real‑world skiing competitions:
$\{\,\textsf{atom}\textsf{Event}_{\text{skier}_{1}},\,\textsf{atom}\textsf{Event}_{\text{skier}_{2}},\,\ldots,\,\textsf{atom}\textsf{Event}_{\text{skier}_{8}}\,\}.$
Abstracting the temporal property yields eight Event instances:
$\{\,\mathrm{event}^{*}_{\text{skier}_{1}},\,\mathrm{event}^{*}_{\text{skier}_{2}},\,\ldots,\,\mathrm{event}^{*}_{\text{skier}_{8}}\,\}.$
Define an index tuple representing the competition start order (this order is usually determined by preliminary results; here we assume $1,2,\ldots,8$):
$\mathrm{idx}\mathcal{T}_{\text{asc}}=(1,2,3,4,5,6,7,8),$
meaning athlete $i$ starts in the ii‑th position. According to the competition rules, we assume there are no physical or logical conflicts that would make a combination infeasible, so we set
$\mathfrak{f}\mathrm{Inf2tupleTS}((,\mathrm{oie}\mathcal{S}),\,\mathrm{idx}\mathcal{T}_{\text{asc}})=\emptyset.$
Next we construct the $\mathrm{OIE}$ instances for each athlete. For $skier_{i}$, the set of feasible intervals (the third element $\mathcal{I}$ of the OIE instance) must satisfy: 1. The interval must lie within the total competition time window $ts_{\text{start}},te_{\text{end}}]$. 2. The interval length (i.e., the skiing time) must be within the skier’s personal capability range $T_{\min}^{(i)},T_{\max}^{(i)}]$.
Hence, for each $i$ we define
$\mathcal{F}_{i}=\{\,(ts,te)\mid ts_{\text{start}}\leq ts<te\leq te_{\text{end}}\ \wedge\ T_{\min}^{(i)}\leq(te-ts)\leq T_{\max}^{(i)}\,\}.$
We form the $\mathrm{OIE}\mathcal{S}$ instance $\mathrm{oie}\mathcal{S}=\{\mathrm{oie}_{1},\ldots,\mathrm{oie}_{8}\}$, where
$\mathrm{oie}_{i}=\Bigl((\,),\ \{\,((ts,te))\mid(ts,te)\in\mathcal{F}_{i}\,\},\ \mathcal{F}_{i},\ \{\,\mathrm{event}^{*}_{\text{skier}_{i}}\,\}\Bigr).$
The core feature of downhill skiing is that athletes must ski one after another; only one person may be on the course at a time. This exactly corresponds to the “unfair” sequentiality emphasized by Complete Sequential Multiplication.
We use the complete sequential multiplication operation $\otimes$ to combine the eight skiers’ $\mathrm{OIE}$ instances according to the given start order $\mathrm{idx}\mathcal{T}_{\text{asc}}$:
$\mathrm{oie}_{\text{downhill}}=\otimes(\mathrm{oie}_{1},\mathrm{oie}_{2},\ldots,\mathrm{oie}_{8}).$
Following Definition 29, we construct the projection from the sequential‑operation result to a set of nn‑ary finitary operations. Let: - $\mathrm{event}\mathcal{S}=\{\textsf{event}_{\text{skier}_{1}},\ldots,\textsf{event}_{\text{skier}_{8}}\}$ be the set of actual event instances. - $\mathrm{operand}\mathcal{S}=\{F_{1}(\mathrm{oie}_{1}),\ldots,F_{8}(\mathrm{oie}_{8})\}$ be the set of operands, where $F_{i}(\mathrm{oie}_{i})$ denotes the competition time of athlete $i$ (i.e., the length tetste-ts of its execution interval). - The operation symbol be $\min$ (take the minimum).
The projection is written as
$\otimes(\mathrm{oie}\mathcal{S},\,\mathrm{idx}\mathcal{T}_{\text{asc}})\xrightarrow[\text{ascending order of TSend}]{\mathrm{event}\mathcal{S},\ \mathrm{operand}\mathcal{S},\ \min}\mathrm{operation}\mathcal{S}.$
By Property 22, because complete sequential multiplication forces a strict order (each athlete’s finish time determines that it must end before the next athlete starts), the resulting operation set $\operatorname{operation}\mathcal{S}$ contains exactly one expression:
$\min\bigl(F_{1}(\operatorname{oie}_{1}),F_{2}(\operatorname{oie}_{2}),\ldots,F_{8}(\operatorname{oie}_{8})\bigr).$
This is precisely the operation used in downhill skiing to determine the champion: compare the competition times of all athletes and take the minimum.
Unlike the 100‑meter dash, the competition‑result operation in downhill skiing depends on the start order. If the start order changes, e.g., swapping athletes 1 and 2:
$\operatorname{idx}\mathcal{T}^{\prime}=(2,1,3,4,5,6,7,8),$
then the result $\operatorname{oie}_{\text{downhill}}^{\prime}$ of complete sequential multiplication is not necessarily permutation equivalent to the original $\operatorname{oie}_{\text{downhill}}$ (by Property 16). Consequently, the projected operation set may also differ (although it still contains only one expression, the order of the operands may change). However, because the $\min$ operation satisfies permutation equivalence (exchanging the order of operands does not affect the resulting value), the mathematical result for the champion remains unchanged regardless of the start order.
6.4.3 Conclusion
By introducing the concepts of Optional Intervals Events (OIE) and the two sequential operations (Complete Sequential Addition $\oplus$ and Complete Sequential Multiplication $\otimes$), we provide a formal method for subjective planning in analyzing the execution of events in the physical world. This chapter focuses on using this framework to explain why two seemingly different types of speed competitions—the 100-meter dash (concurrent type) and downhill skiing (sequential type)—can both determine a champion.
1. Core Modeling and Differences
The 100-meter dash is modeled as Complete Sequential Addition $\oplus$. This operation emphasizes “fairness”: within a given start and end time domain, all participants have an equal opportunity to start simultaneously. The interval combinations in the $\mathcal{F}$ element of the operation result allow for any permutation of the athletes’ ending timestamps, corresponding to the fact that in reality, the order in which athletes cross the finish line can be arbitrary.
Downhill skiing is modeled as Complete Sequential Multiplication $\otimes$. This operation emphasizes “sequentiality”: participants must start strictly in a given order, and the end time of one athlete does not exceed the start time of the next. The interval combinations in the $\mathcal{F}$ element of the operation result have ending timestamps that strictly follow the non-decreasing order of the operands (athletes).
2. Algebraic Operations for Determining the Winner
The common goal of both competitions is to find the minimum completion time among all participants, which corresponds to the champion. This corresponds to a binary operation #\#, defined as $x\ \#\ y=\min(x,y)$. Extended to $n$-ary, this is the minimum operation $\min$.
- Algebraic properties: The operation $\min$ satisfies closure, associativity, and permutation equivalence. That is, for any sequence $(x_{1},\ldots,x_{n})$ and any of its permutations $(x_{\sigma(1)},\ldots,x_{\sigma(n)})$, we have:
$\min(x_{1},\ldots,x_{n})=\min(x_{\sigma(1)},\ldots,x_{\sigma(n)}).$
Therefore, $\min$ is an $n$-ary operation on a commutative semigroup.
3. Operational Equivalence and Determinacy of the Champion
- For the 100-meter dash ($\oplus$): Although the lane assignment (order of operands) can be arbitrarily swapped, leading to different OIE results such as $\operatorname{oie}_{100\text{dash}}$, $\operatorname{oie}_{100\text{dash}_{21}}$, etc., according to Property 3.18, these results all belong to the same equivalence class in the orbit space under permutational equivalence. According to Corollary 5.7, because $\min$ satisfies permutation equivalence, the set of $n$-ary operations $\operatorname{operation}\mathcal{S}$ obtained by projecting from any OIE instance in this equivalence class contains expressions (i.e., $\min$ operations corresponding to all possible finish time orders) that all compute the same value. Therefore, regardless of the actual order of crossing the finish line, the champion is determined by a single, order-independent expression $\min(F_{1},\ldots,F_{n})$.
- For downhill skiing ($\otimes$): The start order is strictly specified by the index tuple $\operatorname{idx}\mathcal{T}$. According to Corollary 5.8, the operation set $\operatorname{operation}\mathcal{S}$ obtained by projecting from the multiplication result $\operatorname{oie}_{\text{downhill}}$ is a singleton set containing only the expression $\min(F_{i_{1}},\ldots,F_{i_{n}})$, where the order matches $\operatorname{idx}\mathcal{T}$. If the start order is changed (e.g., swapping the first two athletes), according to Property 3.19, the resulting OIE may not be permutationally equivalent to the original, so the projected expression may also change in order (e.g., becoming $\min(F_{i_{2}},F_{i_{1}},\ldots)$). However, because the $\min$ operation itself has permutation equivalence, no matter the order of operands in the expression, its computed result (the minimum time) remains unchanged. Therefore, the champion is still uniquely and determinately identified.
4. Conclusion
Although the 100-meter dash and downhill skiing are conceptually opposite in terms of “fairness” and “sequentiality,” their essence as speed competitions—comparing time costs by total ordering—is algebraically consistent. The key point is that the operation that determines the winner, $\min$, is an $n$-ary operation that satisfies permutation equivalence. This algebraic property ensures that:
1. For competitions emphasizing concurrent fairness ($\oplus$), the operation result is not affected by the instantaneous arrangement of participants, thereby supporting the idealized concept of “starting simultaneously.” 2. For competitions emphasizing strict order ($\otimes$), although the operation result depends on the preset order structure, its permutation equivalence ensures that the final comparison result (champion determination) is not affected by the specific arrangement of that order.
Therefore, both competition forms can reliably determine a champion. The fundamental reason is not that they share some specific physical or operational property, but that they both ultimately reduce to the same algebraic operation with strong symmetry (permutation invariance). The OIE and sequential operation framework proposed in this research, by formalizing subjective planning, not only explains this phenomenon but also provides a new theoretical foundation beyond passive observation for broader event scheduling, physical simulation, and computational modeling.
6.5 Algebraic logic in parallel sorting
Array sorting is a pivotal research area in computer science. In this subsection, we employ $\mathrm{OIE}$, Complete Sequential Addition and Complete Sequential Multiplication to analyze the implementation of parallel array sorting via an algebraic approach. Through this example, we demonstrate that $\mathrm{OIE}$ and sequential operations can not only reshape our perspectives on observing the real world in the physical domain, but also directly incorporate subjectivity into program design and many industrial applications.
Suppose there is an array
$arr=(499,375,233,168,100,57,28,14)$
We sort it in ascending order
$(499,375,233,168,100,57,28,14)\rightarrow(14,28,57,100,168,233,375,499)$
In fact, the essence of sorting is to multiply the original sequence by a permutation matrix to yield an ordered sequence, and this example is as following
$\left(499,375,233,168,100,57,28,14\right)\times\color[rgb]{0.75,0.75,0.75}\begin{pmatrix}0&0&0&0&0&0&0&\textbf{\color[rgb]{0,0,0}1}\\ 0&0&0&0&0&0&\textbf{\color[rgb]{0,0,0}1}&0\\ 0&0&0&0&0&\textbf{\color[rgb]{0,0,0}1}&0&0\\ 0&0&0&0&\textbf{\color[rgb]{0,0,0}1}&0&0&0\\ 0&0&0&\textbf{\color[rgb]{0,0,0}1}&0&0&0&0\\ 0&0&\textbf{\color[rgb]{0,0,0}1}&0&0&0&0&0\\ 0&\textbf{\color[rgb]{0,0,0}1}&0&0&0&0&0&0\\ \textbf{\color[rgb]{0,0,0}1}&0&0&0&0&0&0&0\\ \end{pmatrix}\color[rgb]{0,0,0}=\left(14,28,57,100,168,233,375,499\right)$
In standard merge sort algorithm, merge is the fundamental operation at the core. The merging rule of the standard merge sort is to merge only adjacent sorted subintervals, and the merging intervals in each pass are all pairwise combinations of the sorted subintervals obtained from the previous pass. Thus, all merging intervals within the same pass are inevitably non-overlapping and disjoint.
Because it is essentially a sequence multiplied by a transposition matrix for the action of merge, we use matrix
$\mathcal{M}_{([i_{l},i_{r}],[j_{l},j_{r}])}^{(n)}$
to express the transposition matrix for merging two sorted subsequences. In the notation for this matrix variable, $i_{l}$, $i_{r}$ are the left and right boundary indices of one sorted subsequence, and $j_{l}$, $j_{r}$ are those of the other sorted subsequence (We use 1 as the starting index here). For example, the form of $\mathcal{M}_{([1,1],[2,2])}^{(8)}$ in this case is
$\mathcal{M}_{([1,1],[2,2])}^{(8)}=\color[rgb]{0.75,0.75,0.75}\begin{pmatrix}0&\textbf{\color[rgb]{0,0,0}1}&0&0&0&0&0&0\\ \textbf{\color[rgb]{0,0,0}1}&0&0&0&0&0&0&0\\ 0&0&\textbf{\color[rgb]{0,0,0}1}&0&0&0&0&0\\ 0&0&0&\textbf{\color[rgb]{0,0,0}1}&0&0&0&0\\ 0&0&0&0&\textbf{\color[rgb]{0,0,0}1}&0&0&0\\ 0&0&0&0&0&\textbf{\color[rgb]{0,0,0}1}&0&0\\ 0&0&0&0&0&0&\textbf{\color[rgb]{0,0,0}1}&0\\ 0&0&0&0&0&0&0&\textbf{\color[rgb]{0,0,0}1}\\ \end{pmatrix}$
then the merge of 1st and 2nd elements can be expressed as
$\left(499,375,233,168,100,57,28,14\right)\times\mathcal{M}_{([1,1],[2,2])}^{(8)}=\left(375,499,233,168,100,57,28,14\right)$
Then, we use transposition matrix multiplication to denote merge sort for the entire sequence. Its process is
[…]
Translate this process into the form of matrix multiplication
[…]
$\times\color[rgb]{0.75,0.75,0.75}\begin{pmatrix}0&\textbf{\color[rgb]{0,0,0}1}&0&0&0&0&0&0\\ \textbf{\color[rgb]{0,0,0}1}&0&0&0&0&0&0&0\\ 0&0&\textbf{\color[rgb]{0,0,0}1}&0&0&0&0&0\\ 0&0&0&\textbf{\color[rgb]{0,0,0}1}&0&0&0&0\\ 0&0&0&0&\textbf{\color[rgb]{0,0,0}1}&0&0&0\\ 0&0&0&0&0&\textbf{\color[rgb]{0,0,0}1}&0&0\\ 0&0&0&0&0&0&\textbf{\color[rgb]{0,0,0}1}&0\\ 0&0&0&0&0&0&0&\textbf{\color[rgb]{0,0,0}1}\\ \end{pmatrix}\times\begin{pmatrix}\textbf{\color[rgb]{0,0,0}1}&0&0&0&0&0&0&0\\ 0&\textbf{\color[rgb]{0,0,0}1}&0&0&0&0&0&0\\ 0&0&0&\textbf{\color[rgb]{0,0,0}1}&0&0&0&0\\ 0&0&\textbf{\color[rgb]{0,0,0}1}&0&0&0&0&0\\ 0&0&0&0&\textbf{\color[rgb]{0,0,0}1}&0&0&0\\ 0&0&0&0&0&\textbf{\color[rgb]{0,0,0}1}&0&0\\ 0&0&0&0&0&0&\textbf{\color[rgb]{0,0,0}1}&0\\ 0&0&0&0&0&0&0&\textbf{\color[rgb]{0,0,0}1}\\ \end{pmatrix}$
$\times\color[rgb]{0.75,0.75,0.75}\begin{pmatrix}\textbf{\color[rgb]{0,0,0}1}&0&0&0&0&0&0&0\\ 0&\textbf{\color[rgb]{0,0,0}1}&0&0&0&0&0&0\\ 0&0&\textbf{\color[rgb]{0,0,0}1}&0&0&0&0&0\\ 0&0&0&\textbf{\color[rgb]{0,0,0}1}&0&0&0&0\\ 0&0&0&0&0&\textbf{\color[rgb]{0,0,0}1}&0&0\\ 0&0&0&0&\textbf{\color[rgb]{0,0,0}1}&0&0&0\\ 0&0&0&0&0&0&\textbf{\color[rgb]{0,0,0}1}&0\\ 0&0&0&0&0&0&0&\textbf{\color[rgb]{0,0,0}1}\\ \end{pmatrix}\times\begin{pmatrix}\textbf{\color[rgb]{0,0,0}1}&0&0&0&0&0&0&0\\ 0&\textbf{\color[rgb]{0,0,0}1}&0&0&0&0&0&0\\ 0&0&\textbf{\color[rgb]{0,0,0}1}&0&0&0&0&0\\ 0&0&0&\textbf{\color[rgb]{0,0,0}1}&0&0&0&0\\ 0&0&0&0&\textbf{\color[rgb]{0,0,0}1}&0&0&0\\ 0&0&0&0&0&\textbf{\color[rgb]{0,0,0}1}&0&0\\ 0&0&0&0&0&0&0&\textbf{\color[rgb]{0,0,0}1}\\ 0&0&0&0&0&0&\textbf{\color[rgb]{0,0,0}1}&0\\ \end{pmatrix}$
$\times\color[rgb]{0.75,0.75,0.75}\begin{pmatrix}0&0&\textbf{\color[rgb]{0,0,0}1}&0&0&0&0&0\\ 0&0&0&\textbf{\color[rgb]{0,0,0}1}&0&0&0&0\\ \textbf{\color[rgb]{0,0,0}1}&0&0&0&0&0&0&0\\ 0&\textbf{\color[rgb]{0,0,0}1}&0&0&0&0&0&0\\ 0&0&0&0&\textbf{\color[rgb]{0,0,0}1}&0&0&0\\ 0&0&0&0&0&\textbf{\color[rgb]{0,0,0}1}&0&0\\ 0&0&0&0&0&0&\textbf{\color[rgb]{0,0,0}1}&0\\ 0&0&0&0&0&0&0&\textbf{\color[rgb]{0,0,0}1}\\ \end{pmatrix}\times\begin{pmatrix}\textbf{\color[rgb]{0,0,0}1}&0&0&0&0&0&0&0\\ 0&\textbf{\color[rgb]{0,0,0}1}&0&0&0&0&0&0\\ 0&0&\textbf{\color[rgb]{0,0,0}1}&0&0&0&0&0\\ 0&0&0&\textbf{\color[rgb]{0,0,0}1}&0&0&0&0\\ 0&0&0&0&0&0&\textbf{\color[rgb]{0,0,0}1}&0\\ 0&0&0&0&0&0&0&\textbf{\color[rgb]{0,0,0}1}\\ 0&0&0&0&\textbf{\color[rgb]{0,0,0}1}&0&0&0\\ 0&0&0&0&0&\textbf{\color[rgb]{0,0,0}1}&0&0\\ \end{pmatrix}$
$\times\color[rgb]{0.75,0.75,0.75}\begin{pmatrix}0&0&0&0&\textbf{\color[rgb]{0,0,0}1}&0&0&0\\ 0&0&0&0&0&\textbf{\color[rgb]{0,0,0}1}&0&0\\ 0&0&0&0&0&0&\textbf{\color[rgb]{0,0,0}1}&0\\ 0&0&0&0&0&0&0&\textbf{\color[rgb]{0,0,0}1}\\ \textbf{\color[rgb]{0,0,0}1}&0&0&0&0&0&0&0\\ 0&\textbf{\color[rgb]{0,0,0}1}&0&0&0&0&0&0\\ 0&0&\textbf{\color[rgb]{0,0,0}1}&0&0&0&0&0\\ 0&0&0&\textbf{\color[rgb]{0,0,0}1}&0&0&0&0\\ \end{pmatrix}$
[…]
Abbreviated as
[…]
$\quad\times\mathcal{M}_{([1,1],[2,2])}^{(8)}\times\mathcal{M}_{([3,3],[4,4])}^{(8)}\times\mathcal{M}_{([5,5],[6,6])}^{(8)}\times\mathcal{M}_{([7,7],[8,8])}^{(8)}$
$\quad\times\mathcal{M}_{([1,2],[3,4])}^{(8)}\times\mathcal{M}_{([5,6],[7,8])}^{(8)}$
$\quad\times\mathcal{M}_{([1,4],[5,8])}^{(8)}$
[…]
For the merging pass with the merging unit of length $k$ and the merging range of $2k$, the total number of merging tasks is $\lfloor\frac{n}{2k}\rfloor$ for a sequence containing $n$ elements.
The elements involved in different merging operations within the same merging pass are disjoint. Thus, we summarize the rule as follows:
$For\ any\ two\ matrices\ \mathcal{M}_{([a_{l},a_{r}],[b_{l},b_{r}])}^{(n)},\mathcal{M}_{([c_{l},c_{r}],[d_{l},d_{r}])}^{(n)}$
$\ in\ the\ same\ merging\ pass,\ the\ property$
$\quad\{x|x\in[a_{l},b_{r}]\}\cap\{x|x\in[c_{l},d_{r}]\}=\emptyset$
$holds.$
For example, when $k=1$, there are 4 merging operations in this pass:
$\text{merging of position 1 and position 2}:\mathcal{M}_{([1,1],[2,2])}^{(8)},$
$\text{merging of position 3 and position 4}:\mathcal{M}_{([3,3],[4,4])}^{(8)},$
$\text{merging of position 5 and position 6}:\mathcal{M}_{([5,5],[6,6])}^{(8)},$
$\text{merging of position 7 and position 8}:\mathcal{M}_{([7,7],[8,8])}^{(8)}$
This is represented as
$\times\mathcal{M}_{([1,1],[2,2])}^{(8)}\times\mathcal{M}_{([3,3],[4,4])}^{(8)}\times\mathcal{M}_{([5,5],[6,6])}^{(8)}\times\mathcal{M}_{([7,7],[8,8])}^{(8)}\\$
it follows that
$\{1,2\}\cap\{3,4\}\cap\{5,6\}\cap\{7,8\}=\emptyset$
Similarly, for
$\times\mathcal{M}_{([1,2],[3,4])}^{(8)}\times\mathcal{M}_{([5,6],[7,8])}^{(8)}$
it follows that
$\{1,4\}\cap\{5,8\}=\emptyset$
Thus, we can conclude that in the matrix multiplication corresponding to each pass of Merge Sort, the index positions undergoing transposition for each transposition matrix operation are disjoint. This indicates that the matrix multiplication corresponding to each pass satisfies the commutative law. The corresponding practical operation is that one may choose to merge positions 1 and 2 first and positions 7 and 8 last, or alternatively merge positions 7 and 8 first and positions 1 and 2 last.
Then, we can leverage the commutative law to add parentheses to the matrix multiplication expression
$\left(\mathcal{M}_{([1,1],[2,2])}^{(8)}\times\mathcal{M}_{([3,3],[4,4])}^{(8)}\times\mathcal{M}_{([5,5],[6,6])}^{(8)}\times\mathcal{M}_{([7,7],[8,8])}^{(8)}\right)$
$\times\left(\mathcal{M}_{([1,2],[3,4])}^{(8)}\times\mathcal{M}_{([5,6],[7,8])}^{(8)}\right)$
$\times\mathcal{M}_{([1,4],[5,8])}^{(8)}$
Matrix multiplications inside the parentheses satisfy the commutative law, while matrix operations outside the parentheses do not.
We now construct an $\mathrm{OIE}\mathcal{S}$ instance for all matrices in the matrix multiplication of Merge Sort
$\mathrm{oie}\mathcal{S}_{merge}=\{\mathrm{oie}_{11\_22},\ \mathrm{oie}_{33\_44},\ \mathrm{oie}_{55\_66},\ \mathrm{oie}_{77\_88},\ \mathrm{oie}_{12\_34},\ \mathrm{oie}_{56\_78},\ \mathrm{oie}_{14\_58}\}$
an $\mathrm{Operand}\mathcal{S}$ instance
$\mathrm{operand}\mathcal{S}_{merge}=\{\ \mathcal{M}_{([1,1],[2,2])}^{(8)},\ \mathcal{M}_{([3,3],[4,4])}^{(8)},\ \mathcal{M}_{([5,5],[6,6])}^{(8)},$
$\qquad\mathcal{M}_{([7,7],[8,8])}^{(8)},\ \mathcal{M}_{([1,2],[3,4])}^{(8)},\ \mathcal{M}_{([5,6],[7,8])}^{(8)},\mathcal{M}_{([1,4],[5,8])}^{(8)}\ \}$
and an $\mathrm{Event}\mathcal{S}$ instance
$\mathrm{event}\mathcal{S}_{merge}=\{\textsf{event}_{11\_22},\ \textsf{event}_{33\_44},\ \textsf{event}_{55\_66},\ \textsf{event}_{77\_88},\ \textsf{event}_{12\_34},\ \textsf{event}_{56\_78},\ \textsf{event}_{14\_58}\}$
According to Definition 29 and Corollary 4, we can use following sequential operation to build projection using only complete sequential multiplication
$\mathrm{oie}_{merge}=\otimes\left(\otimes(\mathrm{oie}_{11\_22},\mathrm{oie}_{33\_44},\mathrm{oie}_{55\_66},\mathrm{oie}_{77\_88}),\ \otimes(\mathrm{oie}_{12\_34},\mathrm{oie}_{56\_78}),\ \mathrm{oie}_{14\_58}\right)$
or using complete sequential multiplication outside the parentheses while complete sequential addition inside the parentheses
$\mathrm{oie}_{merge}=\otimes\left(\oplus|_{\alpha_{1}}^{\beta_{1}}(\mathrm{oie}_{11\_22},\mathrm{oie}_{33\_44},\mathrm{oie}_{55\_66},\mathrm{oie}_{77\_88}),\ \oplus|_{\alpha_{2}}^{\beta_{2}}(\mathrm{oie}_{12\_34},\mathrm{oie}_{56\_78}),\ \mathrm{oie}_{14\_58}\right)$
In cases where both complete sequential addition and complete sequential multiplication are employed, the part running complete sequential addition theoretically underpin the practical applicability of parallel technologies. Since the time complexity of all merge tasks in a single merge operation is of the same order of magnitude, for a certain additive subexpression in the arithmetic expression, we directly set the filter domain to the range corresponding to the time complexity of a single merge task. In this way, we transform the parallelization of merge sort into the following problem:
Analyze the time complexity of parallel merge sort for sorting $n$ elements with $k$ processes ($1\leq k\leq n$) under the constraint that only tasks on the same layer of the divide-and-conquer binary tree are executed in parallel, while different layers are processed serially (the next layer starts only after all tasks of the current layer complete).
Merge sort’s execution forms a divide-and-conquer binary tree with $\log_{2}n$ layers (indexed by $d$, $0\leq d\leq\log_{2}n-1$). The $d$-th layer has $2^{d}$ subtasks, each processing a subarray of length $n/2^{d}$, and the total elements processed per layer is always $n$ ($2^{d}\times n/2^{d}=n$). For the $d$-th layer, the time complexity $T(d)$ depends on $k$ and $2^{d}$: - If $2^{d}\leq k$ (sufficient processes), all subtasks run in parallel, so $T(d)=\mathcal{O}(n/2^{d})$. - If $2^{d}>k$ (insufficient processes), subtasks run in batches of $k$, so $T(d)=\lceil 2^{d}/k\rceil\times\mathcal{O}(n/2^{d})=\mathcal{O}(n/k)$ (constant factors are ignored in big-$\mathcal{O}$).
Let $d_{0}=\log_{2}k$ (critical layer where $2^{d_{0}}=k$). We split the tree into pre-critical layers ($0\leq d\leq d_{0}-1$, $2^{d}<k$) and post-critical layers ($d_{0}\leq d\leq\log_{2}n-1$, $2^{d}\geq k$). The total time of pre-critical layers is:
$T_{\text{pre}}=\sum_{d=0}^{\log_{2}k-1}\mathcal{O}\left(\frac{n}{2^{d}}\right)=\mathcal{O}\left(n\times\sum_{d=0}^{\log_{2}k-1}\frac{1}{2^{d}}\right)=\mathcal{O}(n),$
since the geometric series sum converges to a constant (2). The total number of post-critical layers is $\log_{2}n-\log_{2}k=\log_{2}(n/k)$, so their total time is:
$T_{\text{post}}=\log_{2}\left(\frac{n}{k}\right)\times\mathcal{O}\left(\frac{n}{k}\right)=\mathcal{O}\left(\frac{n}{k}\log n\right)$
where $\log_{2}(n/k)$ is simplified to $\log n$ (logarithm base does not affect big-$\mathcal{O}$).
The total time complexity is the sum of $T_{\text{pre}}$ and $T_{\text{post}}$:
$\boxed{T_{\text{total}}=\mathcal{O}(n)+\mathcal{O}\left(\frac{n}{k}\log n\right)}$
For $1\leq k\leq n$ and large $n$, $\mathcal{O}(n/k\log n)$ is the dominant term, so the final time complexity is:
$\boxed{T_{\text{total}}=\mathcal{O}\left(\frac{n}{k}\log n\right)}.$
An illustrative example: $n=16$ ($\log_{2}16=4$), $k=4$ ($d_{0}=2$). Pre-critical layers ($d=0,1$) have $T(0)=\mathcal{O}(16)$, $T(1)=\mathcal{O}(8)$ (sum $\mathcal{O}(n)$); post-critical layers ($d=2,3$) have $T(2)=\mathcal{O}(4)$, $T(3)=\mathcal{O}(4)$ (sum $\mathcal{O}(16)$), matching $\mathcal{O}(n/k\log n)=\mathcal{O}(16/4\times 4)=\mathcal{O}(16)$.
In conclusion, under the constraint of parallel execution on the same layer and serial execution across layers, the time complexity of merge sort for $n$ elements with $k$ processes ($1\leq k\leq n$) is $\mathcal{O}\left(\frac{n}{k}\log n\right)$, where a larger $k$ (within $n$) reduces the complexity, and pre-critical layers contribute only a negligible linear term for large $n$.
This concludes our analysis of parallelized merge sort. In fact, for all sorting algorithms that fall into the category of selection sort or those based on the divide-and-conquer method—such as bubble sort and Shell sort—once we analyze the sorting steps using array multiplication, we can identify matrix multiplication subexpressions that satisfy the commutative law to varying degrees in the operational formulas. Thus, with the framework of Optional Intervals Event (OIE), we can rigorously derive feasible parallelization strategies for all such algorithms.
From a broader perspective, matrix multiplication is a method widely applied in engineering practices, including data processing, object rotation and translation, and many other fields. Matrix multiplication itself is merely a form of algebraic operation, and numerous operations across all engineering fields can be abstracted into algebraic expressions of n-ary operations in key scenarios. It is precisely in these scenarios that the concepts of Optional Intervals Event($\mathrm{OIE}$) and sequential operations can be effectively introduced.
6.6 Applications in Probability Theory and Stochastic Distributions
Consider a bag containing $n$ balls, of which $m$ are red and nmn-m are black Let $k$ participants ($1<k\leq m<n$) each draw one ball without replacement. We distinguish two sampling scenarios: simultaneous sampling, where all $k$ participants reach into the bag concurrently, and sequential sampling, where participants draw one after another according to a predetermined order.
Classical probability theory establishes that each participant’s marginal probability of drawing a red ball is $m/n$ in both scenarios. For simultaneous sampling, this is due to the constrained, non-repetitive random distribution; for sequential sampling, it follows from Bayes’ theorem.
The $\mathrm{OIE}$ framework reveals that this equivalence of outcomes arises despite fundamentally distinct algebraic architectures: simultaneous sampling corresponds to Complete Sequential Addition ($\oplus$) with its single-orbit permutation equivalence, while sequential sampling corresponds to Complete Sequential Multiplication ($\otimes$) with its strict sequential ordering and potentially multiple orbital equivalence classes. The key insight is that although the algebraic paths differ, both operations project onto the same probability distribution in the observable outcome space.
We model each participant $i$ as an $\textsf{Atom}\mathrm{OIE}$ instance $\textsf{atom}\mathrm{Oie}_{i}$ mapping to the atomic event $\textsf{atom}\textsf{Event}_{i}$ (the action of drawing a ball). Each $\textsf{atom}\mathrm{Oie}_{i}$ carries a feasible interval set $\mathcal{I}_{\textsf{atom}\mathrm{Oie}_{i}}=\{(\mathrm{ts},\mathrm{te})\mid\alpha\leq\mathrm{ts}<\mathrm{te}\leq\beta\land\mathrm{te}-\mathrm{ts}=\delta\}$, where $\alpha,\beta)$ denotes the shared temporal domain of the sampling process and $\delta>0$ represents the non-zero duration required to physically execute the draw.
In the simultaneous scenario, all $k$ participants possess equal opportunity to commence their draw within the shared domain $\alpha,\beta)$, with no privileged order of precedence. This embodies the semantics of Complete Sequential Addition $\oplus|_{\alpha}^{\beta}$ (Definition 20). Executing the operation
$\mathrm{oie}_{\mathsf{simul}}=\oplus|_{\alpha}^{\beta}(\textsf{atom}\mathrm{Oie}_{1},\textsf{atom}\mathrm{Oie}_{2},\ldots,\textsf{atom}\mathrm{Oie}_{k})$
yields a result whose orbit space $\mathcal{O}(\mathrm{oie}\mathcal{S},\oplus|_{\alpha}^{\beta})$ contains exactly one equivalence class (Property 19). Consequently, for any two index tuples $\mathrm{idx}\mathcal{T}_{1}$ and $\mathrm{idx}\mathcal{T}_{2}$, the results are permutationally equivalent:
$\oplus|_{\alpha}^{\beta}(\mathrm{oie}\mathcal{S},\mathrm{idx}\mathcal{T}_{1})\stackrel{{\scriptstyle\mathcal{M}}}{{\sim}}\oplus|_{\alpha}^{\beta}(\mathrm{oie}\mathcal{S},\mathrm{idx}\mathcal{T}_{2}).$
This algebraic symmetry reflects that the assignment of indices represents merely a descriptive convention, not an ontological distinction, since all participants act within the same temporal window without sequential constraint.
Conversely, the sequential scenario enforces a strict order wherein participant $i$ must complete their draw before participant $i+1$ begins. This corresponds to Complete Sequential Multiplication $\otimes$ (Definition 22), which applies an ascending ordered filter (Definition 21) requiring $\mathrm{te}_{i}\leq\mathrm{ts}_{i+1}$ for all consecutive pairs. The operation
$\mathrm{oie}_{\mathsf{seq}}=\otimes(\textsf{atom}\mathrm{Oie}_{1},\textsf{atom}\mathrm{Oie}_{2},\ldots,\textsf{atom}\mathrm{Oie}_{k})$
generates a distinct algebraic structure: by Property 16, changing the index order generally produces a result that is not permutationally equivalent to the original. Thus the orbit space $\mathcal{O}(\mathrm{oie}\mathcal{S},\otimes)$ may contain multiple equivalence classes (Property 20), each corresponding to a specific permutation of the drawing order. The algebraic structure explicitly encodes the temporal “unfairness” of precedence.
But decisively, despite these structural differences, both operations yield identical probability distributions when projected onto the outcome space. Under $\oplus$, the single-orbit property ensures that every participant appears in every ordinal position with equal frequency across the feasible set $\mathcal{F}$, directly yielding the uniform marginal probability $m/n$. Under $\otimes$, the same marginal probability emerges from the combinatorial properties of conditional probability chains: the product of conditional probabilities arranged in any order (via Bayes’ theorem) preserves the individual marginal $m/n$ for each position.
This distinction becomes clear when examining the projection to $n$-ary operations (Definition 28). For simultaneous sampling, the projection generates a set of $k$-ary operations respecting permutation equivalence (Corollary 2), reflecting that the expression’s order of drawing does not affect the final allocation. For sequential sampling, the projection yields a single ordered expression (Property 22), yet when the operation $\odot$ for the $\mathrm{operation}\mathcal{S}$ in the projection is defined as probability calculation (specifically, the counting measure over red ball allocations), both projections evaluate to the identical probability value $m/n$.
This contrasts with the speed competitions analyzed in Section 6.4, where $\oplus$ (100-meter dash) and $\otimes$ (downhill skiing) yield different operational characteristics regarding lane assignments and starting orders. In the sampling scenario, the physical indistinguishability of balls (except for color) and the symmetry of the sampling frame erase the operational distinctions between concurrent and sequential execution at the level of final probability distributions. The OIE framework thus illuminates how identical stochastic outcomes can emerge from algebraically distinct event structures—one collapsing all permutations into equipotentiality ($\oplus$), the other preserving strict causal ordering ($\otimes$)— “Providing a rigorous formal distinction between process symmetry and outcome symmetry in probabilistic sampling”.
7 A class of operation table for complete sequential addition
This chapter presents the construction of an “operation table for Complete Sequential Addition (CSA)” as a visual and structural demonstration of its behavior within a finite algebraic setting. The content is illustrative and expository in nature; its purpose is to offer an intuitive way to understand the combinatorial structure of CSA, rather than to furnish essential proofs for the core theory of the paper. The Cayley table and the Full CSA Diagram introduced here serve as visual supplements to the algebraic properties established earlier, and are intended for pedagogical and structural‑analysis purposes.
7.1 A finite commutative semigroup with absorbing element and reflexive absorption property & cayley table
In subsection 6.4, we have analyzed the binary operation #\#. Since it satisfies the properties of closure, associativity, and commutativity, the #\# operation performed on n operands is equivalent to the n-ary finitary operation that satisfies permutation equivalence.
Essentially, the operation #\# forms a commutative semigroup. In this subsection, we extend the #\# operation to a class of special binary operations that also form commutative semigroups, and we construct their corresponding Cayley tables. Furthermore, in the next subsubsection, we will establish a connection between these Cayley tables and certain special addition operations.
[…]
$\mathcal{S}=\{v_{1},v_{2},\cdots,v_{n}\}$
[…]
  1. 1.
    Closure: For all $a,b\in\mathcal{S}$, $a\oslash b\in\mathcal{S}$;
  2. 2.
    Associativity: For all $a,b,c\in\mathcal{S}$, $(a\oslash b)\oslash c=a\oslash(b\oslash c)$;
  3. 3.
[…]
  4. 4.
    Absorbing Element Property: There exists an element $v_{abs}\in\mathcal{S}$ such that $a\oslash v_{abs}=v_{abs}\oslash a=v_{abs}$ for all $a\in\mathcal{S}$;
  5. 5.
    Reflexive Absorption Property: For all $\in\mathcal{S}, aa$$\oslash a=v_{abs}$;
  6. 6.
    Uniqueness: For any two distinct non-empty subsets $\mathcal{A}_{1},\mathcal{A}_{2}\subseteq\{v_{1},v_{2},\dots,v_{n}\}$, $\oslash(\mathcal{A}_{1})\neq\oslash(\mathcal{A}_{2})$.
Under the uniqueness condition, the cardinality of the set $\mathcal{S}$ is $|\mathcal{S}|=2^{n}$.
Now we proceed to construct the Cayley table for this finite group. The top priority is to determine the arrangement of elements for the row and column headers. We use properties of the Newton binomial theorem to build this boundary.
For positive integer $n$, given n different variables of $\mathcal{S}$, we build the structure of line and column as follows
$\left(v_{abs},v_{1},v_{2},\cdots,v_{n},\ast(\mathcal{C}_{1}),\ast(\mathcal{C}_{2}),\cdots,\ast(\mathcal{C}_{n})\right)$
[…]
Algorithm 1 Generate algorithm for N-dimensional Cayley table for complete sequential addition
[…]
3:
4:Push [ $v_{err}$ ] into binomial_theorem_collectionbinomial\_theorem\_collection
5:for $i\leftarrow 0$ to N1N-1 do
6:  $cur\_combos\leftarrow[\ ]$
7:  if $i=0$ then
8:   $cur\_combos\leftarrow[\ \left[v_{1}\right],\left[v_{2}\right],\cdots\left[v_{N}\right]\ ]$
9:  else
10:   for pre_combopre\_combo of comb_collection[i1]comb\_collection[i-1] do
11:     $starting\_idx\leftarrow$ index of pre_combopre\_combo’s last element + 1
12:     for $j\leftarrow starting\_idx+1$ to N1N-1 do
13:      $cur\_combo\leftarrow pre\_combo\oplus OIE_{j}|_{\alpha}^{\beta}$
14:      Push cur_combocur\_combo into cur_comboscur\_combos
15:     end for
16:   end for
17:  end if
18:  Push cur_comboscur\_combos into comb_collectioncomb\_collection
19:  Push cur_comboscur\_combos into binomial_theorem_collectionbinomial\_theorem\_collection
20:end for
21:
[…]
23:for comboscombos of binomial_theorem_collectionbinomial\_theorem\_collection do
24:  for combocombo of comboscombos do
25:   Push combocombo into all_combosall\_combos
26:  end for
27:end for
28:
[…]
30:Init asterisk_cayley_table[SIZE][SIZE]asterisk\_cayley\_table[SIZE][SIZE]
31:for $i\leftarrow 0$ to SIZE1SIZE-1 do
32:  for $j\leftarrow 0$ to SIZE1SIZE-1 do
33:   $asterisk\_cayley\_table[i][j]\ \leftarrow\ all\_combos[i]\oplus\ all\_combos[j]|_{\alpha}^{\beta}$
34:   Sort asterisk_cayley_table[i][j]asterisk\_cayley\_table[i][j] with element index with asc order
35:  end for
36:end for
Figure 1: 5-dimensional Full CSA Diagram
Figure 2: 7-dimensional Full CSA Diagram
7.2 Cayley Table Representation of Complete Sequential Addition
In Section 6.1, we constructed a finite commutative semigroup $(\mathcal{S},\oslash)$ with an absorbing element and reflexive absorption property, along with its Cayley table. This section demonstrates that the Complete Sequential Addition (CSA) operation $\oplus$ can be naturally embedded into this semigroup structure.
Let:
  • $\mathrm{oie}_{\mathrm{void}}$ denote the void optional intervals event,
  • $\mathrm{oie}_{1},\mathrm{oie}_{2},\dots,\mathrm{oie}_{n}$ denote $n$ distinct non‑void OIE instances.
Define a mapping $\Phi$:
$\Phi(\mathrm{oie}_{\mathrm{void}})=v_{\mathrm{abs}},\ \Phi(\mathrm{oie}_{i})=v_{i}\quad(i=1,\dots,n).$
For any non‑empty subset $\mathrm{A}\subseteq\{\mathrm{oie}_{1},\dots,\mathrm{oie}_{n}\}$, let $\oplus(\mathrm{A})$ denote the result of applying Complete Sequential Addition to all OIE instances in $\mathrm{A}$. Define:
$\Phi\bigl(\oplus(\mathrm{A})\bigr)=v_{1}\oslash v_{2}\oslash\cdots\oslash v_{k}\bigl(\{v_{i}\mid\mathrm{oie}_{i}\in\mathrm{A}\}\bigr).$
The mapping $\Phi$ preserves the operations, i.e., for any two OIE instances $a,b$:
$\Phi(\oplus(a,b))=\Phi(a)\oslash\Phi(b).$
Interpreting the row/column headers of the Cayley table constructed in Section 6.1 as the $\Phi$-images of OIE instances:
  • $v_{\mathrm{abs}}$ corresponds to $\mathrm{oie}_{\mathrm{void}}$,
  • $v_{i}$ corresponds to $\mathrm{oie}_{i}$,
  • $\oslash(\mathrm{C}_{k})$ corresponds to $\oplus(\mathrm{C}_{k})$.
Then the entry $Z=X\oslash Y$ in the cell $(X,Y)$ of the Cayley table signifies:
$\oplus(\Phi^{-1}(X),\Phi^{-1}(Y))\text{ yields a result corresponding to }\Phi^{-1}(Z).$
The structure of the table reflects the properties of CSA:
  1. 1.
    First row/column entirely $v_{\mathrm{abs}}$: corresponds to the absorption property of $\mathrm{oie}_{\mathrm{void}}$.
  2. 2.
    Diagonal entries all $v_{\mathrm{abs}}$: corresponds to the reflexive absorption $a\oplus a=\mathrm{oie}_{\mathrm{void}}$.
  3. 3.
    Symmetry: reflects the permutational equivalence of CSA.
  4. 4.
    Closure: every entry in the table corresponds to some CSA result.
The Complete Sequential Addition, as an $n$-ary finitary operation, can be fully represented by the Cayley table of the finite commutative semigroup $(\mathcal{S},\oslash)$. This table encodes the algebraic properties of CSA (commutativity, absorption, etc.) in a finite structure, providing a concise algebraic framework for analysing the combinatorial nature of CSA.
FIG. 1 (the Full CSA Diagram) is a geometric visualisation of this Cayley table, illustrating the complete state space and operational relationships of CSA over a given set of basic $\mathrm{OIE}$ instances. And FIG. 2 (the Full CSA Diagram) is a 7-dimensional pic.
7.3 “And” in natural language & “Composition” in programming
The finite commutative semigroup structure expounded in Section 7.1 exhibits a striking correspondence with the implicit operational semantics underlying natural language conjunction and programmatic composition. This subsection elucidates these correspondences, demonstrating that the reflexive absorption property $a\oslash a=v_{abs}$ encodes a universal constraint across linguistic, computational, and physical domains.
In natural language, the conjunction “and” operates under a fundamental presupposition of referential distinctness. When a speaker asserts a proposition of the form “Event A and Event B” there exists an implicit requirement that the referents of A and B be distinguishable within the discourse context. The utterance “Dr.A and Dr.A submitted a paper” is semantically anomalous precisely because it violates this distinctness condition. While natural language processing systems often treat “and” as an idempotent, associative operator (interpreting $\{a,\ a,\ b\}$ as $\{a,\ b\}$), this elides a crucial cognitive constraint: the mental model construction underlying linguistic comprehension cannot sustain identical, indistinguishable participants occupying distinct participant roles simultaneously. This linguistic constraint manifests as self-referential absorption: the composition of an entity with itself fails to produce a meaningful augmentation of the conceptual structure. Instead, it collapses into a null interpretation– analogous to the absorption element $v_{abs}$ in our algebraic structure. The finite commutative semigroup $(\mathcal{S},\oslash)$ thus provides a formal correlate to this linguistic phenomenon, where the reflexive absorption property encodes the ungrammaticality of conjoining indistinguishable referents.
In computer programming, composition operators (function composition, concurrent process spawning, resource allocation) enforce the distinctness constraint syntactically and operationally rather than merely pragmatically. Consider the following paradigmatic cases:
  1. 1.
    Resource Allocation Semantics: The expression bind(socket, address) composed with itself on the same socket descriptor raises a runtime exception (EADDRINUSE), not an idempotent no-op. This reflects the operational reality that a resource cannot be bound twice–a physical constraint encoded as an absorbing error state.
  2. 2.
    Process Composition: In parallel programming frameworks, constructs like par[A, B] effect concurrent execution. If $A=B$ and the operation requires distinct thread contexts, the composition yields a scheduling failure or deduplication, not parallelized self-composition.
  3. 3.
    Functional Composition with Effects: While $f\circ f$ is mathematically valid, in effectful computation (state monads, resource-transforming functions), composing an action with itself may violate constraints.
These examples demonstrate that programming composition, particularly in the presence of state and resources, adheres to the finite commutative semigroup with absorption structure. The void element corresponds to the computational bottom type, error monad, or exception state– proactively excluding invalid compositions before execution, mirroring our OIE framework’s shift from passive observation to active planning.
The epistemological shift proposed in our $\mathrm{OIE}$ framework is thus not merely mathematical but transdisciplinary: it formalizes the universal constraint that composition of indistinguishable entities is not merely redundant but ontologically invalid in any system that schedules, plans, or allocates resources. The Cayley table of Section 6.1, is not merely an algebraic curiosity but a grammar of feasibility applicable to language, computation, and physics alike.
In conclusion, by embedding Complete Sequential Addition within the finite commutative semigroup with absorbing element, we provide a single formalism that subsumes:
  • Linguistic well-formedness constraints on “and”
  • Computational safety constraints on composition
  • Physical feasibility constraints on event scheduling
This unification supports our central thesis: event analysis must transcend passive observation and embrace the algebraic structure of proactive, constraint-aware planning. The reflexive absorption property is not an artifact of our formalism but a feature capturing a deep truth across domains: identity, once used, cannot be reused without consequence.
8 In-depth insights
8.1 “Simultaneous starting” from the Observer’s Viewpoint
The preceding analysis compels us to confront a fundamental epistemological limitation: the notion of “simultaneous starting” as a pointwise equality of timestamps is an observer-centric construct that cannot be empirically verified within any finite precision. As noted in the Introduction, even the most advanced atomic clock (with systematic uncertainty $8\times 10^{-19}$ seconds as of 2024) cannot guarantee simultaneity at the $10^{-1000}$ second scale. This limitation is not merely technological but ontological—observation is, by its nature, a post-hoc act that depends on event execution. The observer is perpetually passive and asynchronous relative to the phenomenon under study.
Traditional physics circumvents this problem by introducing an ad hoc measurement tolerance: if the difference between two starting timestamps falls below a predetermined threshold (e.g., the instrument’s resolution), they are deemed “simultaneous.” This pragmatic approach, however, conflates ontological simultaneity (the existential equality of instants) with epistemic indistinguishability (the observer’s inability to resolve differences). The former is a metaphysical ideal; the latter is a cognitive limitation. Our framework severs this conflation by rejecting the primacy of the observer and instead formalizing simultaneity as a planning construct rather than an observational verdict.
Through the lens of Optional Intervals Events, “simultaneous starting” is not the equality of points $t_{start}^{A}=t_{start}^{B}$ but rather the equipotentiality of access to a shared temporal domain. Specifically, given a domain-filtering interval $(\alpha,\beta)$, the Complete Sequential Addition $\oplus|_{\alpha}^{\beta}$ ensures that every participant OIE retains at least one feasible interval whose start timestamp can assume any value in $\alpha,\beta]$ (Property 3.9, Step 3). The operation does not enforce that all participants actually start at the same moment; it merely guarantees that the opportunity structure is symmetric. This is the algebraic embodiment of “equal opportunity”—a weaker but rigorously definable condition than pointwise equality.
Critically, this formulation reveals that the traditional pursuit of “simultaneity at a point” is a pseudo-problem from the observer’s viewpoint. Consider the Complete Sequential Multiplication $\otimes$, which explicitly orders events: by construction, each participant’s interval must terminate before the next begins ($te_{i}\leq ts_{i+1}$). Yet, if the observer’s measurement granularity $\Delta t$ satisfies $\Delta t>\max_{i}(te_{i}-ts_{i})$, the entire sequence becomes observationally indistinguishable from a truly simultaneous start. A coarse-grained observer would conclude “all events began together,” while a fine-grained observer would detect sequentiality. Both conclusions are observer-relative truths, not objective facts about the world. The world itself is agnostic to our observational resolutions; it is we who impose categories of “simultaneous” or “sequential” ex post facto.
Our OIE framework inverts this relationship. Instead of deriving structure from observation, we prescribe structure through planning. The void element $\mathrm{oie}_{\mathrm{void}}$ serves as an absorbing witness to ontological impossibility: if a combination of events violates physical or logical constraints (e.g., $a\oplus a$), the result is not a noisy measurement but an explicit $\mathrm{oie}_{\mathrm{void}}$. This is the algebraic equivalent of a refutation: the plan is infeasible, independent of any observer’s capacity to detect the failure. In contrast, traditional physics would require executing the events, observing the collision (e.g., resource deadlock), and only then concluding incompatibility. Our approach eliminates this passive trial-and-error by embedding feasibility as a structural invariant within the algebra itself.
Thus, the observer’s role is demoted from arbiter of simultaneity to consumer of planning outputs. Armed with the feasible set $\mathcal{F}_{\mathrm{oie}_{\mathrm{res}}}$ produced by $\oplus$ or $\otimes$, the observer may subsequently choose a specific 2-tuple $(ts,te)$ from the abstracted interval set $\mathcal{I}_{\mathrm{oie}_{\mathrm{res}}}$ to instantiate a concrete Event. This two-stage process—first planning (generating OIEs), then projecting (selecting timestamps)—decouples the formal structure from observational artifacts. The observer never “discovers” simultaneity; they merely sample from a pre-validated space of possibilities where the distinction between simultaneous and sequential has been rendered mathematically irrelevant.
In conclusion, the pursuit of pointwise simultaneity is an observer-centric and passive. Our algebraic formalism replaces it with a planning-centric, active paradigm where “simultaneous starting” can be reconceptualized as equipotentiality within a domain, and where observer limitations are acknowledged not as approximations to an ideal but as derivative quantities layered atop a more fundamental, observer-independent structure. The world need not conform to our clocks; rather, our clocks are tools for navigating a world whose algebra structure is introduced.
8.2 Observation, Relativistic Physics and All Observable Entities in Physics
Within the empirical tradition of physics, “observation” is regarded as the fundamental means of acquiring objective knowledge. However, a deeper analysis of the act of observation itself reveals inherent subjective factors. Taking the judgment of whether two events “start simultaneously” as an example, the data on which this operation relies—whether timestamps recorded by atomic clocks or start signals captured by photoelectric sensors—are not completely objective “raw facts.” Every stage in the generation, transmission, processing, and interpretation of these data is shaped by multiple human‑dependent factors: instrumental precision, environmental noise, the observer’s chosen reference frame, data‑processing algorithms, and even theoretical presuppositions.
Relativistic physics, while addressing the problem of simultaneity by introducing spacetime geometry and Lorentz transformations—thereby parameterizing the observer’s state of motion and enabling the “computation” of time coordinates in different reference frames—does not fundamentally resolve the subjectivity of observation. It merely adds more variables and more complex mathematical relations to the computational model. For the “precise” spacetime coordinates $(t,x,y,z)$ thus computed, we still cannot obtain them with infinite precision through any observational means. Relativity corrected the notion of absolute spacetime, but it remains based on the assumption of “ideal observation”—that the quantities computed by the formulas can, in principle, be approached arbitrarily closely by infinitely refined measurements. Yet one of the core theses of this paper is that we are not pursuing an infinitely approachable “accurate value”; rather, we fundamentally acknowledge that, no matter how precise the theoretical computation, any physical quantity that must be “observed” by humans or instruments inevitably carries irreducible uncertainty, approximation, and model‑dependence.
This realization can be greatly extended. It is not only temporal “instants” that cannot be observed precisely; the same holds for spatial “coordinate points.” Any physical process used for localization (e.g., laser ranging, GPS signals) has its inherent limit of accuracy. Going further, temperature, pressure, electromagnetic‑field strength, particle spin, quantum states…every element in physics that needs to be observed yields “observed values” that are not direct mirror images of the object’s properties, but products of the interaction between the object and the observational system (including instruments, environment, and the theoretical interpretive framework).
This is precisely the deeper significance and expansive potential of the Optional Intervals Event (OIE) model and sequential‑operations framework proposed in this paper. The core idea of the OIE 4‑tuple $(\mathcal{C},\mathcal{F},\mathcal{I},\mathcal{A})$ is to transform the traditional description of events, which relies on “precise observed values,” into a pre‑planning and structural representation of possibility intervals. $\mathcal{I}$ is no longer a single determinate interval, but a set of optional intervals; $\mathcal{F}$ further prescribes how these intervals interrelate and constrain one another in multi‑event combinations.
This model possesses a natural potential for multidimensional extension:
  1. 1.
    From one‑dimensional time to multi‑dimensional physical quantities: We can generalize the notion of “interval” in the model from a mere time interval $(t_{s},t_{e})$ to an interval of any physical quantity. For instance, one may define a “spatial OIE” whose $\mathcal{I}$ represents the set of possible spatial‑coordinate ranges of an object, or a “temperature OIE” whose $\mathcal{I}$ represents the set of possible temperature ranges of a system.
  2. 2.
    Description of composite physical quantities: For a complex event, one can build OIE models for each of its key physical quantities and use sequential operations ($\oplus$ or $\otimes$) to characterize the concurrent or sequential constraints among them. For example, describing a chemical reaction might involve a “reaction‑time OIE,” a “temperature OIE,” a “pressure OIE,” and a “concentration OIE” simultaneously; the $\oplus$ operation could indicate that these parameters can vary independently within some feasible domain, while the $\otimes$ operation could indicate that they must change according to a specific order or functional relationship.
  3. 3.
    From determinism to possibility‑based planning: This breaks the paradigm often pursued in traditional physical simulation of a single determinate trajectory, and instead explicitly acknowledges and handles the inherent uncertainty and multiple possibilities in physical processes. It is particularly crucial for understanding chaotic systems, quantum processes, and any system that has measurement limits or control errors.
The establishment of this framework has profound implications for several fields:
  • Theoretical physics: It provides a new formal tool for thinking about physical theories that go beyond exact solutions and observational limits. It can help formalize issues in measurement theory, the “fuzziness” in interpretations of quantum mechanics, and the possible non‑exactness of spacetime itself in frontier areas such as quantum gravity.
  • Experimental physics: It offers a new paradigm for experimental design and data analysis. An experiment is no longer merely to “measure” a value, but to “constrain” or “verify” a possibility space represented by OIEs. The interpretation of experimental results would be directly linked to updates of $\mathcal{I}$ or $\mathcal{F}$.
  • Engineering fields: In robotics planning, autonomous driving, aerospace, distributed‑system scheduling, etc., systems must make decisions under limited sensor accuracy and environmental uncertainty. The OIE model provides a rigorous mathematical foundation for this kind of “uncertainty‑based planning,” enabling systems to explicitly handle the feasible intervals of various physical quantities (position, time, battery level, signal strength, etc.) and their combinatorial constraints, thereby making more robust and safer plans.
In summary, the OIE model and its algebraic operations shift our perspective from “pursuing unattainable precise observation” to “managing an unavoidable space of uncertainty.” It no longer asks “What is the value?” but rather “Where might the value be, and how do these possibilities interact?” This shift not only deepens the understanding of the observer’s role since relativity, but also provides a promising formal path for addressing the ubiquitous limitations of observation in physics and across a broad range of scientific and technological disciplines.
9 Comparison with other algebraic methods of describing events & Comparison with other physical or philosophical understandings of simultaneity
9.1 Comparison with other algebraic methods of describing events
Model Type Differences from OIE and Problems OIE Can Solve That They Cannot
Process Algebra - Different focus: Process algebra emphasizes event ordering, synchronization, and communication, while OIE focuses on temporal intervals and their optionality. - Temporal expressiveness: Process algebra lacks native constructs for representing flexible time windows or interval-based uncertainty. OIE explicitly models optional intervals and domain-filtering operations. - Problem-solving capability: OIE can schedule events under temporal uncertainty and interval constraints, which are not naturally expressible in classical process algebra.
Petri Nets - Representation granularity: Petri nets model concurrency via tokens and places, but do not inherently capture interval-based temporal constraints. OIE explicitly models start/end timestamps and interval sets. - Temporal planning: Petri nets are not designed for proactive interval planning or feasibility checking before execution. OIE supports constraint-aware planning and void-result detection for infeasible schedules. - Problem-solving capability: OIE can handle problems where events have multiple feasible time windows and must be arranged under domain constraints—a scenario not naturally supported by standard Petri nets.
Table 1: Comparison between OIE and Classical Concurrency Models
9.2 Comparison with other physical or philosophical understandings of simultaneity
School / Figure Core Tenet vs. OIE Re-casting
Newton
Absolute time supplies a unique world-wide instant; simultaneity = $t_{\mathrm{A}}=t_{\mathrm{B}}$. $\mathrm{OIE}$ drops the absolute clock: only a common launch window $\alpha,\beta]$ is guaranteed, and a point $t$ is a later sampling artifact.
Einstein
Simultaneity is frame-dependent but still point-wise; $t_{\mathrm{A}}=t_{\mathrm{B}}$ within one inertial frame. $\mathrm{OIE}$ keeps frame flexibility by choosing $(\alpha,\beta)$ but replaces point equality with equipotentiality: every participant keeps at least one feasible start inside the shared window.
Penrose
causal-set view
Light-cones abolish a global “same moment”; only partial order $a\preceq b$ matters. $\mathrm{OIE}$ pre-computes the allowed causal stories in the set $\mathcal{F}$ and collapses impossible chains into $\mathit{oie}_{\mathrm{void}}$ before any geometry is drawn.
Wang,Shouren “When you look at flowers, they bloom; when you don’t look, they fade into nothingness.” Simultaneity and event existence depend on conscious perception—events gain determinacy only through the observer’s awareness. $\mathrm{OIE}$ rejects pure perceptual dependence but retains the “subjective constructive” core.
Bridgman
operationalism
Simultaneity = “pointer difference $\varepsilon$”; $\varepsilon$ is an experimental tolerance. $\mathrm{OIE}$ turns the tolerance into a first-class interval width and blocks contradictory plans algebraically rather than post-hoc.
School / Figure Core Tenet vs. OIE Re-casting
Zeilinger
quantum-tag
A measurement assigns a time-label with $\Delta t\,\Delta E\geq\hbar/2$; precision is knowledge-limited. $\mathrm{OIE}$ builds the uncertainty set $\mathcal{I}$ explicitly and treats measurement as a random draw from it, so no further $\hbar$ correction is needed.
Davidson
event ontology
Events are already-happened particulars; “same moment” is a predicate of best descriptions. $\mathrm{OIE}$ treats events as schedulable programs; the algebra decides which descriptions (intervals) are admissible and returns $\mathit{oie}_{\mathrm{void}}$ for self-contradictory narratives.
OIE framework Simultaneity is not only observed but also planned: $\oplus$ guarantees every agent equal access to the launch domain $\alpha,\beta]$, while $\otimes$ imposes a strict order. Point equality is abandoned in favor of interval equipotentiality and hard algebraic feasibility.
10 Derived implications
10.1 Introduce category theory
Since we introduced the natural isomorphism of category theory in defining complete sequential addition and complete sequential multiplication, how to use category theory to describe OIE and the two sequential operations at a more abstract level is clearly a promising direction for further research.
10.2 The extensions of complete sequential addition and complete sequential multiplication
This paper only introduces two types of sequential operations, both of which contain the word “complete”. The reason why we named them this way is that both the starting timestamp and the ending timestamp are deeply involved in the operation rules. Therefore, if they are “incomplete”, then does there exist “sequential addition based on a filtering 2-tuple and the starting timestamp” and “sequential multiplication based on the ending timestamp”?
Judging from the names, in fact, if there is no regulation on when the 100 metres must end, the 100 metres seems like a “sequential addition based on the start time” that we conjecture.
10.3 Time discretization
In this paper, the understanding of time is to regard time as a real number.
Is time distributed as a real number? If time is discretely distributed, are the viewpoints in this paper correct?
Since both Event and $\mathrm{OIE}$ share the same time model, we speculate that a discrete time model will not affect the conclusions of this paper. However, this also requires further in-depth analysis.
10.4 Time variableization
Sequential operations are targeted at time. If we abstract time once again and just regard it as a variable, then by replacing this variable with other physical elements, can the conclusions of this paper still be applied?
For example, if we replace time with location and consider having a meal as an event, then for a person having lunch in Beijing and a person having lunch in Shanghai, can we abstract an operation like this:
$Lunch_{Beijing}\oplus Lunch_{Shanghai}$
Or from a mathematical perspective, if we directly regard time as a variable without physical significance, what is the deeper theoretical foundation of this paper?
The above issues, we believe, are also worthy of in-depth discussion.
10.5 Introduce probability / probability density
For the $\mathrm{Intvl2Tuple}\mathcal{S}$ instance of an $\mathrm{OIE}$ instance, if it is a finite set, what is the probability of each set element (that is, each interval) being used? If it is an infinite set, what is the probability density? This paper does not cover these aspects.
After introducing probability, both types of sequential operations will yield more probabilistic properties. Therefore, this is a problem with great potential for application.
10.6 Other types of sequential operations
Do sequential subtraction and sequential division exist? If they do, what are their algebraic properties?
10.7 Full CSA Diagram
The N-dimensional Full CSA diagram is not only a kind of Cayley table for complete sequential addition.
There is a class of operations, they satisfy the closure property, the commutative law, and the associative law, but the operation between two identical numbers is not allowed. Their Cayley tables are all N-dimensional Full CSA diagrams.
For example, the four major oceans of the Earth consist of the Pacific Ocean, the Atlantic Ocean, the Indian Ocean, and the Arctic Ocean, which can be expressed as
$\text{Pacific Ocean}+\text{Atlantic Ocean}$
$\qquad+\text{Indian Ocean}+\text{Arctic Ocean}$
Each ocean can only appear once in the expression. An expression with repeated occurrences is incorrect. For example, the Pacific Ocean should not participate in the composition of the world’s oceans twice. If participating twice is wrong, then
$\text{Pacific Ocean}+\text{Pacific Ocean}=Error$
Obviously, the 4-dimensional Full CSA diagram can be used to represent the ways of combining the four major oceans.
Let’s take another classic problem in the field of computer science. “Composition” and “Inheritance” are two important fundamental means for handling code reuse.
Inheritance specifically refers to a way of creating new classes in object-oriented programming. A subclass can inherit the attributes and methods of its superclass and thus expand the functions of the superclass. Inheritance represents an “Is-a” relationship.
Composition is not limited to object-oriented programming. It usually means that a class (or structure) contains an object (or structure) of another class as its member variable, and the required functions are implemented through these member objects. Composition represents a “Has-a” relationship.
Suppose a class $Class_{A}$ is composed of three classes
$Class_{A}=Class_{1}+Class_{2}+Class_{3}$
Then their objects are
$Object_{A}=Object_{1}+Object_{2}+Object_{3}$
In this way, through the composition of the three objects, $Object_{A}$ can implement multiple functions. Obviously, the same class cannot participate in the composition twice.
So we believe that the algebraic properties of composition in the field of programming have many similarities with complete sequential addition, and the N-dimensional Full CSA diagram can be used. Since composition can be regarded as addition, could inheritance be a certain kind of multiplication? We think this is possible and worthy of further analysis. From this perspective, there is still a lot of structured work that has not been done in the field of object-oriented programming in computer science alone.
In conclusion, similar scenarios can be found in many fields, so the applicable range of the N-dimensional Full CSA diagram is very wide.
Another aspect worthy of in-depth study is the diagram itself. As the number of dimensions increases, many complex and beautiful curves and shapes appear. What can these shapes and curves be used for, and what special properties do high-dimensional diagrams have? These are also issues worthy of observation.
10.8 Non-classical mechanics
The contents discussed in this article are all within the scope of classical Newtonian mechanics. So, are they still applicable within the scope of non-classical mechanics?
10.9 Optimization of Cartesian Product
Both “complete sequential addition” and “complete sequential multiplication” involve the Cartesian product of sets. Although when applying sequential operations, we mostly do not perform Cartesian products with heavy computation loads, we only analyze algebraic properties to make decisions or explore the derived conclusions.
However, if someone forcefully needs to traverse the results of the Cartesian product, numerous existing optimization schemes are available for Cartesian product computation, such as the application of database technologies and various pruning algorithms. While we do not recommend allocating computing resources to Cartesian product calculations when using the Optional Intervals Event framework, it remains a worthwhile research topic to explore how to introduce database-based solutions, or alternatively, integrate the OIE framework and sequential operations into database technologies, if Cartesian product operations are deemed necessary for sequential computations in a handful of specific scenarios.
Appendix A The simulation of $\mathrm{OIE}$ and 2 sequential operations
(Property 3) Ordering of starting timestamp and ending timestamp:
(Definition 4) Event with undetermined interval:
(Definition 6) Optional Intervals Event:
(Definition 7) The bound 2-tuple of a non-empty finite $\mathrm{2Tuple}\mathcal{T}$ instance:
(Property 5) The relationship between the $\mathcal{F}$(2nd element) and $\mathcal{I}$(3rd element) of $\mathrm{OIE}$:
(Definition 8) Equality of OIE instances:
(Definition 11) Set of $\mathcal{I}$ of finite $\mathrm{OIE}\mathcal{S}$ instance:
(Definition 12) Void $\mathrm{OIE}$:
(Definition 14)Function to get the set that is naturally isomorphic to the Cartesian product of all members of a finite $\mathrm{2Tuple}\mathcal{SS}$ instance in an index order:
(Definition 15) Set of tuples of infeasible interval 2-tuples of all members of a finite $\mathrm{OIE}\mathcal{S}$ instance under an index order:
(Definition 18 / function 1) Get the minimum 1st item of a $\mathrm{2Tuple}\mathcal{T}$ instance:
(Definition 18 / function 2) Get the maximum 2nd item of a $\mathrm{2Tuple}\mathcal{T}$ instance:
(Definition 19) Domain-filtered subset of a non-empty set of tuples of 2-tuples under a domain-filtering 2-tuple:
(Definition 20) Complete Sequential Addition:
(Definition 21) Ascending ordered filtered subset of the set of tuples of 2-tuples:
(Definition 22) Complete Sequential Multiplication:
(Property 11) Sequential operations involving identical $\mathrm{OIE}$ instances result in an $\mathrm{OIE}_{\mathrm{void}}$ instance:
(Property 12) The sequential operations involving $\mathrm{OIE}_{\mathrm{void}}$ instances result $\mathrm{oie}_{\mathrm{void}}$:
(Definition 24) Natural Complete Sequential Addition:
(Definition 27) The 1st type of implementation of an OIE instance:
(Definition 28) The 2nd type of implementation of an OIE instance:
For test cases running of complete sequential addition and complete sequential multiplication, visit
Appendix B The Generation of N-dimensional Full CSA Diagram
The simulated code implementation of Algorithm 1:
Generate the binomial theorem collection as edge of Full CSA diagram:
https://github.com/yleen/LiZhongYuan_diagram/blob/master/cayley_table/main.py#L46.